### v1

# Wall crossing and the Fourier–Mukai transform for Grassmann flops

Abstract.
We prove the crepant transformation conjecture for relative Grassmann flops over a smooth base $B$. We show that the $I$-functions of the respective GIT quotients are related by analytic continuation and a symplectic transformation. We verify that the symplectic transformation is compatible with Iritani’s integral structure, that is, that it is induced by a Fourier–Mukai transform in $K$-theory.
1. Introduction
The connection between birational geometry and quantum cohomology has been a source of great activity and interest in Gromov–Witten theory since the early development of the theory. As first described in Ruan’s crepant resolution conjecture in [Rua99] and later extended to more general crepant transformations in [LR01, CIT09, CR13, LLW10], it is expected that for a general crepant rational map $f:X_{-}\dashrightarrow X_{+}$, the quantum cohomology of $X_{-}$ should be equivalent to that of $X_{+}$ in a specific sense.
More specifically, suppose we have rational maps of smooth projective varieties
[…]
such that $f_{-}^{*}(K_{X_{-}})=f_{+}^{*}(K_{X_{+}})$. The crepant transformation conjecture predicts that generating functions of genus zero Gromov–Witten invariants of $X_{-}$ and $X_{+}$ are equal after analytic continuation. Versions of the conjecture have now been proven in many specific instances, and for certain classes of crepant transformations (see e.g. [LR01, BG09, CIJ18, ILLW12]).
One can further refine the crepant transformation conjecture to take into account Iritani’s integral structure (see [Iri09]) or the rational structure of Katzarkov–Kontsevich–Pantev (see [KKP08]). With respect to this structure, the correspondence should be compatible with the Fourier–Mukai transform
$\mathbb{FM}={f_{+}}_{*}f_{-}^{*}:K^{0}(X_{-})\to K^{0}(X_{+}).$
In this paper, we prove a crepant transformation conjecture for relative Grassmann flops and show it is compatible with the Fourier–Mukai transform.
1.1. Grassmann flops
A rich source of examples of crepant transformations as in (1.0.1) arises from variation of GIT. Namely, suppose a reductive group $G$ acts on a variety $V$. Then we can define two GIT quotients
$X_{\pm}=[V\sslash_{\pm}G],$
where $+/-:G\to\mathbb{C}^{*}$ are characters lying in adjacent maximal chambers with respect to the wall and chamber structure on $\operatorname{Hom}(G,\mathbb{C}^{*})\otimes_{\mathbb{Z}}\mathbb{R}$. Under certain numerical conditions, the rational map $X_{+}\dashrightarrow X_{-}$ will be crepant. In this context, the relation between the Gromov–Witten theory of $X_{+}$ and $X_{-}$ is known as wall crossing, as the variation of GIT between the two varieties involves crossing a codimension-one wall in the $G$-ample cone.
In the present paper, we focus on a particular variation of GIT known as a Grassmann flop. This is one of the simplest examples of variation of GIT with respect to a quotient by a non-abelian group. In fact, we work in the setting of relative Grassmann flops, as we describe below.
The setup is as follows: given a smooth projective variety $B$, let $F=\oplus_{i=1}^{n}M_{i}$ and $E=\oplus_{j=1}^{n}L_{j}$ be sums of line bundles on $B$. For $r<n$, define the vector bundle $V\to B$ to be the total space of
$\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
There is a natural (fiberwise) action of $G=\operatorname{GL}_{r}$ on $V$. Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$ respectively. Then
$X_{+}=$ $\operatorname{tot}(S_{+}\otimes p^{*}F^{\vee})\to\operatorname{Gr}(r,E)$
$X_{-}=$ $\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}(r,F^{\vee})$
where $p:\operatorname{Gr}(r,E)\to B$ (resp. $p:\operatorname{Gr}(r,F^{\vee})\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp. $F^{\vee}$) and $S_{\pm}$ denotes the rank-$r$ tautological bundle on the associated Grassmann bundle. We will work equivariantly with respect to the action of $\mathbb{T}=(\mathbb{C}^{*})^{2n}$ induced by the natural action on $E\oplus F$.
Example 1.1.
We give an example of a relative Grassmann flop realizing a variation of GIT of quiver varieties.
Fix integers $0<r<k<n$ and let $B$ be the Grassmannian $\operatorname{Gr}(k,n).$ Let $F\to B$ be the trivial rank $k$ bundle $\mathcal{O}_{B}\otimes\mathbb{C}^{k}$ and let $E\to B$ be the rank $k$ tautological bundle on $B$. In this case, $X_{+}$ is
$X_{+}=\operatorname{tot}(\hat{S}_{r}^{\oplus k})$
where $\hat{S}_{r}$ denotes the rank $r$ tautological bundle on $\operatorname{Fl}(r,k,n).$
On the other hand, $X_{-}$ is the total space of a vector bundle over $\operatorname{Gr}(r,k)\times\operatorname{Gr}(k,n).$ If we let $\bar{S}_{r}$ denote the tautological rank $r$ bundle from the first factor and $S_{k}$ the tautological rank $k$ bundle from the second factor, then
$X_{-}=\operatorname{tot}(\bar{S}_{r}^{\vee}\otimes S_{k}).$
This example can be realized as an instance of variation of GIT of quiver varieties. Consider the quiver
[…]
The associated quiver varieties (for various choices of stability conditions) are quotients of
$\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{r})\times\operatorname{Hom}(\mathbb{C}^{r},\mathbb{C}^{k})\times\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{n})$
by the natural right action of $\operatorname{GL}_{r}\times\operatorname{GL}_{k}$ where $\operatorname{GL}_{r}$ acts trivially on $\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{n})$ and $\operatorname{GL}_{k}$ acts trivially on $\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{r})$. Then $X_{+}$ and $X_{-}$ from above are GIT quotients of this representation associated to the characters
$+:(g_{1},g_{2})\mapsto\det(g_{1})\cdot\det(g_{2})$
and
$-:(g_{1},g_{2})\mapsto\det(g_{1})^{-1}\cdot\det(g_{2})$
respectively.
1.2. Statement of theorem
With $X_{\pm}$ as above, consider the infinite dimensional (symplectic) vector space
$\mathcal{H}_{X_{\pm}}=H^{*}_{\mathbb{T}}(X_{\pm})(\!(z^{-1})\!)[\![Q,q_{\pm}]\!].$
Here $Q$ and $q_{\pm}$ are Novikov parameters with respect to the base $B$ and fiber directions of $X_{\pm}\to B$, respectively.
Using the relative quasimap theory of [Oh21], one can define $I$-functions
$I_{X_{\pm}}(Q,q_{\pm},z)$
lying in $\mathcal{H}_{X_{\pm}}$. These functions are shown to lie on Givental’s Lagrangian cone $\mathcal{L}_{X_{\pm}}\subset\mathcal{H}_{X_{\pm}}$ for the respective spaces and may therefore be expressed as generating functions of genus-zero Gromov–Witten invariants of $X_{\pm}$. See Section 3 for details.
One can then extend $\mathcal{H}_{X_{\pm}}$ by ~$\widetilde{\mathcal{H}}_{X_{\pm}}=\mathcal{H}_{X_{\pm}}[\log z][z^{-1/2}]$, and following Iritani in [Iri09], define the map
$\Psi:K^{0}_{\mathbb{T}}(X)$$\to\widetilde{\mathcal{H}}_{X}$
$A$ $\mapsto z^{-\mu}z^{\rho}\left(\hat{\Gamma}_{X}\cup(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\right)$
where $\mathbbm{i}=\sqrt{-1}$, $\mu$ is the grading operator, $\rho=c_{1}^{\mathbb{T}}(X_{\pm})$, and $\hat{\Gamma}_{X}$ is the Gamma class—a characteristic class associated with the Gamma function $\Gamma(1+z)$.
We prove the following:
Theorem 1.2 (Theorem 5.4).
There exists a linear symplectic isomorphism
$\mathbb{U}:\widetilde{\mathcal{H}}_{X_{-}}\to\widetilde{\mathcal{H}}_{X_{+}}$
such that the following diagram commutes:
(1.2.1) ${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\mathbb{U}}$
There is a path $\hat{\gamma}$ from a neighborhood of $q_{+}=0$ to a neighborhood of $q_{-}=q_{+}^{-1}=0$ such that
(1.2.2) $\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}},$
where $\widetilde{I_{X_{+}}}$ denotes the analytic continuation of $I_{X_{+}}$ along $\hat{\gamma}.$
Theorem 1.2 generalizes previous known cases of the crepant transformation conjecture in two directions. First, it is to our knowledge one of the first examples of a wall crossing result for a variation of GIT quotients by a non-abelian group—in this case $\operatorname{GL}_{r}$. Second, it provides an example of the conjecture for fiber bundles over an arbitrary smooth base $B$. Thus our result is new even in the special case of the quotient by $\operatorname{GL}_{1}=\mathbb{C}^{*}$, where the two GIT quotients are toric bundles over $B$.
1.3. Strategy of proof
The proof relies on the abelian/non-abelian correspondence in Gromov–Witten theory, which allows us to construct the $I$-functions $I_{X_{\pm}}$ for the GIT quotients $X_{\pm}$ in terms of $I$-functions of the associated abelian quotients
$X_{T,\pm}:=V\sslash_{\pm}T,$
where $T$ is the maximal torus in $G$ consisting of diagonal matrices. More precisely, the function $I_{X_{\pm}}(Q,q_{\pm},z)$ is obtained as the specialization of
$I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,\pm},\ldots,q_{r,\pm},z)$
at $q_{1,\pm}=\cdots=q_{r,\pm}=q_{\pm}.$ Here $I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ is a twist of the $I$-function for the associated abelian quotient by the root bundle for $G$, which we denote by $\mathfrak{g}/\mathfrak{t}$. See Section 3 for details.
The technical heart of the proof relies on computing the analytic continuation of $I_{X_{+}}$ along the path $\hat{\gamma}$ and using this to determine $\mathbb{U}$. It is not clear how to compute $\widetilde{I_{X_{+}}}$ directly. Instead we compute the analytic continuation of $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\ldots,q_{r,+},z)$ along each of the variables $q_{k,+}$ independently. That is to say, we analytically continue $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\ldots,q_{r,+},z)$ along a concatenated path $\gamma_{1}\star\cdots\star\gamma_{r}$, where each $\gamma_{k}$ leaves the variables $q_{i,+}$ constant for $i\neq k$, and goes from 00 to $\infty$ in the $q_{k,+}$ direction. We then homotope the path $\gamma_{1}\star\cdots\star\gamma_{r}$ to a path $\hat{\gamma}$ which lies entirely in the locus $q_{1,+}=\cdots=q_{r,+},$ thereby obtaining $\widetilde{I_{X_{+}}}$.
Finally, we must verify that the linear map $\mathbb{U}$ identifying $I_{X_{-}}$ with $\widetilde{I_{X_{+}}}$ is compatible with the Fourier–Mukai transform. This involves an explicit computation of $\mathbb{FM}$ on generators of $K^{0}_{\mathbb{T}}(X_{-})$. The proof that the map $\mathbb{U}$ in (1.2.1) satisfies (1.2.2) then involves a nontrivial identity of anti-symmetric functions (Lemma 5.2).
1.4. Relation to other work
The structure of the proof follows that given in [CIT09, CR10, CIR14, CIJ18]. In particular the formulation of the correspondence and the proof via localization is closely related to [CIJ18]. The new ingredients in the current article are the homotopy of the path of analytic continuation, the generalization to the relative setting, and the combinatorial identity verifying agreement with $\mathbb{FM}$, which does not appear in the case of abelian variation of GIT.
In the final stages of writing this paper, we learned of a related result by Lutz–Shafi–Webb (see [LSW]). While both papers use the idea of homotoping the path of analytic continuation, there are significant differences in the two results and their proofs. The most important such difference is in the analysis of the symplectic transformation $\mathbb{U}$. In this paper, we compare the map $\mathbb{U}$ to the Fourier–Mukai transform $\mathbb{FM}$. In [LSW], they instead compare $\mathbb{U}$ with the symplectic transformation of the associated abelian wall crossing. This leads to two different proofs that the map $\mathbb{U}$ is a symplectic isomorphism and has a well-defined non-equivariant limit. See Remark 5.5 for an explicit comparison of approaches.
The two papers are in fact complementary, in that our respective results may be combined to show that the abelian/non-abelian correspondence is compatible with the Fourier–Mukai transform in $K$-theory. Such a compatibility is not a priori obvious, as it is not clear how best to relate the kernel $\tilde{X}$ of the Fourier–Mukai transform of the Grassmann flop to the kernel of Fourier–Mukai transform of the associated abelian wall crossing. We hope to revisit this issue in the future to obtain a more direct and conceptual understanding of this compatibility.
1.5. Acknowledgements
We would like to thank Tom Coates, Wendelin Lutz, Jeongseok Oh, Yongbin Ruan, Ed Segal, Qaasim Shafi, and Rachel Webb for the useful discussions.
N. Priddis is supported by the Simons Foundation Travel Grant 586691. M. Shoemaker is supported by the Simons Foundation Travel Grant 958189. Y. Wen is supported by a KIAS Individual Grant (MG083902) at the Korea Institute for Advanced Study.
2. Geometric setup
Let $E=\oplus_{j=1}^{n}L_{j}$ and $F=\oplus_{i=1}^{n}M_{i}$ be sums of line bundles over a smooth projective variety $B$. Choose $r<n$, and define the vector bundle $V\to B$ to be the total space of
$\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
This vector bundle comes equipped with a (right) action of $G=\operatorname{Aut}(\mathbb{C}^{r})=\operatorname{GL}_{r}$ given fiberwise by
$(X,Y)\cdot g=(g^{-1}X,Yg).$
Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$ respectively. Then
$X_{+}=$ $\operatorname{tot}(S_{+}\otimes p^{*}F^{\vee})\to\operatorname{Gr}(r,E)$
$X_{-}=$ $\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}(r,F^{\vee})$
where $p:\operatorname{Gr}(r,E)\to B$ (resp. $p:\operatorname{Gr}(r,F^{\vee})\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp. $F^{\vee}$) and $S_{\pm}$ denotes the rank-$r$ tautological bundle on the associated Grassmann bundle. In other words, the fiber of
$\operatorname{Gr}(r,E)\to B$
over a closed point $b\in B$ is the Grassmannian $\operatorname{Gr}(r,E_{b})$ where $E_{b}$ is the fiber of $E\to B$ over $b$.
The birational map
$f:X_{-}\dashrightarrow X_{+}$
is a relative version of what is known as a Grassmann flop. We construct a resolution of this map in Section 2.3 following [BLVdB16], and compute the associated Fourier–Mukai transform in $K$-theory.
2.1. Cohomology
Let $(\mathbb{C}^{*})^{n}$ act on $F$ by scaling, such that the $i$th factor scales $M_{i}$. Let a different copy of $(\mathbb{C}^{*})^{n}$ act on $E$ similarly. This induces an action of the torus $\mathbb{T}:=(\mathbb{C}^{*})^{2n}$ on both $X_{+}$ and $X_{-}$. Unless otherwise specified, all cohomology groups are $\mathbb{T}$-equivariant. We define $z_{i}=c_{1}^{\mathbb{T}}(M_{i}^{\vee})$ and $x_{i}=c_{1}^{\mathbb{T}}(L_{i}^{\vee})$.
Let $R=\mathbb{C}^{r}$ denote the right representation of $G$ (viewing $\mathbb{C}^{r}$ as row vectors), given by
(2.1.1) $v\cdot g=vg.$
The $G$-equivariant vector bundle $R\times V\to V$ induces a vector bundle on both $X_{+}$ and $X_{-}$ which, by abuse of notation, we also denote by $R$. On $X_{-}$ it is (the pullback of) the tautological bundle $S_{-}\to\operatorname{Gr}(r,F^{\vee})$ and on $X_{+}$ it is (the pullback of) the dual of the tautological bundle $S_{+}\to\operatorname{Gr}(r,E)$. Denote the equivariant Chern roots of the dual $R^{\vee}$ by $y_{1},\ldots,y_{r}$.
Let
$\pi_{\pm}:X_{\pm}\to B$
denote the projection. To simplify notation we will denote $\pi_{\pm}^{*}(x_{i})$ (resp. $\pi_{\pm}^{*}(z_{j})$) simply by $x_{i}$ (resp. $z_{j}$) when no confusion will result. The equivariant cohomology of $X_{+}$ (resp. $X_{-}$), denoted $H^{*}_{\mathbb{T}}(X_{\pm})$, is generated as an algebra by elements of
$H^{*}(B)[z_{i},x_{i}]_{i=1}^{n}$
and $\sigma_{1},\ldots,\sigma_{r}$, the elementary symmetric polynomials in $y_{1},\ldots,y_{r}$, subject to the relations:
(2.1.2) $I_{+}=\left\{\left[\frac{\prod_{i=1}^{n}(1-x_{i})}{\prod_{j=1}^{r}(1+y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}$
in $H^{*}_{\mathbb{T}}(X_{+})$, and
(2.1.3) $I_{-}=\left\{\left[\frac{\prod_{i=1}^{n}(1+z_{i})}{\prod_{j=1}^{r}(1-y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}.$
in $H^{*}_{\mathbb{T}}(X_{-})$. In the above formulas, $\left[-\right]_{l}$ denotes the degree $l$ part. We will denote by $R_{\mathbb{T}}$ the ring $H^{*}_{\mathbb{T}}(pt)$, and by $S_{\mathbb{T}}$ its localization with respect to non-zero homogeneous elements. We will also make use of the completion $\hat{S}_{\mathbb{T}}$, defined by
$\hat{S}_{\mathbb{T}}=\{\sum_{d\in\mathbb{Z}}a_{d}|\text{there exists $d_{0}\in\mathbb{Z}$ such that $a_{d}=0$ for $d<d_{0}$}\},$
where $a_{d}$ lies in the degree $d$ homogeneous part of $S_{\mathbb{T}}.$ We will often work with the localized equivariant cohomology ring $H^{*}_{\mathbb{T}}(X_{\pm})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}},$ or its completion $H^{*}_{\mathbb{T}}(X_{\pm})\otimes_{R_{\mathbb{T}}}\hat{S}_{\mathbb{T}},$
The connected components of the $\mathbb{T}$-fixed locus of $X_{+}$ (resp. $X_{-}$) are indexed by subsets of $\{1,\ldots,n\}$ of size $r$, corresponding to choosing an $r\times r$ minor of $\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ (resp. $\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$). Thus each connected component is isomorphic to $B$. Let $D_{+}$ (resp. $D_{-}$) denote the set indexing these fixed loci. We may represent an element of $D_{\pm}$ by a function $\delta:\{1,\ldots,r\}\to\{1,\ldots,n\}$ such that $\delta(i)<\delta(j)$ for all $i<j$. We will further denote $\delta(i)$ by $\delta_{i}$.
Lemma 2.1.
  1. (1)
    For $\delta^{-}\in D_{-}$, let $B_{\delta^{-}}$ denote the associated fixed locus, isomorphic to $B$. Then $R|_{B_{\delta^{-}}}\cong\oplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}$.
  2. (2)
    For $\delta^{+}\in D_{+}$, let $B_{\delta^{+}}$ denote the associated fixed locus, isomorphic to $B$. Then $R|_{B_{\delta^{+}}}\cong\oplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}$.
Proof.
We will prove the first statement. The fixed locus may be expressed as the GIT quotient
$\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})\sslash_{\theta_{-}}G.$
There is a tautological map of vector bundles $q^{*}(\oplus_{i=1}^{r}M_{\delta^{-}_{i}})\to R^{\vee}$ over the stack
$[\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})/G,]$
where $q:[\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})/G,]\to B$ is the projection. The $\theta_{-}$-stable locus of $\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})$ is exactly the locus where this map is an isomorphism. ∎
Corollary 2.2.
  1. (1)
    In $H^{*}_{\mathbb{T}}(X_{-})$, up to a choice of ordering, the restriction of $y_{i}$ to $B_{\delta^{-}}$ is $-z_{\delta^{-}_{i}}$.
  2. (2)
    In $H^{*}_{\mathbb{T}}(X_{+})$, up to a choice of ordering, the restriction of $y_{i}$ to $B_{\delta^{+}}$ is $-x_{\delta^{+}_{i}}$.
2.2. The associated abelian quotients
Let $T\subset G$ denote the maximal torus in $G$ consisting of diagonal matrices. The characters $\det^{\pm}$ restrict to characters of $T$. We define the abelian quotients $X_{T,\pm}$ associated to $X_{\pm}$ to be the GIT quotients $V\sslash_{\pm}T$. They are fiber products over $B$ of $r$ copies of $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)$ and $r$ copies of $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)$, resp.:
$X_{T,{-}}$ $=\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)\times_{B}\cdots\times_{B}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)$
$X_{T,{+}}$ $=\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)\times_{B}\cdots\times_{B}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right).$
There is a series of intermediate GIT quotients defined as follows. For $0\leq j\leq r$, let $\theta_{j}\in\operatorname{Hom}(T,\mathbb{C}^{*})$ denote the character
$\theta_{j}((t_{1},\ldots,t_{r}))=\prod_{i=1}^{j}t_{i}\cdot\prod_{i=j+1}^{r}t_{i}^{-1},$
and let $X_{T,j}$ denote the GIT quotient $V\sslash_{\theta_{j}}T$:
$X_{T,j}=\prod_{i=1}^{j}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)\times_{B}\prod_{i=j+1}^{r}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)$
where the products above are fiber products over $B$. Then $X_{T,{+}}=X_{T,r}$ and $X_{T,{-}}=X_{T,0}$.
The equivariant cohomology ring $H^{*}_{\mathbb{T}}(X_{T,j})$ is given by the quotient of
$H^{*}_{\mathbb{T}}(B)[z_{i},x_{i}]_{i=1}^{n}[y_{1},\ldots,y_{r}]$
by the relations
[…]
The $\mathbb{T}$-fixed loci of $X_{T,\pm}$ are indexed by (not necessarily injective) functions $f:\{1,\ldots,r\}\to\{1,\ldots,n\}$. For such an $f$, the associated fixed locus $B_{f}$ is the subspace of $X_{T,+}$ such that, on the $i$th factor of $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)$ in the fiber product, the homogeneous coordinates are all zero except for in the $f(i)$th factor of $\mathbb{P}(E)$. In other words,
$B\cong B_{f}=\mathbb{P}(L_{f(1)})\times_{B}\cdots\times_{B}\mathbb{P}(L_{f(r)})\subset\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E)\subset X_{T,+}.$
An analogous description holds for $X_{T,-}$ and indeed all $X_{T,j}$.
We will make use of the following weak form of the abelian/non-abelian correspondence in cohomology:
Proposition 2.3.
The pullback
$H^{*}_{\mathbb{T}}(X_{\pm})\to(H^{*}_{\mathbb{T}}(V^{s,\pm}(G)/T))^{W}$
is an isomorphism, where $V^{s,\pm}(G)$ denotes the stable locus of $V$ with respect to $\det^{\pm}:G\to\mathbb{C}^{*}$ and $W$ denotes the Weyl group. Furthermore, the pullback map
$H^{*}_{\mathbb{T}}(X_{T,\pm})\to H^{*}_{\mathbb{T}}(V^{s,\pm}(G)/T)$
is surjective.
We note that although not originally stated in this context, the result holds equivariantly, as one can check from the explicit presentation of $H^{*}_{\mathbb{T}}(X_{\pm})$ given in Section 2.1. The proposition allows us to lift cohomology classes in $X_{\pm}$ to $W$-invariant classes in $H^{*}_{\mathbb{T}}(X_{T,\pm})$.
For future use we will use $F_{\pm}$ to denote the set of functions $f:\{1,\ldots,r\}\to\{1,\ldots,n\}$, i.e. $F_{\pm}$ indexes the set of fixed loci of $X_{T,\pm}$. We will use $In_{\pm}$ to denote the set of injective functions $\{1,\ldots,r\}\to\{1,\ldots,n\}$, noting that $In_{\pm}$ indexes the fixed loci of $V^{s,\pm}(G)/T.$
2.3. Fourier–Mukai transform
We can define a resolution of the birational map
$f:X_{-}\dashrightarrow X_{+}$
as follows. Let
$\tilde{V}=\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$
and let $\tilde{G}=GL_{r}\times GL_{r}$. Consider the right action of $\tilde{G}$ on $\tilde{V}$ given by
$(X,U,Y)\cdot(g_{1},g_{2})=(g_{1}^{-1}X,g_{2}^{-1}Ug_{1},Yg_{2}).$
Let $\tilde{\theta}\in\hom(\tilde{G},\mathbb{C}^{*})$ denote the character
$\tilde{\theta}(g_{1},g_{2})=\det(g_{1})^{-1}\det(g_{2}).$
Define $\tilde{X}$ to be the GIT quotient
$\tilde{X}=\tilde{V}\sslash_{\tilde{\theta}}\tilde{G}.$
The morphism $f_{-}:\tilde{X}\to X_{-}$ is induced by the morphism of stacks
$[\tilde{V}/\tilde{G}]\to[V/G]$
defined by the map
$\tilde{V}$ $\to V$
$(X,U,Y)$ $\mapsto(X,YU)$
and the homomorphism
$\tilde{G}$ $\to G$
$(g_{1},g_{2})$ $\mapsto g_{1}.$
There is a similar description of the map $f_{+}:\tilde{X}\to X_{+}.$ The $\mathbb{T}$-action on $E$ and $F$ induces an action on $\tilde{X}$ for which the maps $f_{\pm}:\tilde{X}\to X_{\pm}$ are equivariant. We define
(2.3.1) $\mathbb{FM}:K^{0}_{\mathbb{T}}(X_{-})\to K^{0}_{\mathbb{T}}(X_{+})$
to be $\mathbb{FM}=(f_{+})_{*}\circ f_{-}^{*}.$
In [BLVdB16], it is proven that the associated functor $\mathbb{FM}:D^{b}(X_{-})\to D^{b}(X_{+})$, between derived categories is an equivalence in the case that $B$ is a point. In [BCF+21] it is shown that this equivalence may also be described using windows. The arguments naturally extend to our setting. We will use this fact only to conclude that the map (2.3.1) is an isomorphism and preserves the pairing.
We compute $\mathbb{FM}$ using localization as in [Tho92]. The fixed loci of $\tilde{X}$ are indexed by a pair of size-$r$ subsets of $\{1,\ldots,n\}$, corresponding to choosing $r\times r$ minors of both $\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ and $\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$, i.e. they are indexed by elements of $D_{-}\times D_{+}$.
Let $R_{1},R_{2}$ be the representations of $\tilde{G}$ obtained by composing the projection $\pi_{1},\pi_{2}:\tilde{G}\to G$ to $G$ with the right representation $R$ of $G$. Again we abuse notation and also use $R_{1/2}$ to denote the induced vector bundles on $\tilde{X}$. With these conventions, we observe that $f_{-/+}^{*}(R)=R_{1/2}$.
A simple generalization of Lemma 2.1, gives the following.
Lemma 2.4.
For $(\delta^{-},\delta^{+})\in D_{-}\times D_{+}$, let $B_{(\delta^{-},\delta^{+})}$ denote the associated fixed locus, isomorphic to $B$. Then $f_{+}$ (resp. $f_{-}$) maps $B_{(\delta^{-},\delta^{+})}$ isomorphically onto $B_{\delta^{+}}$ (resp. $B_{\delta^{-}}$). Furthermore, $(R_{1})|_{B_{(\delta^{-},\delta^{+})}}\cong\oplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}$ and $(R_{2})|_{B_{(\delta^{-},\delta^{+})}}\cong\oplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}$.
In order to ease notation, in this section we will also use $L_{i}$ or $M_{j}$ to denote the pullbacks of these respective line bundles to $X_{+}$, $X_{-}$, and $\tilde{X}.$ With this convention, for $\delta^{-}\in D_{-}$, let $e_{\delta^{-}}$ denote the class
$e_{\delta^{-}}=$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\mathcal{H}om(M_{j_{-}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})^{\vee}$
(2.3.2) $\cong$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}$
in $K^{0}_{\mathbb{T}}(X_{-})$, where for $A\in K^{0}_{\mathbb{T}}(X_{-})$, we denote $\wedge^{\bullet}A:=\oplus_{i=0}^{\operatorname{rk}(A)}(-1)^{i}\wedge^{i}A$.
Corollary 2.5.
The $\mathbb{T}$-localized equivariant $K$-theory of $X_{-}$ is spanned by the classes $\{\pi_{-}^{*}(A)\otimes e_{\delta^{-}}|A\in K^{0}(B),\delta^{-}\in D_{-}\}$.
Proof.
By Lemma 2.1, one checks that the class $\pi_{-}^{*}(A)\otimes e_{\delta^{-}}$ restricts to zero on $B_{\delta_{0}^{-}}$ for $\delta_{0}^{-}\neq\delta^{-}$. Furthermore, it restricts to the invertible class $\wedge^{\bullet}N_{B_{\delta^{-}}|\operatorname{Gr}(r,F^{\vee})}^{\vee}$ on $B_{\delta^{-}}$. The result then follows from the localization theorem in $K$-theory. ∎
Proposition 2.6.
We have
(2.3.3) $\mathbb{FM}(e_{\delta^{-}})=\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}.$
Remark 2.7.
Note that although the right hand side of the equation above is the same formal expression as the right hand side of (2.3.2), the right hand side of (2.3.3) is to be interpreted as a class in $K^{0}_{\mathbb{T}}(X_{+})$.
Proof.
Note first that, in $K^{0}_{\mathbb{T}}(\tilde{X})$, the class
$\prod_{i_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{i_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}\cdot\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}(R_{2}^{\vee}\otimes R_{1})^{\vee}$
restricts to $\wedge^{\bullet}N_{B_{\delta^{-},\delta^{+}}|\tilde{X}}^{\vee}$ on $B_{(\delta^{-},\delta^{+})}$, and restricts to zero on $B_{(\delta^{-}_{1},\delta^{+}_{1})}$ for $(\delta^{-}_{1},\delta^{+}_{1})\neq(\delta^{-},\delta^{+})$.
It follows that
$f_{-}^{*}(e_{\delta^{-}})=$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}$
$=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\left(\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}}{\prod_{i_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{i_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}\cdot\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}(R_{2}^{\vee}\otimes R_{1})^{\vee}}\right)$
$=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\left(\frac{1}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}(R_{2}^{\vee}\otimes R_{1})^{\vee}}\right)$
$=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\left(\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R_{2}^{\vee})^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}(M_{j}^{\vee}\otimes R_{2}^{\vee})^{\vee}}\right)$
where the second equality is by the localization theorem applied to $K^{0}_{\mathbb{T}}(\tilde{X})$, the fourth is Lemma 2.4. Pushing forward via $f_{+}$, we obtain
$(f_{+})_{*}f_{-}^{*}(e_{\delta^{-}})=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{+}})_{*}i_{\delta^{+}}^{*}\left(\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R)^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}(M_{j}^{\vee}\otimes R^{\vee})^{\vee}}\right)$
$=$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}$
where the last equality is the localization theorem applied to $K^{0}_{\mathbb{T}}(X_{+})$.
For $(\delta^{-},\delta^{+})\in D_{-}\times D_{+}$, define
(2.3.4) $C_{\delta^{-},\delta^{+}}=$ $\prod_{i=1}^{r}e^{(n-r)(x_{\delta^{+}_{i}}-z_{\delta^{-}_{i}})/2}\prod_{j_{-}\notin\delta^{-}}\frac{\sin((x_{\delta^{+}_{i}}-z_{j_{-}})/2\mathbbm{i})}{\sin((z_{\delta^{-}_{i}}-z_{j_{-}})/2\mathbbm{i})},$
where $\mathbbm{i}=\sqrt{-1}$.
Definition 2.8.
Define $\mathbb{U}_{H}:H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to be the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{\delta^{-}})\in H^{*}_{\mathbb{T}}(B_{\delta^{-}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to
(2.3.5) $\sum_{\delta^{+}\in D_{+}}C_{\delta^{-},\delta^{+}}\frac{\phi_{\delta^{+},\delta^{-}}(\alpha)}{e_{\mathbb{T}}(N_{\delta^{+}})}$
where $\phi_{\delta^{+},\delta^{-}}:H^{*}_{\mathbb{T}}(B_{\delta^{-}})\to H^{*}_{\mathbb{T}}(B_{\delta^{+}})$ is the canonical isomorphism induced by the projections $\pi_{\delta^{+}}:B_{\delta^{+}}\to B$ and $\pi_{\delta^{-}}:B_{\delta^{-}}\to B$.
Proposition 2.9.
The following diagram commutes:
[…]
Proof.
We will prove $\operatorname{ch}^{\mathbb{T}}(\mathbb{FM}(e_{\delta^{-}}))=\mathbb{U}_{H}(\operatorname{ch}^{\mathbb{T}}(e_{\delta^{-}}))$. The general statement then follows by Corollary 2.5 and the projection formula. Observe that
$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{-}}^{*}e_{\delta^{-}})=$ $\operatorname{ch}^{\mathbb{T}}\left(i_{\delta^{-}}^{*}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}\right)$
$=$ $\operatorname{ch}^{\mathbb{T}}\left(\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes M_{\delta^{-}_{i}})^{\vee}\right)$
$=$ $\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\left(1-e^{z_{\delta^{-}_{i}}-z_{j_{-}}}\right)$
and
$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{+}}^{*}\mathbb{FM}(e_{\delta^{-}}))=$ $\operatorname{ch}^{\mathbb{T}}i_{\delta^{+}}^{*}\left(\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}\right)$
$=$ $\operatorname{ch}^{\mathbb{T}}\left(\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes L_{\delta^{+}_{i}})^{\vee}\right)$
$=$ $\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\left(1-e^{x_{\delta^{+}_{i}}-z_{j_{-}}}\right)$
The ratio of the two expressions above, after identifying $B_{\delta^{-}}$ and $B_{\delta^{+}}$, is
$\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{\left(1-e^{x_{\delta^{+}_{i}}-z_{j_{-}}}\right)}{\left(1-e^{z_{\delta^{-}_{i}}-z_{j_{-}}}\right)}$
$=$ $\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{e^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\left(e^{-(x_{\delta^{+}_{i}}-z_{j_{-}})/2}-e^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\right)}{e^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\left(e^{-(z_{\delta^{-}_{i}}-z_{j_{-}})/2}-e^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\right)}$
$=$ $C_{\delta^{-},\delta^{+}}.$
By the localization theorem in cohomology (see [AB84]), it follows that the diagram commutes. ∎
Corollary 2.10.
The map $\mathbb{U}_{H}$ is invertible and has a well-defined non-equivariant limit. Furthermore, we have
${\pi_{-}}_{*}\left(\alpha\cup\operatorname{Td}_{X_{-}/B}\right)={\pi_{+}}_{*}\left(\mathbb{U}_{H}(\alpha)\cup\operatorname{Td}_{X_{+}/B}\right).$
Proof.
Let $\overline{X}_{+}$ and $\overline{X}_{-}$ denote the GIT quotients from Section 2 in the special case that the base is a point, so $\overline{X}_{+}\cong\overline{X}_{-}=\operatorname{tot}(S)$, where $S$ is the tautological bundle over $\operatorname{Gr}(r,\mathbb{C}^{n})$. Let
$\overline{\mathbb{FM}}:\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{-})\to\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{+})$
denote the Fourier–Mukai transform in this case, where $\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{\pm})$ denotes the completion of $K^{0}_{\mathbb{T}}(\overline{X}_{\pm})$ along the augmentation ideal (see [EG00]).
By definition of the $\mathbb{T}$ action on $B$, $E$, and $F$, the (completion of the) equivariant cohomology of $B$ is given by
$\hat{H}^{*}_{\mathbb{T}}(B)=H^{*}(B)[\![z_{i},x_{i}]\!]_{i=1}^{n}$
where recall that $x_{i}=c_{1}^{\mathbb{T}}(L_{i}^{\vee})$ and $z_{i}=c_{1}^{\mathbb{T}}(M_{i}^{\vee})$. The equivariant line bundles $L_{i}\to B$ and $M_{i}\to B$ for $1\leq i\leq n$ determine a map
$v:[B/(\mathbb{C}^{*})^{2n}]\to[\{pt\}/(\mathbb{C}^{*})^{2n}].$
This induces the following commutative diagram of stacks:
${\left[X_{\pm}/(\mathbb{C}^{*})^{2n}\right]}$${\left[\overline{X}_{\pm}/(\mathbb{C}^{*})^{2n}\right]}$${\left[B/(\mathbb{C}^{*})^{2n}\right]}$[${\left[\{pt\}/(\mathbb{C}^{*})^{2n}\right].}$$\scriptstyle{\pi_{\pm}}$$\scriptstyle{\tilde{v}_{\pm}}$$\scriptstyle{\overline{\pi}_{\pm}}$$\scriptstyle{v}$
Taking cohomology, we obtain
[…]
This diagram commutes. In fact, after identifying
$\hat{H}^{*}_{\mathbb{T}}(X_{\pm}):=\prod_{i=0}^{\infty}H^{i}_{\mathbb{T}}(X_{\pm})$
with $H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{\pm},$ we have the identification
${\pi_{\pm}}_{*}=\operatorname{id}_{B}\otimes_{\mathbb{C}}{\overline{\pi}_{\pm}}_{*}$
where $\operatorname{id}_{B}$ denotes the identity map on $H^{*}_{\mathbb{T}}(B)$.
After taking completions with respect to the augmentation ideal, the Chern map
$\hat{\operatorname{ch}}^{\mathbb{T}}:\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{\pm})\to\hat{H}^{*}_{\mathbb{T}}(\overline{X}_{\pm})$
is an isomorphism (see [EG00]; here we use $\hat{\operatorname{ch}}^{\mathbb{T}}$ to denote the induced map defined on the completion $\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{\pm})$). Because it is induced by a derived equivalence, the map $\overline{\mathbb{FM}}$ is also an isomorphism. Proposition 2.9 implies that $\overline{\mathbb{U}}_{H}:=\hat{\operatorname{ch}}^{\mathbb{T}}\circ\overline{\mathbb{FM}}\circ(\hat{\operatorname{ch}}^{\mathbb{T}})^{-1}$ is also an isomorphism, and may be defined before localizing with respect to $\mathbb{T}$. In other words, $\overline{\mathbb{U}}_{H}$ gives a well-defined map
$\overline{\mathbb{U}}_{H}:\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{-}\to\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{+}.$
Again using the identification $\hat{H}^{*}_{\mathbb{T}}(X_{\pm})=H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{\pm}$ together with the explicit formula for $\mathbb{U}_{H}$ in (2.3.5), we also see that $\mathbb{U}_{H}=\operatorname{id}_{B}\otimes_{\mathbb{C}}\overline{\mathbb{U}}_{H}$ which implies the first statement.
Given $\overline{\alpha}\in\hat{H}^{*}_{\mathbb{T}}(\overline{X}_{-})$, let $A=(\operatorname{ch}^{\mathbb{T}})^{-1}(\overline{\alpha})\in\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{-})$. Applying the equivariant Hirzebruch–Riemann–Roch theorem (see (3.1.2) below) twice, we see that for $A\in\hat{K}^{0}_{\mathbb{T}}(X_{-})$,
(2.3.7) ${\overline{\pi}_{-}}_{*}\left(\overline{\alpha}\operatorname{Td}_{\overline{X}_{-}}\right)$ $=\chi(A)$
$=\chi(\overline{\mathbb{FM}}(A))$
$={\overline{\pi}_{+}}_{*}\left(\operatorname{ch}^{\mathbb{T}}(\overline{\mathbb{FM}}(A))\operatorname{Td}_{\overline{X}_{+}}\right)$
$={\overline{\pi}_{+}}_{*}\left(\mathbb{U}_{H}(\overline{\alpha})\operatorname{Td}_{\overline{X}_{+}}\right).$
The pullback $\tilde{v}_{\pm}^{*}(\operatorname{Td}_{\overline{X}_{\pm}})$ is the relative Todd class $\operatorname{Td}_{X_{\pm}}$. This fact together with (2.3.7) allows us to upgrade (2.3.6) to the following commutative diagram:
[…]
The second claim follows. ∎
3. Gromov–Witten theory and $I$-functions
Let $X$ be a smooth semi-projective variety. In what follows, $X$ will also be equipped with an action by a torus $\mathbb{T}$. Following [CIJS15], we assume that all $\mathbb{T}$-weights appearing in the representation $H^{0}(X,\mathcal{O}_{X})$ lie in a strictly convex cone, and that $H^{0}(X,\mathcal{O}_{X})^{\mathbb{T}}=\mathbb{C}.$ This implies in particular that $X^{\mathbb{T}}$ is compact.
Following standard practice, we define integrals over $X$ via localization. For $\gamma\in H^{*}_{\mathbb{T}}(X)$, define
$\int_{X}\gamma:=\sum_{F}\int_{F}\frac{(\gamma)|_{F}}{e_{\mathbb{T}}(N_{F|X})}\in S_{\mathbb{T}}$
where the sum is over all connected components $F$ in $X^{\mathbb{T}}$ and the class $e_{\mathbb{T}}(-)$ denotes the $\mathbb{T}$-equivariant Euler class. Define a pairing $\left\langle-,-\right\rangle_{X}$ on $H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ by
$\left\langle\alpha,\beta\right\rangle_{X}:=\sum_{F}\int_{F}\frac{(\alpha\cup\beta)|_{F}}{e_{\mathbb{T}}(N_{F|X})}.$
For a ring $R$, let
$R[\![Q]\!]=\left\{\sum_{d\in\operatorname{Eff}(X)}a_{d}Q^{d}\mid a_{d}\in R\right\}$
denote the Novikov ring of $X$. Let
$\mathcal{H}_{X}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}(\!(z^{-1})\!)[\![Q]\!]$
denote Givental’s symplectic vector space, equipped with the symplectic form
(3.0.1) $\Omega(f,g)=\operatorname{Res}_{z=0}\left\langle f(-z),g(z)\right\rangle_{X}dz.$
This defines a polarization
$\mathcal{H}_{X}=\mathcal{H}_{X}^{+}\oplus\mathcal{H}_{X}^{-}$
with $\mathcal{H}_{X}^{+}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}[z][\![Q]\!]$ and $\mathcal{H}_{X}^{-}=z^{-1}H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}[\![z^{-1}]\!][\![Q]\!]$.
For $\alpha_{1},\ldots,\alpha_{m}\in H^{*}_{\mathbb{T}}(X)$, $k_{1},\ldots,k_{m}\geq 0$, and $d\in\operatorname{Eff}(X)$, define the degree-$d$ genus-0 Gromov–Witten invariant of $X$:
$\left\langle\alpha_{1}\psi_{1}^{k_{1}},\ldots,\alpha_{m}\psi_{m}^{k_{m}}\right\rangle_{0,m,d}^{X}=\int_{[\mathcal{M}_{0,m}(X,d)]^{vir}}\prod_{i=1}^{m}ev_{i}^{*}(\alpha_{i})\psi_{i}^{k_{i}}.$
If $X$ is not compact, the integral on the right hand side is defined after localization.
Definition 3.1.
Define the $\mathbb{T}$-equivariant genus-zero descendent potential to be
(3.0.2) $\mathcal{F}_{X}({\bm{t}}(z))=\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle{\bm{t}}(\psi_{1}),\ldots,{\bm{t}}(\psi_{m})\right\rangle_{0,m,d}^{X},$
for ${\bm{t}}(z)=\sum_{a=0}^{s}t_{a}z^{a}\in H^{*}_{\mathbb{T}}(X)[z]$. Extending linearly, $\mathcal{F}_{X}$ defines a function on $\mathcal{H}_{X}^{+}$.
Definition 3.2.
After applying the Dilaton shift $\sum_{a=0}^{s}t_{a}z^{a}=z1+\sum_{a=0}^{s}q_{a}z^{a}$, we define Givental’s overruled Lagrangian cone $\mathcal{L}_{X}$ to be the graph of $\mathcal{F}_{X}$ with respect to the above polarization:
$\mathcal{L}_{X}=\{(p,q)\in T^{*}\mathcal{H}_{X}^{+}\mid p=d_{q}\mathcal{F}_{X}\}.$
Let $\{\phi_{i}\}_{i\in I}$ be a basis for $H^{*}_{\mathbb{T}}(X)$ and let $\{\phi^{i}\}_{i\in I}$ denote the dual basis. Points of $\mathcal{L}_{X}$ are of the form
$-z1+{\bm{t}}(z)+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},{\bm{t}}(\psi_{2}),\ldots,{\bm{t}}(\psi_{m+1})\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
Definition 3.3.
The $J$-function of $X$ is the slice of $\mathcal{L}_{X}$ obtained by intersecting with $-z1+t+\mathcal{H}_{X}^{-}$ for $t=\sum_{i\in I}t^{i}\phi_{i}\in H^{*}_{\mathbb{T}}(X)$:
$J^{X}(t,z)=-z1+t+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\sum_{i\in I}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},t,\ldots,t\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
The function $J^{X}(t,z)$ determines the entire Lagrangian cone. Via the string equation, dilaton equation, and topological recursion relations, any point on $\mathcal{L}_{X}$ may be expressed as
$J^{X}(\tau,z)+\sum_{i\in I}C^{i}(z)\frac{\partial}{\partial{t^{i}}}J^{X}(t,z)|_{t=\tau}$
where $C^{i}(z)\in S_{\mathbb{T}}[z][\![Q]\!]$ are polynomials in $z$.
3.1. Integral lattice
For $\phi\in H^{2p}_{\mathbb{T}}(X)$ a homogeneous element, let
$\deg(\phi)=2p$
denote the real degree. We will use the following three operators heavily in what follows.
First, we define $\mu$ to be the operator
$\mu(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X)}{2}\right)\phi,$
where $\dim$ denotes the dimension over $\mathbb{C}$.
Secondly, we define $\deg_{0}:H^{*}_{\mathbb{T}}(X)\to H^{*}_{\mathbb{T}}(X)$ be the degree operator, defined by
$\deg_{0}(\phi)=\deg(\phi)\cdot\phi.$
Finally, we define $\rho$ denote the operator of multiplication by $c_{1}(T_{X})$.
For $A\to X$ a vector bundle, define the multiplicative class $\hat{\Gamma}(A)$ to be
$\hat{\Gamma}(A)=\prod_{i=1}^{s}\Gamma(1+a_{i})$
where $a_{1},\ldots,a_{s}$ are the Chern roots of $A$, and $\Gamma(z)$ is the Gamma function expanded as a power series at $z=0$. Denote $\hat{\Gamma}(TX)$ by $\hat{\Gamma}_{X}$.
Define the map
$\Psi:K^{0}_{\mathbb{T}}(X)$$\to\widetilde{\mathcal{H}}_{X}$
$A$ $\mapsto z^{-\mu}z^{\rho}\left(\hat{\Gamma}_{X}\cup(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\right)$
where
~$\widetilde{\mathcal{H}}_{X}=\mathcal{H}_{X}[\log z][z^{-1/2}].$
The image of $\Psi$ defines a lattice in ~$\widetilde{\mathcal{H}}_{X}$. For future use, we will also define the map
(3.1.1) $\psi:\hat{H}^{*}_{\mathbb{T}}(X)$$\to\widetilde{\mathcal{H}}_{X}$
$\phi$ $\mapsto z^{-\mu}z^{\rho}\left(\hat{\Gamma}_{X}\cup(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\phi\right)$
so that $\Psi=\psi\circ\operatorname{ch}^{\mathbb{T}}$.
Define the equivariant Euler characteristic to be
$\chi(A):=\sum_{i=0}^{\dim(X)}(-1)^{i}\operatorname{ch}^{\mathbb{T}}(H^{i}(X,A))$
where, as before, we use the equivariant Chern character with respect to $\mathbb{T}$. If $X$ is compact, then by [EG00, Tho92] we have the following equivariant Hirzebruch–Riemann–Roch theorem
(3.1.2) $\chi(A)=\int_{X}\operatorname{ch}^{\mathbb{T}}(A)\operatorname{Td}_{X},$
as an equality in $\hat{S}_{\mathbb{T}}$ (as defined in Section 2.1). When $X$ is not compact the equivariant Hirzebruch–Riemann–Roch theorem is not known in general, but the following simple case follows easily from the projection formula. Suppose that $Z$ is a smooth projective variety equipped with an action of $\mathbb{T}$, that $V\to Z$ is an equivariant vector bundle, and that $X$ is the total space of $V$. Suppose further that the $\mathbb{T}$-fixed locus of $X$ is contained in $Z$. Then the projection formula in cohomology implies
$\int_{X}\alpha=\int_{Z}s^{*}(\alpha)/e_{\mathbb{T}}(V)$
where $s:Z\to V$ denotes the zero section. On the other hand, the projection formula in $K$-theory implies
$\chi(A)=\chi(s^{*}(A)/\wedge^{\bullet}(V^{\vee})).$
Applying the Hirzebruch–Riemann–Roch theorem (3.1.2) to $Z$ then implies (3.1.2) for $X$.
Following [CIJ18], consider a modified Euler characteristic $\chi_{z}$ given by
$\chi_{z}(C,D):=\left(\frac{2\pi\mathbbm{i}}{z}\right)^{\lambda\partial_{\lambda}}\chi(C,D)$
where $\lambda\partial_{\lambda}:=\sum_{i}\lambda_{i}\frac{\partial}{\partial\lambda_{i}}$ for some basis $\lambda_{1},\ldots,\lambda_{s}$ of $H^{2}(B\mathbb{T})$, .
Lemma 3.4 (Proposition 2.10 of [Iri09]).
The map $\Psi$ relates the pairings in $K$-theory and cohomology as follows:
$\chi_{z}(C,D)=\frac{1}{(2\pi)^{\dim X}}\left\langle\Psi(C)(-z),\Psi(D)(z)\right\rangle_{X}.$
Proof.
Following [CIJ18, Proposition 3.2], we use the following equalities: $z^{-\lambda d\lambda}\left\langle\alpha,\beta\right\rangle_{X}=\left\langle z^{-\mu}\alpha,z^{\mu}\beta\right\rangle_{X}$, $\left\langle z^{\rho}\alpha,z^{\rho}\beta\right\rangle_{X}=\left\langle\alpha,\beta\right\rangle_{X}$, and $e^{2\pi\mathbbm{i}\mu}\rho=-\rho e^{2\pi\mathbbm{i}\mu}$. The proof then follows from the equivariant Hirzebruch–Riemann–Roch theorem.
$\chi_{z}(C,D)=z^{-\lambda\partial\lambda}\frac{1}{(2\pi\mathbbm{i})^{\dim X}}\int_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C^{\vee}\otimes D)\operatorname{Td}_{X}$
$=z^{-\lambda\partial\lambda}\frac{1}{(2\pi\mathbbm{i})^{\dim X}}\int_{X}e^{\pi\mathbbm{i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C^{\vee}\otimes D)$
$=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\int_{X}e^{-\pi\mathbbm{i}\dim(X)/2}e^{\pi\mathbbm{i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C^{\vee}\otimes D)$
$=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\left\langle e^{\pi\mathbbm{i}\rho}e^{\pi\mathbbm{i}\mu}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\right\rangle_{X}$
$=\frac{1}{(2\pi)^{\dim X}}\left\langle z^{-\mu}z^{-\rho}e^{\pi\mathbbm{i}\rho}e^{\pi\mathbbm{i}\mu}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\right\rangle_{X}$
$=\frac{1}{(2\pi)^{\dim X}}\left\langle(e^{-\pi\mathbbm{i}}z)^{-\mu}(e^{-\pi\mathbbm{i}}z)^{\rho}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\right\rangle_{X}$
$=\frac{1}{(2\pi)^{\dim X}}\left\langle\Psi(C)(e^{-\pi\mathbbm{i}}z),\Psi(D)(z)\right\rangle_{X}$
Let us return now to the specific geometry of the relative Grassmann flop of the previous section. Let $X_{+}$ and $X_{-}$ be as in Section 2. Let $\psi_{\pm}$ denote (3.1.1) for these particular spaces.
Definition 3.5.
Define $\mathbb{U}:\mathcal{H}_{X_{-}}\to\mathcal{H}_{X_{+}}$ to be
$\mathbb{U}=\psi_{+}\circ\mathbb{U}_{H}\circ\psi_{-}^{-1},$
where $\mathbb{U}_{H}$ is the map in (2.3.5).
Corollary 3.6.
The map $\mathbb{U}$ preserves the symplectic pairing. In particular,
$\left\langle f(-z),g(z)\right\rangle_{X_{-}}=\left\langle\mathbb{U}f(-z),\mathbb{U}g(z)\right\rangle_{X_{+}}.$
Proof.
By the Leray–Hirsch theorem, any class in $\hat{H}^{*}_{\mathbb{T}}(X_{-})$ may be decomposed as a sum of classes of the form $\pi_{-}^{*}(\beta)\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C})$ for some $\overline{C}\in K^{0}_{\mathbb{T}}(\overline{X}_{-})$ where the map $\tilde{v}_{-}$ appears in the proof of Corollary 2.10. We may therefore assume without loss of generality that $f(z)=\psi_{-}(\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C}))$ and $g(z)=\psi_{-}(\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{D}))$. Using the same chain of equalities from the previous proof, the projection formula, and (2.3.8), one computes that
$\left\langle\psi_{-}(\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C}))(e^{-\pi\mathbbm{i}}z),\psi_{-}(\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{D}))(z)\right\rangle_{X_{-}}$
$=$ $K\cdot\chi_{z}(\overline{C},\overline{D})$
and
$\left\langle\mathbb{U}\psi_{-}(\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C}))(e^{-\pi\mathbbm{i}}z),\mathbb{U}\psi_{-}(\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{D}))(z)\right\rangle_{X_{+}}$
$=$ $K\cdot\chi_{z}(\overline{\mathbb{FM}}(\overline{C}),\overline{\mathbb{FM}}(\overline{D}))$
where
$K=\left(z^{-\lambda\partial\lambda}\frac{(2\pi)^{\dim X}}{(2\pi\mathbbm{i})^{\dim B}}\int_{B}\left((-2\pi\mathbbm{i})^{\deg_{0}/2}\beta_{1}\right)\cup\left((2\pi\mathbbm{i})^{\deg_{0}/2}\beta_{2}\operatorname{Td}_{B}\right)\right).$
Because $\overline{\mathbb{FM}}$ is induced from an equivalence of categories,
$\chi_{z}(\overline{C},\overline{D})=\chi_{z}(\overline{\mathbb{FM}}(\overline{C}),\overline{\mathbb{FM}}(\overline{D})).$
The claim follows.
3.2. $I$-functions for Grassmann bundles
In this section, we record $I$-functions for $X_{\pm}$ and the associated abelian quotients. In each case an $I$-function is obtained by modifying the $J$-function of the base $B$.
3.2.1. $r=1$
We begin with the case $r=1$, in which
$X_{+}=$ $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)$
$X_{-}=$ $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right).$
In this case $X_{\pm}$ is a toric bundle over $B$. $I$-functions for split toric bundles were computed by Brown in [Bro14] in terms of the $J$-function for the base $B$. Let $J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}$ be the $J$-function for $B$. Define the $H^{*}_{\mathbb{T}}(X_{\pm})$-valued function $I_{X_{\pm}}(Q,q,z)$ to be
(3.2.1) $I_{X_{+}}(Q,q_{+},z)=$ $\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{e\in\mathbb{Z}}q_{+}^{-y/z+e}M_{\beta,-e}$
(3.2.2) $I_{X_{-}}(Q,q_{-},z)=$ $\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{d\in\mathbb{Z}}q_{-}^{y/z+d}M_{\beta,d}$
where
(3.2.3) $M_{\beta,d}=\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y+hz)}{\prod_{h=-\infty}^{d-\beta\cdot M_{i}}(z_{i}+y+hz)}\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y+hz)}{\prod_{h=-\infty}^{-d+\beta\cdot L_{k}}(-x_{k}-y+hz)}.$
Remark 3.7.
Despite having the same formal expression under the change of variables $q_{+}=q_{-}^{-1}$, we emphasize that $I_{X_{+}}$ and $I_{X_{-}}$ take very different forms due to the fact that they lie in different cohomology rings. In particular, in $I_{X_{+}}$, the expression $M_{\beta,-e}$ vanishes whenever $e<\operatorname{min}_{k}(-\beta\cdot L_{k})$ by (2.2.1). On the other hand in $I_{X_{-}}$ the expression $M_{\beta,d}$ vanishes when $d<\operatorname{min}_{i}(\beta\cdot M_{i})$.
The following is a special case of Brown’s result for toric bundles (see [Bro14]).
Theorem 3.8.
The function $I_{X_{\pm}}(Q,q_{\pm},-z)$ lies in the $\mathbb{T}$-equivariant Lagrangian cone $\mathcal{L}_{X_{\pm}}$.
3.2.2. The associated abelian quotients
Now assume $r>1$. By iterating the result of the previous section, we obtain $I$-functions for $X_{T,\pm}$, and for all $X_{T,j}$ for $0\leq j\leq s$. Let
$L_{\beta,\vec{d}}=$ $\prod_{j=1}^{r}\left(\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}.\right)$
Define $I_{X_{T,j}}(Q,d_{1},\ldots,d_{j},e_{j+1},\ldots,e_{r},z)$ to be
(3.2.4) $I_{X_{T,j}}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{\stackrel{{\scriptstyle(e_{1},\ldots,e_{j})\in\mathbb{Z}^{j}}}{{(d_{j+1},\ldots,d_{r})\in\mathbb{Z}^{s-j}}}}\prod_{i=1}^{j}q_{i,+}^{-y_{i}/z+e_{i}}\prod_{i=j+1}^{r}q_{i,-}^{y_{i}/z+d_{i}}L_{\beta,d_{1},\ldots,d_{j},-e_{j+1},\ldots,-e_{r}}.$
Again by [Bro14], we see that $I_{X_{T,j}}(Q,d_{1},\ldots,d_{j},e_{j+1},\ldots,e_{r},-z)$ lies on the Lagrangian cone for $X_{T,j}$.
3.2.3. The non-abelian quotients
The abelian/non-abelian correspondence in Gromov–Witten theory is the principle that one can recover the Gromov–Witten theory of the non-abelian quotient from that of the associated abelian quotient. There is a long history of results in this direction, e.g. [BCFK05, BCFK08, CFKS08, Web23, Web21, CLS22]. The modern formulation of the expected relationship is, roughly, that the (genus zero) Gromov–Witten theory for the non-abelian quotient may be obtained from that of the associated abelian quotient by the following steps:
  1. (1)
    twisting the theory of the abelian quotient by the Euler class of the root bundle $\mathfrak{g}/\mathfrak{t}$;
  2. (2)
    specializing those Novikov parameters which are identified via the natural map
    $H_{2}(BT)\to H_{2}(BG).$
Let
$N_{\beta,\vec{d}}=\prod_{j=1}^{r}$ $\left(\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}\right)$
(3.2.5) $\cdot$ $\prod_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{r}\frac{\prod_{h=-\infty}^{d_{i}-d_{j}}(y_{i}-y_{j}+hz)}{\prod_{h=-\infty}^{0}(y_{i}-y_{j}+hz)}.$
Define $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,\vec{q}_{+},z)$ and $I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,\vec{q}_{-},z)$ to be
$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\ldots,-e_{r}}$
$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\ldots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\ldots,d_{r}}.$
Then, define
$I_{X_{\pm}}(Q,q_{\pm},z)=I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,\vec{q}_{+},z)|_{q_{1,\pm}=\cdots=q_{s,\pm}=q_{\pm}}.$
The following proposition is a corollary of the work of Coates–Lutz–Shafi in [CLS22] on the abelian/non-abelian correspondence. One could also prove it directly by generalizing the work of Oh in [Oh21] on $I$-functions of Grassmann bundles.
Proposition 3.9.
The function $I_{X_{\pm}}(Q,q_{\pm},-z)$ lies on the $\mathbb{T}$-equivariant Lagrangian cone $\mathcal{L}_{X_{\pm}}$.
Proof.
As described in [CLS22], $I_{X_{\pm}}$ is the Givental–Martin modification of $I_{X_{T,\pm}}$, so the proposition essentially follows from [CLS22, Theorem 1.11]. More precisely, we apply the argument of Theorem 5.11 of [CLS22], which provides an abelian/non-abelian correspondence for twisted theories over a flag bundle. Because that theorem is not stated in exactly the generality which we need here, we summarize the argument below and explain the necessary adjustments to the case at hand. We will prove the proposition for $X_{+}$.
The space $X_{+}$ is the total space of a vector bundle over the Grassmann bundle
$Z:=\operatorname{Gr}(r,E)=\operatorname{tot}\left(\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}G.$
The associated abelian quotient is the toric bundle given by the $r$-fold fiber product of $\mathbb{P}(E)$ over $B$:
$Z_{T}:=\operatorname{tot}\left(\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}T=\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E).$
Let $I_{Z_{T}}$ denote the Brown $I$-function for $Z_{T}$ as in [Bro14], and let $I_{Z_{T}}^{e_{\mathbb{T}}^{-1}((S_{T}\otimes p^{*}F^{\vee}))}$ denote the corresponding $((S_{T}\otimes p^{*}F^{\vee}),e^{-1}_{\mathbb{T}}(-))$-twisted $I$-function, where we use $S_{T}$ to denote the vector bundle on $Z_{T}$ induced by the right standard representation $R$ from (2.1.1) restricted to $T\subset G$. Then $I_{Z_{T}}^{e_{\mathbb{T}}^{-1}((S_{T}\otimes p^{*}F^{\vee}))}$ is exactly the $I$-function $I_{X_{T,+}}$ from the previous section.
Following the notation and discussion in Section 5.2 of [CLS22] (in particular Theorem 5.11), we apply a further twist to $I_{X_{T,+}}=I_{Z_{T}}^{e^{-1}_{\mathbb{T}}((S_{T}\otimes p^{*}F^{\vee})}$ by $(\Phi,\mathbf{c}^{\prime})$ where $\Phi\to Z_{T}$ is the root bundle with respect to $G$ and $\mathbf{c}^{\prime}$ is the $\mathbb{C}^{*}$-equivariant Euler class with respect to the $\mathbb{C}^{*}$-action scaling the fibers of $\Phi.$ Denote the equivariant parameter of this $\mathbb{C}^{*}$ action by $\lambda$. Specializing variables to $q_{1,+}=\cdots=q_{s,+}=q_{+},$ projecting via the map on Givental spaces defined in [CLS22]
$p:\mathcal{H}_{Z_{T}}^{W}\to\mathcal{H}_{Z},$
and taking the nonequivariant limit $\lambda\mapsto 0$, we obtain $I_{X_{+}}(Q,q_{+},z)$.
By the proof of [CLS22, Theorem 5.11], this function lies in the $((S_{+}\otimes p^{*}F^{\vee}),e^{-1}_{\mathbb{T}})$-twisted Lagrangian cone of $Z=\operatorname{Gr}(r,E)$. While the theorem in loc. cit. is only stated for the twisted theory $(L,e_{\mathbb{C}^{*}}(-))$, where $L$ is a line bundle and $e_{\mathbb{C}^{*}}(-)$ is an equivariant Euler class, by [CLS22, Remark 3.7] the result holds for other multiplicative classes as well. The further extension to the case that $L$ is replaced by the sum of line bundles also follows the same argument.
By definition, the $((S_{+}\otimes p^{*}F^{\vee}),e^{-1}_{\mathbb{T}})$-twisted Lagrangian cone of $\operatorname{Gr}(r,E)$ is equal to the $\mathbb{T}$-equivariant Lagrangian cone of $X_{+}$.
4. Analytic continuation: the abelian case
In this section we explicitly compute the analytic continuation of the $I$-functions of the associated abelian quotients $X_{T,\pm}$ and show that they are related by symplectic transformation. This was proven in [CIJ18] in the case $B$ is a point.
4.1. Projective bundles
Let us consider the special case of $r=1$ first. We will analytically continue the function $I_{X_{+}}$ to $q_{+}=\infty$, and show that the resulting function is related to $I_{X_{-}}$ by a symplectic transformation.
We begin by writing $I_{X_{\pm}}$ as $\psi(H_{X_{\pm}})$ where $\psi$ is as in (3.1.1). The relative Euler sequence for $X_{\pm}$ is
$0\to\mathcal{O}_{X_{\pm}}\to R\otimes p^{*}E\oplus R^{\vee}\otimes p^{*}F^{\vee}\to TX_{\pm}/B\to 0$
where $R$ is the vector bundle defined in Section 2.1. Because the Gamma class is multiplicative, we have $\hat{\Gamma}$, we have that
$\hat{\Gamma}_{X_{\pm}}=$ $\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)$
$=$ $\hat{\Gamma}_{B}\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y))\prod_{i=1}^{n}\Gamma(1+(z_{i}+y)).$
Note that the hypergeometric modification (3.2.3) is equal to
(4.1.1) $\frac{1}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}\prod_{j=1}^{n}\frac{\Gamma(1+(-x_{j}-y)/z)\Gamma(1+(z_{j}+y)/z)}{\Gamma(1+(-x_{j}-y)/z+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/z-\beta\cdot M_{j}+d)}$
Define $H^{B}(Q,z)$ to be
$H^{B}$ $:=(2\pi\mathbbm{i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}$
$=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}(TB\oplus E\oplus F^{\vee})$ and
$\mu_{X}(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\right)\phi.$
Define the $H^{*}_{\mathbb{T}}(X_{+})$-valued function
(4.1.2) $H_{X_{+},\beta}(q_{+})=$
$\sum_{e\in\mathbb{Z}}\frac{q_{+}^{-y/2\pi\mathbbm{i}+e}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi\mathbbm{i}+\beta\cdot L_{j}+e)\Gamma(1+(z_{j}+y)/2\pi\mathbbm{i}-\beta\cdot M_{j}-e)}.$
Similarly, define the $H^{*}_{\mathbb{T}}(X_{-})$-valued function
(4.1.3) $H_{X_{-},\beta}(q_{-})=$
$\sum_{d\in\mathbb{Z}}\frac{q_{-}^{y/2\pi\mathbbm{i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/2\pi\mathbbm{i}-\beta\cdot M_{j}+d)}.$
Then a simple check shows that
(4.1.4) $I_{X_{\pm}}(Q,q_{\pm},z)$ $=\psi\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{\pm},\beta}\right).$
We will fix $\beta$ and relate $H_{X_{+},\beta}$ and $H_{X_{-},\beta}$ via analytic continuation and symplectic transformation. Recall Definition 2.8 in the special case $r=1$. In this case $\mathbb{U}_{H}:H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{B_{k}|X_{-}})\in H^{*}_{\mathbb{T}}(B_{k}^{-})$ to
(4.1.5) $\sum_{l=1}^{n}e^{(n-1)\cdot(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2\mathbbm{i})}\phi_{kl}(\alpha)/e_{\mathbb{T}}(N_{B_{l}|X_{+}})$
where $\phi_{kl}:H^{*}_{\mathbb{T}}(B_{k}^{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(B_{l}^{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{k}^{+}:B_{k}^{+}\to B$ and $\pi_{l}^{-}:B_{l}^{-}\to B$. By restricting $H_{X_{+},\beta}$ to the fixed points of the $\mathbb{T}$-action and using Mellin–Barnes integrals, we may analytically continue $H_{X_{+},\beta}$ to $q_{+}=\infty.$ Because the method of computation is by now somewhat standard (see e.g. [CIJ18]), we relegate the proof to Appendix A. We obtain the following.
Proposition 4.1.
We have the equality
(4.1.6) $\mathbb{U}_{H}H_{X_{-},\beta}=\widetilde{H_{X_{+},\beta}}$
for all $\beta$, where $\widetilde{H_{X_{+},\beta}}$ denotes the analytic continuation of $H_{X_{+},\beta}$ along the path described above, under the substitution $q_{+}^{-1}=q_{-}$.
Recall Definition 3.5, $\mathbb{U}=\psi_{+}\circ\mathbb{U}_{H}\circ\psi_{-}^{-1}.$ We conclude the following in the case $r=1.$
Theorem 4.2.
The linear transformation $\mathbb{U}$ is symplectic, has a well-defined non-equivariant limit, and satisfies:
$\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}},$
where $\widetilde{I_{X_{+}}}$ denotes the analytic continuation of $I_{X_{+}}$ along the path $\gamma$ and we set $q_{+}^{-1}=q_{-}$. Furthermore, the following diagram commutes:
${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\mathbb{U}}$
Proof.
The proof of the equality is immediate from Definition 3.5, (4.1.4) and (4.1.6). The fact that $\mathbb{U}$ is symplectic and has a well-defined non-equivariant limit is Corollary 3.6. The commuting diagram follows from Proposition 2.9. ∎
4.2. Products of projective bundles
We next consider, for general $r>0$, the wall crossing between $I$-functions of the associated abelian quotients $X_{T,\pm}$. While the result follows simply by iterating Theorem 4.2 $r$ times, the computations in this section will play a role when we consider the wall crossing of the Grassmann flop.
To pass from $X_{T,-}$ to $X_{T,+}$ via variation of GIT, one must cross $r$ distinct walls, one for each factor in $T=(\mathbb{C}^{*})^{r}$. Recall the definition of $X_{T,j}$ from Section 2.2, and set $X_{T,-}=X_{T,0}$ and $X_{T,+}=X_{T,r}$. Observe that the wall crossing from $X_{T,j}$ to $X_{T,j+1}$ is a particular example of the wall crossing from the previous section.
In this case for any $j$ we have the relationship between Gamma functions:
$\hat{\Gamma}(TX_{T,j})=\hat{\Gamma}_{B}\hat{\Gamma}(TX_{T}/B)$
with
$\hat{\Gamma}(TX_{T}/B)=\prod_{k=1}^{r}\prod_{j=1}^{n}\left(\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\right).$
Define $H^{B}_{T}(Q,z)$ to be
$H^{B}_{T}$ $:=(2\pi\mathbbm{i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X_{T}}}z^{\mu_{X_{T}}}J^{B}$
$=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,T,\beta},$
where $\rho_{X_{T}}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}(TB\oplus(E\oplus F^{\vee})^{\oplus r})$ and
$\mu_{X_{T}}(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X_{T,\pm})}{2}\right)\phi.$
Define the $H^{*}_{\mathbb{T}}(X_{T,+})$-valued, resp. $H^{*}_{\mathbb{T}}(X_{T,-})$-valued, functions
$H_{X_{T,+},\beta}=$
$\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{q_{k,+}^{(-y_{k}/2\pi\mathbbm{i}+e_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})}$
$H_{X_{T,-},\beta}=$
$\sum_{(d_{1},\ldots,d_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{q_{k,-}^{(y_{k}/2\pi\mathbbm{i}+d_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}+d_{k})}$
Then as in the case of $r=1$, we have
(4.2.1) $I_{X_{T,\pm}}(Q,q_{\pm},z)$ $=\psi_{X_{T,\pm}}\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}H_{B,T,\beta}H_{X_{T,\pm},\beta}\right)$
Recall that the $\mathbb{T}$-fixed loci of $X_{T,\pm}$ are indexed by $F_{\pm}$.
Let $\mathbb{U}_{H}^{T}:H^{*}_{\mathbb{T}}(X_{T,-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{T,+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined by sending $\alpha/e_{\mathbb{T}}(N_{f^{-}})\in H^{*}_{\mathbb{T}}(B_{f^{-}})$ to
$\sum_{f^{+}\in F_{+}}\prod_{i=1}^{r}e^{(n-1)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+x_{f^{+}_{i}})/2\mathbbm{i})}{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+z_{f^{-}_{i}})/2\mathbbm{i})}\frac{\phi_{f^{+},f^{-}}(\alpha)}{e_{\mathbb{T}}(N_{f^{+}})}$
where $\phi_{f^{+},f^{-}}:H^{*}_{\mathbb{T}}(B_{f^{-}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(B_{f^{+}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{f^{+}}:B_{f^{+}}\to B$ and $\pi_{f^{-}}:B_{f^{-}}\to B$, and $N_{f^{\pm}}$ is the normal bundle $N_{B_{f^{\pm}}|X_{\pm}}$.
A straightforward generalization of the analytic continuation in Appendix A (essentially repeating it $r$ times, one for each $q_{i}$ variable) yields the following:
$\mathbb{U}_{H}^{T}H_{X_{T,-},\beta}=\widetilde{H_{X_{T,+},\beta}}$
for all $\beta$, after identifying $q_{k,+}^{-1}=q_{k,-}$. Here the path of analytic continuation is a concatenation
$\gamma_{1}^{T}\star\cdots\star\gamma_{r}^{T}$
where $\gamma_{k}^{T}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$, and in the $\log(q_{k,+})$ plane follows the path $\gamma$ of analytic continuation from the previous section.
Define
$\mathbb{U}^{T}=\psi_{X_{T,+}}\circ\mathbb{U}_{H}^{T}\circ\psi_{X_{T,-}}^{-1}$
in analogy to the previous section. By Theorem 4.2, we obtain:
Theorem 4.3.
The linear transformation $\mathbb{U}^{T}$ is symplectic, has a well-defined non-equivariant limit, and satisfies:
$\mathbb{U}^{T}I_{X_{T,-}}=\widetilde{I_{X_{T,+}}}.$
5. Grassmann flops
In this section we compute the analytic continuation and symplectic transformation for the $I$-functions $X_{+}$ and $X_{-}$ appearing in the Grassmann flop.
5.1. Analytic continuation
The computation proceeds by first computing the analytic continuation for the twisted $I$-functions of the abelian quotients, similar to the previous section, and then deforming the path to the locus $q_{1,+}=\cdots=q_{r,+}$.
Let $S_{-}$ denote the rank-$r$ tautological bundle on $X_{-}$. We have the exact sequence
$0\to S_{-}\otimes S_{-}^{\vee}\to S_{-}\otimes p^{*}E\oplus S_{-}^{\vee}\otimes p^{*}F^{\vee}\to TX_{-}/B\to 0$
and similarly for $X_{+}$. We view $H^{*}_{\mathbb{T}}(X_{\pm})$ as a subspace of $(H^{*}_{\mathbb{T}}(V^{s}(G)/T))$ via Proposition 2.3. After pulling back to $V^{s}(G)/T$, we may write the Gamma class $\hat{\Gamma}_{X_{\pm}}$ as $\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)$ where in this case
$\hat{\Gamma}(TX_{\pm}/B)=\frac{\prod_{k=1}^{r}\prod_{j=1}^{n}\left(\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\right)}{\hat{\Gamma}(\mathfrak{g}/\mathfrak{t})}$
with
$\hat{\Gamma}(\mathfrak{g}/\mathfrak{t})=\prod_{i\neq k}\hat{\Gamma}(1+y_{k}-y_{i}).$
Recall that the $I$-functions for $X_{\pm}$ are specializations at $q_{1,\pm}=\cdots=q_{r,\pm}$ of $I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ where
$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\ldots,-e_{r}}$
$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\ldots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\ldots,d_{r}}.$
with $N_{\beta,\vec{d}}$ defined in (3.2.5).
The hypergeometric factor $N_{\beta,\vec{d}}$ can be rewritten as
$\frac{1}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}\prod_{i\neq k}\frac{\Gamma(1+(y_{k}-y_{i})/z+d_{k}-d_{i})}{\Gamma(1+(y_{k}-y_{i})/z)}\cdot$
$\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{\hat{\Gamma}(1+(-y_{k}-x_{j})/z)\Gamma(1+(y_{k}+z_{j})/z)}{\Gamma(1+(-y_{k}-x_{j})/z+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/z-\beta\cdot M_{j}+d_{k})},$
Generalizing the definition of $H^{B}(Q,z)$ to the case of $r>1$, we define
$H^{B}$ $:=(2\pi\mathbbm{i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}$
$=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}(TB\oplus(E\oplus F^{\vee})^{\oplus r})$ and $\mu_{X}$ acts on a homogeneous element $\phi$ by
$\mu_{X}(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\right)\phi.$
Denote by $H_{X_{+},\beta}$ the function
$\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\left(q_{k,+}^{(-y_{k}/2\pi\mathbbm{i}+e_{k})}\prod_{i<k}\Gamma(1+(y_{k}-y_{i})/2\pi\mathbbm{i}-e_{k}+e_{i})\Gamma(1-(y_{k}-y_{i})/2\pi\mathbbm{i}+e_{k}-e_{i})\right.$
$\left.\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})^{-1}\right)$
$=$ $\frac{\pi^{r(r-1)}}{\prod_{i<k\leq r}\sin((y_{k}-y_{i})/2\mathbbm{i})}\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}$
$\prod_{k=1}^{r}\left(q_{k,+}^{(-y_{k}/2\pi\mathbbm{i}+e_{k})}\prod_{i<k}((y_{k}-y_{i})/2\pi\mathbbm{i}-e_{k}+e_{i})\right.$
$\left.\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})^{-1}\right)$
As before we have that
(5.1.1) $I_{X_{+}}(Q,q_{+},z)$ $=\psi_{+}\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{+},\beta}\right)\vline_{q_{1,+}=\cdots=q_{r,+}=q_{+}}.$
An analogous statement holds for $X_{-}$.
Restrict $H_{X_{+},\beta}$ to a fixed locus $B_{\delta}$, where $\delta\in D_{+}$ is a size $r$ subset of $\{1,\ldots,n\}$. Then
(5.1.2) $H^{+}_{\beta,\delta}:=H_{X_{+},\beta}|_{B_{\delta}}=\frac{\pi^{r(r-1)}}{\prod_{i<k\leq r}\sin((x_{\delta_{i}}-x_{\delta_{k}})/2\mathbbm{i})}K^{+}_{\beta,\delta}$
where $K^{+}_{\beta,\delta}$ is
$\sum_{e_{1}\geq-\beta\cdot L_{\delta_{1}}}\cdots\sum_{e_{r}\geq-\beta\cdot L_{\delta_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}\prod_{k=1}^{r}\left(q_{k,+}^{(x_{\delta_{k}}/2\pi\mathbbm{i}+e_{k})}\prod_{i<k}((-x_{\delta_{k}}+x_{\delta_{i}})/2\pi\mathbbm{i}-e_{k}+e_{i})\right.$
$\left.\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})^{-1}\right).$
For any $f\in F_{-}$, define $K^{-}_{\beta,f}$ to be
$\sum_{d_{1}\geq\beta\cdot M_{f_{1}}}\cdots\sum_{d_{r}\geq\beta\cdot M_{f_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}d_{k}}\prod_{k=1}^{r}\left(q_{k,-}^{(-z_{f_{k}}/2\pi\mathbbm{i}+d_{k})}\prod_{i<k}((-z_{f_{k}}+z_{f_{i}})/2\pi\mathbbm{i}+d_{k}-d_{i})\right.$
$\left.\prod_{j=1}^{n}\Gamma(1+(z_{f_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d_{k})^{-1}\Gamma(1+(-z_{f_{k}}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}+d_{k})^{-1}\right).$
Then for $\delta^{-}\in D_{-}$, after choosing an ordering on the elements and viewing $\delta^{-}$ as a function, we have
(5.1.3) $H^{-}_{\beta,\delta^{-}}:=H_{X_{-},\beta}|_{B_{\delta^{-}}}=\frac{\pi^{r(r-1)}}{\prod_{i<k\leq r}\sin((z_{\delta^{-}_{i}}-z_{\delta^{-}_{k}})/2\mathbbm{i})}K^{-}_{\beta,\delta^{-}}$
The computation of the analytic continuation of $K_{\beta,\delta}^{+}$ is almost identical to that of $H_{\beta,T,f^{+}}$ from the previous section. In this case the path of analytic continuation will be given by $\gamma_{1}\star\cdots\star\gamma_{r}$ where $\gamma_{k}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$ and in the $\log(q_{k,+})$ plane, will go through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi\mathbbm{i}$ when $\mathfrak{R}(\log(q_{k,+}))$ is close to 0. See Appendix A for details.
Let
$C^{K}_{f^{-},f^{+}}=\prod_{i=1}^{r}e^{(n-r)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+x_{f^{+}_{i}})/2\mathbbm{i})}{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+z_{f^{-}_{i}})/2\mathbbm{i})}$
Generalizing the computation from Appendix A, we see that the analytic continuation of $K^{+}_{\beta,\delta^{+}}$ along the path $\gamma_{1}\star\cdots\star\gamma_{r}$ is given by
(5.1.4) $\sum_{f^{-}\in F_{-}}C^{K}_{f^{-},\delta^{+}}K^{-}_{\beta,f^{-}},$
after identifying $q_{k,+}^{-1}=q_{k,-}$. We emphasize that here the sum is over all functions $f^{-}:\{1,\ldots,r\}\to\{1,\ldots,n\}$.
Then by (5.1.2) and (5.1.3), if we define
$C^{H}_{f^{-},\delta^{+}}=C^{K}_{f^{-},\delta^{+}}\prod_{i<k\leq r}\frac{\sin((z_{f^{-}_{i}}-z_{f^{-}_{k}})/2\mathbbm{i})}{\sin((x_{\delta^{+}_{i}}-x_{\delta^{+}_{k}})/2\mathbbm{i})},$
the analytic continuation of $H^{+}_{\beta,\delta^{+}}$ is given by
$\widetilde{H^{+}_{\beta,\delta^{+}}}=\sum_{f^{-}\in F_{-}}C^{H}_{f^{-},\delta^{+}}H^{-}_{\beta,f^{-}}.$
Note that the numerator of $C^{H}_{f^{-},\delta^{+}}$ vanishes unless $f^{-}_{i}\neq f^{-}_{k}$ for $i\neq k$. Therefore we can replace the summation above to be only over all injective functions $\delta^{-}\in In_{-}$:
(5.1.5) $\widetilde{H^{+}_{\beta,\delta^{+}}}=\sum_{\delta^{-}\in In_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}.$
In order to connect the above computation to the actual $I$-functions for $X_{+}$ and $X_{-}$ via (5.1.1), we must homotope the path $\gamma_{1}\star\cdots\star\gamma_{r}$ of analytic continuation to a new path which lies entirely within the diagonal $q_{1,+}=\cdots=q_{r,+}$.
Proposition 5.1.
The path of analytic continuation $\gamma_{1}\star\cdots\star\gamma_{r}$ used in (5.1.5) is homotopic to a path $\hat{\gamma}$ contained entirely in the locus $q_{1,+}=\cdots=q_{r,+}$, via a homotopy which is
  1. (1)
    contained within the domain for which ${H^{+}_{\beta,\delta^{+}}}$ is analytic, and
  2. (2)
    independent of $\beta$ and $\delta^{+}$.
After specializing the Novikov parameters $q_{1,\pm}=\cdots=q_{r,\pm}=q_{\pm}$, we can therefore analytically continue $H^{+}_{\beta,\delta^{+}}$ from $q_{+}=0$ to $q_{+}=\infty$. Upon setting $q_{+}^{-1}=q_{-}$, we obtain the following:
(5.1.6) $\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in In_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}.$
Proof.
First consider the function $f_{\beta,\delta,k}(q_{k,+})$ defined as
$\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi\mathbbm{i}+e_{k})}(-1)^{(r-1)e_{k}}}{\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})}$
$=$ $\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\left(\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi\mathbbm{i}+e_{k})}(-1)^{(n+r-1)e_{k}+\sum_{j=1}^{n}\beta\cdot M_{j}}\prod_{j=1}^{n}\sin((x_{\delta_{k}}-z_{j})/2\mathbbm{i})}{\pi^{n}}\right.$
$\cdot\left.\prod_{j=1}^{n}\frac{\Gamma((x_{\delta_{k}}-z_{j})/2\pi\mathbbm{i}+\beta\cdot M_{j}+e_{k})}{\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})}\right)$
This satisfies the differential equation
$\left[\prod_{j=1}^{n}(\partial_{k}-x_{j}/2\pi\mathbbm{i}+\beta\cdot L_{j})-q_{k,+}(-1)^{n-r+1}\prod_{j=1}^{n}(\partial_{k}-z_{j}/2\pi\mathbbm{i}+\beta\cdot M_{j})\right]\Phi(q_{k,+})=0,$
where $\partial_{k}=q_{k,+}\frac{\partial}{\partial q_{k,+}}$. This differential equation has singularities at $0,\infty,$ and $(-1)^{n-r+1}$. Following (A.0.3), our path of analytic continuation in the $\log(q_{k,+})$ plane goes through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi\mathbbm{i}$ when $\mathfrak{Re}(\log(q_{k,+}))$ is close to 0.
Consider the product
$\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})$
viewed as an analytic function of the variables $\log(q_{1,+}),\ldots,\log(q_{r,+})$. We see that it may be analytically continued everywhere except along the poles:
(5.1.7) $\log(q_{k,+})=(n-r+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}.$
Define the differential operator
$\hat{\Delta}=\prod_{k=1}^{r}\left(\prod_{i<k}(-\partial_{k}+\partial_{i})\right).$
Then
$K^{+}_{\beta,\delta}=\hat{\Delta}\left(\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})\right),$
and therefore $K^{+}_{\beta,\delta}$ has the same poles at (5.1.7).
[…]
Figure 1.
Let $\gamma_{1}$ denote the path that keeps $\log(q_{2,+}),\ldots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ from Figure 1 in $\log(q_{1,+})$. Let $\gamma_{2}$ denote the path that keeps $\log(q_{1,+})$ constant (equal to $\infty$), keeps $\log(q_{3,+}),\ldots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ in $\log(q_{2,+})$. Define $\gamma_{k}$ for $k\leq r$ similarly. The analytic continuation for $K^{+}_{\beta,\delta}$ computed in (5.1.4) involved the concatenated path
$\gamma_{1}\star\cdots\star\gamma_{r}.$
We must show that this composition of paths is homotopy equivalent to a path $\hat{\gamma}$ which follows $\gamma$ in the plane $\log(q_{1,+})=\cdots=\log(q_{r,+})$ via a homotopy which is contained in
$\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}\}_{k=1}^{r}.$
We verify this in the case $r=2$. An induction argument proves the claim for general $r$. First, define a path $\widetilde{\gamma}_{1}$ which agrees with $\gamma_{1}$ except that
  • for all points on the path such that $\mathfrak{Re}(\log(q_{1,+}))\geq 0$ we let $\mathfrak{Im}(\log(q_{1,+}))=(n-r)\pi\mathbbm{i}$;
  • while the value of $\mathfrak{Re}(\log(q_{2,+}))$ remains constant, the value of $\mathfrak{Im}(\log(q_{2,+}))$ is defined to be equal to $\mathfrak{Im}(\log(q_{1,+}))$ for all points on the path.
Similarly, define a path $\widetilde{\gamma}_{2}$ which agrees with $\gamma_{2}$ except that
  • for all points on the path such that $\mathfrak{Re}(\log(q_{2,+}))\leq 0$ we let $\mathfrak{Im}(\log(q_{2,+}))=(n-r)\pi\mathbbm{i}$;
  • while the value of $\mathfrak{Re}(\log(q_{1,+}))$ remains constant, the value of $\mathfrak{Im}(\log(q_{1,+}))$ is defined to be equal to $\mathfrak{Im}(\log(q_{2,+}))$ for all points on the path.
See Figure 2.
[…]
Figure 2. The projections of the paths $\widetilde{\gamma}_{1}$ and $\widetilde{\gamma}_{2}$ to the $\log(q_{1,+})$ and $\log(q_{2,+})$ planes respectively.
It is clear that $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ is homotopic to $\gamma_{1}\star\gamma_{2}$ in $\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}\}_{k=1}^{r}.$ By the construction, the imaginary parts $\mathfrak{Im}(\log(q_{1,+}))$ and $\mathfrak{Im}(\log(q_{2,+}))$ are equal for all points on the path $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$.
Finally, we homotope $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ further by modifying the real part of the path so that $\mathfrak{Re}(\log(q_{1,+}))=\mathfrak{Re}(\log(q_{2,+}))$ for all points on the path (see Figure 3). This second homotopy is also contained entirely in $\mathbb{C}^{2}\setminus\{\log(q_{k,+})=(n-2+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}\}_{k=1}^{2}.$
The resulting path $\hat{\gamma}$ from $(\log(q_{1,+}),\log(q_{2,+}))=(-\infty,-\infty)$ to (,)($\infty$,$\infty$)( ∞ , ∞ ) lies entirely in the plane $\log(q_{1,+})=\log(q_{2,+})$ as desired.
$\mathfrak{R}e_{\mathbb{T}}(\log(q_{1,+}))$$\mathfrak{R}e_{\mathbb{T}}(\log(q_{2,+}))$(,($\infty$,$\infty$)( - ∞ , - ∞ )(,($\infty$,$\infty$)( ∞ , ∞ )
Figure 3. The (real part of the) original path of analytic continuation $\gamma_{1}\star\gamma_{2}$ (in purple) and the new path $\hat{\gamma}$ (in blue).
5.2. Simplification
After specializing the Novikov parameters, we observe that $H^{{\pm}}_{\beta,\delta}|_{q_{1,\pm}=\cdots=q_{r,\pm}}=H^{{\pm}}_{\beta,\delta^{\prime}}|_{q_{1,\pm}=\cdots=q_{r,\pm}}$ whenever $\delta$ and $\delta^{\prime}$ are equal up to a permutation. This allows for a dramatic simplification to the formula (5.1.6). The simplification relies on the following identity.
Lemma 5.2.
We have the following identity of (bi)-anti-symmetric functions:
(5.2.1) $\prod_{i=1}^{r}\prod_{l<i}\left((z_{i}-z_{l})(x_{l}-x_{i})\right)=\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l}).$
Proof.
Denote the left and right hand sides of (5.2.1) by $L(\underline{x},\underline{z})$ and $R(\underline{x},\underline{z})$ respectively.
We first claim that $R(\underline{x},\underline{z})$ is divisible by $(z_{i}-z_{l})$ for $i\neq l$. Consider the substitution $z_{i}=z_{l}=z$. Group the elements $\sigma\in S_{r}$ in pairs by grouping $\sigma$ with $\tilde{\sigma}=\sigma\circ\tau_{il}$ where $\tau_{il}$ is the transposition $(il)$. The elements $\tilde{\sigma}$ and $\sigma$ have opposite signs. Since $\sigma(l)=\tilde{\sigma}(i)$ and $\sigma(i)=\tilde{\sigma}(l)$, we see that
$\left(\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l})+\operatorname{sgn}(\tilde{\sigma})\prod_{l=1}^{r}\prod_{j\neq\tilde{\sigma}(l)}(x_{j}-z_{l})\right)|_{z_{i}=z_{l}=z}=0,$
and the claim follows.
On the other hand, $R(\underline{x},\underline{z})$ may also be written as
$\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{j=1}^{r}\prod_{l\neq\sigma(j)}(x_{j}-z_{l}),$
from which one can apply the same argument to conclude that $R(\underline{x},\underline{z})$ is divisible by $(x_{i}-x_{l})$ for $i\neq l$.
Because $R(\underline{x},\underline{z})$ is a polynomial of degree $r(r-1)$, the above divisibility claims imply that it is equal to a constant times $L(\underline{x},\underline{z})$. We consider the coefficient of the monomial $\prod_{i=1}^{r}(z_{i}x_{i})^{i-1}$. A simple induction argument shows that in $R(\underline{x},\underline{z})$, this monomial only appears in the summand indexed by $\sigma=id$, with a coefficient of $(-1)^{r(r-1)/2}.$ In $L(\underline{x},\underline{z})$ the monomial arises from the product
$\prod_{i=1}^{r}\prod_{l<i}(-z_{l}\cdot x_{l}).$
We see that the coefficients match, thus the two functions agree. ∎
Corollary 5.3.
For $\delta^{-}:\{1,\ldots,r\}\to\{1,\ldots,n\}$ an injective function, let $S_{r}\cdot\delta^{-}$ denote the set of all possible functions obtained by pre-composing $\delta^{-}$ with a permutation of $\{1,\ldots,r\}.$ The identity
(5.2.2) $\sum_{f^{-}\in S_{r}\cdot\delta^{-}}C^{H}_{f^{-},\delta^{+}}=C_{\delta^{-},\delta^{+}}.$
holds, where $C_{\delta^{-},\delta^{+}}$ was defined in (2.3.4).
Proof.
By reindexing the sets $\{x_{1},\ldots,x_{n}\}$ and $\{z_{1},\ldots,z_{n}\}$, we may assume without loss of generality that $\delta^{-}=\delta^{+}=\{1,\ldots,r\}$. Then in this case (5.2.2) simplifies to the equation
(5.2.3) $\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\left(\prod_{\stackrel{{\scriptstyle 1\leq j\leq n}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2\mathbbm{i})}{\sin((z_{\sigma(k)}-z_{j})/2\mathbbm{i})}\prod_{i<k}\frac{\sin((z_{\sigma(k)}-z_{\sigma(i)})/2\mathbbm{i})}{\sin((x_{k}-x_{i})/2\mathbbm{i})}\right)$
$=$ $\prod_{k=1}^{r}\prod_{j>r}\frac{\sin((x_{k}-z_{j})/2\mathbbm{i})}{\sin((z_{k}-z_{j})/2\mathbbm{i})}$
or, equivalently,
(5.2.4) $\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\left(\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2\mathbbm{i})}{\sin((z_{\sigma(k)}-z_{j})/2\mathbbm{i})}\prod_{i<k}\frac{\sin((z_{\sigma(k)}-z_{\sigma(i)})/2\mathbbm{i})}{\sin((x_{k}-x_{i})/2\mathbbm{i})}\right)=1.$
Observe that
$\prod_{k=1}^{r}\left(\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{1}{\sin((z_{\sigma(k)}-z_{j})/2\mathbbm{i})}\prod_{i<k}\sin((z_{\sigma(k)}-z_{\sigma(i)})/2\mathbbm{i})\right)$
$=$ $\prod_{k=1}^{r}\prod_{j>k}\frac{1}{\sin((z_{\sigma(k)}-z_{\sigma(j)})/2\mathbbm{i})}$
$=$ $\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{j<k}\frac{1}{\sin((z_{j}-z_{k})/2\mathbbm{i})}$
and (5.2.4) is then equivalent to the equality
(5.2.5) $\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\sin((x_{k}-z_{j})/2\mathbbm{i})=\prod_{k=1}^{r}\prod_{i<k}\sin((z_{i}-z_{k})/2\mathbbm{i})\sin((x_{k}-x_{i})/2\mathbbm{i})$
Noting that
$\sin((a-b)/2\mathbbm{i})=\frac{e^{a/2+b/2}(e^{a}-e^{b})}{2\mathbbm{i}},$
under the change of variables $X_{i}=e^{x_{i}},Z_{i}=e^{z_{i}}$, (5.2.5) becomes
(5.2.6) $\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}(X_{k}-Z_{j})=\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\prod_{i<k}(Z_{i}-Z_{k})(X_{k}-X_{i}).$
Lemma 5.2 then gives the desired equality.
5.3. Main theorem
Combining (5.1.6) with (5.2.2), we conclude that
(5.3.1) $\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in D_{-}}C_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}.$
where the path of analytic continuation is along $\hat{\gamma}$ from Proposition 5.1 and the sum is now over all unordered size $r$ subsets of $\{1,\ldots,n\}$.
Let $\mathbb{U}_{H}:H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined in (2.3.5). Then (5.3.1) implies that:
(5.3.2) $\mathbb{U}_{H}H_{X_{-},\beta}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}=\widetilde{H_{X_{+},\beta}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}$
for all $\beta$, where the path of analytic continuation is along $\hat{\gamma}$ and we set $q_{+}^{-1}=q_{-}$.
Theorem 5.4.
The linear transformation $\mathbb{U}$ is symplectic, has a well-defined non-equivariant limit, and satisfies:
$\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}},$
where $\widetilde{I_{X_{+}}}$ denotes the analytic continuation of $I_{X_{+}}$ along the path $\hat{\gamma}$ and we set $q_{+}^{-1}=q_{-}$. Furthermore, the following diagram commutes:
${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\mathbb{U}}$
Proof.
Equation (5.1.1) and the analogous statement for $X_{-}$ together with (5.3.2) immediately implies that $\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}}.$ The fact that $\mathbb{U}$ is symplectic and has a well-defined non-equivariant limit is Corollary 3.6. The commuting diagram the follows from Proposition 2.9.
Remark 5.5 (Big $I$-functions and the comparison with [LSW]).
In the special case that $B$ is a point, the wall crossing from $X_{+}$ to $X_{-}$ is the standard Grassmann flop. In this case, by Webb [Web23, Section 6.3], one can obtain big $I$-functions of $X_{\pm}$, denoted $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$, as a modification of $I_{X_{\pm}}(q_{\pm},z)$. The respective big $I$-functions generate the entire Lagrangian cones $\mathcal{L}_{X_{\pm}}$, and therefore fully determine the genus-zero Gromov–Witten theory of $X_{\pm}$. A straightforward generalization of the computations above shows that the transformation $\mathbb{U}$ from Theorem 5.4 also identifies these big $I$-functions:
(5.3.3) $\mathbb{U}\mathbb{I}_{X_{-}}(q_{-},\mathbf{x},z)=\widetilde{\mathbb{I}_{X_{+}}}(q_{+},\mathbf{x},z),$
where $\widetilde{\mathbb{I}_{X_{+}}}$ denotes the analytic continuation in $q_{+}$ as in Theorem 5.4.
In the case that $B$ is a point, Lutz–Shafi–Webb [LSW] have obtained a different proof that there exists a symplectic transformation, which we will denote here by $\mathbb{U}^{LSW}$, identifying the big $I$-functions of $X_{\pm}$ after the analytic continuation. Rather than comparing to the Fourier–Mukai transform $\mathbb{FM}$ however, the symplectic transformation in $\mathbb{U}^{LSW}$ is obtained directly from wall crossing of the associated abelian GIT quotients $X_{T,\pm}$. Their result is related to Theorem 5.4 as follows. Let $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$ be the big $I$-functions of the previous paragraph, and let $g^{*}\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)$ denote the $I$-functions used in [LSW]. Comparing explicit formulas, we have
(5.3.4) $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)|_{q_{\pm}=\mathbf{y}_{\pm}}=\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)|_{\mathbf{q}=1},$
Under the identification $q_{\pm}\mapsto\mathbf{y}_{\pm}$, the path $\delta$ of analytic continuation used in loc. cit. matches that given in Proposition 5.1. In [LSW] the abelian/non-abelian correspondence is used to show that there is a unique linear map $\mathbb{U}^{LSW}$ satisfying
$\mathbb{U}^{LSW}\mathbb{I}^{-}_{G}(\mathbf{y}_{-}\mathbf{x},z)|_{\mathbf{q}=1}=\widetilde{\mathbb{I}^{+}_{G}}(\mathbf{y}_{+}\mathbf{x},z)|_{\mathbf{q}=1}.$
From this, (5.3.4), and (5.3.3), we conclude that $\mathbb{U}=\mathbb{U}^{LSW}$, thus providing a consistency check between the two results.
5.4. Central charge
Motivated by the notion of central charge as given in e.g. [KA23, Iri09], we make a similar definition. The above wall crossing result takes an especially simple form using this language.
Definition 5.6.
For $E\in K^{0}_{\mathbb{T}}(X_{\pm})$, define the relative quasimap central charge
$Z_{X_{\pm}}(E)(Q,q_{\pm},z):=\left\langle I_{X_{\pm}}(Q,q_{\pm},-z),\Psi_{\pm}(E)\right\rangle_{X_{\pm}}.$
With this definition, we obtain as an immediate corollary that the central charge of an element $E$ and of its Fourier–Mukai transform are related by analytic continuation.
Corollary 5.7.
For $E\in K^{0}_{\mathbb{T}}(X_{\pm})$, the analytic continuation of $Z_{X_{+}}(\mathbb{FM}(E))(Q,q_{+},z)$ to $q_{+}=\infty$ is $Z_{X_{-}}(E)(Q,q_{-},z)$.
Proof.
We use the diagram in Theorem 4.2 and the fact that $\mathbb{U}$ preserves the symplectic pairing.
$\widetilde{Z_{X_{+}}}(\mathbb{FM}(E))(Q,q_{+},z)=$ $\left\langle\widetilde{I_{X_{+}}}(Q,q_{+},-z),\Psi_{+}(\mathbb{FM}(E))\right\rangle_{X_{+}}$
$=$ $\left\langle\mathbb{U}I_{X_{-}}(Q,q_{-},-z),\mathbb{U}\Psi_{-}(E)\right\rangle_{X_{+}}$
$=$ $\left\langle I_{X_{-}}(Q,q_{-},-z),\Psi_{-}(E)\right\rangle_{X_{-}}$
$=$ $Z_{X_{-}}(E)(Q,q_{-},z).$
Appendix A Proof of Proposition 4.1
Using the identity
(A.0.1) $\Gamma(1-z)\Gamma(z)=\frac{\pi}{\sin(\pi z)},$
the denominators of $H_{X_{+},\beta}$ may be rewritten as
$\frac{\prod_{i=1}^{n}\sin((-z_{i}-y)/2\mathbbm{i}))}{\pi^{n}}\frac{\prod_{i=1}^{n}(-1)^{\beta\cdot M_{i}+e}\Gamma((-z_{i}-y)/2\pi\mathbbm{i}+\beta\cdot M_{i}+e)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi\mathbbm{i}+\beta\cdot L_{j}+e)}.$
Let $B_{l}^{+}$ denote the $l$th fixed locus, isomorphic to $B$. By Corollary 2.2, the restriction of the relative hyperplane class $-y$ to $B_{l}^{+}$ is $x_{l}$. Thus $H_{\beta,l}^{+}:=H_{X_{+},\beta}|_{B_{l}^{+}}$ is given by
(A.0.2) $\sum_{e\geq-\beta\cdot L_{l}}q_{+}^{e+x_{l}/2\pi\mathbbm{i}}\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi^{n}}\frac{\prod_{i=1}^{n}(-1)^{\beta\cdot M_{i}+e}\Gamma((-z_{i}+x_{l})/2\pi\mathbbm{i}+\beta\cdot M_{i}+e)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+x_{l})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e)}$
This satisfies the differential equation
$\left[\prod_{j=1}^{n}(\partial_{q}-x_{j}/2\pi\mathbbm{i}+\beta\cdot L_{j})-q_{+}(-1)^{n}\prod_{j=1}^{n}(\partial_{q}-z_{j}/2\pi\mathbbm{i}+\beta\cdot M_{j})\right]\Phi(q_{+})=0,$
where $\partial_{q}=q_{+}\frac{\partial}{\partial q_{+}}$. This differential equation has singularities at $0,\infty,$ and $(-1)^{n}$. Following [CIJ18], our path $\gamma$ of analytic continuation in the $\log(q_{+})$ will go through $\mathfrak{Im}(\log(q_{+}))=(n-1)\pi\mathbbm{i}$ when $\mathfrak{R}(\log(q_{+}))$ is close to 0. See Figure 1 for a picture of the path. We will use the Mellin–Barnes method to compute this analytic continuation of $H_{\beta,l}^{+}$ to $q_{+}=\infty$.
We may rewrite (A.0.2) as a contour integral
(A.0.3) $\frac{1}{2\pi\mathbbm{i}}\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi^{n}}\int_{C}\Gamma(s)\Gamma(1-s)q_{+}^{s+x_{l}/2\pi\mathbbm{i}}\cdot$
$e^{-\pi\mathbbm{i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot s)}\frac{\prod_{i=1}^{n}\Gamma((-z_{i}+x_{l})/2\pi\mathbbm{i}+\beta\cdot M_{i}+s)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+x_{l})/2\pi\mathbbm{i}+\beta\cdot L_{j}+s)},$
where $C$ encloses all integers $e\geq-\beta\cdot L_{l}$.
Closing the curve on the other side, we obtain residues at integers $d<-\beta\cdot L_{l}$ and at
$-(-z_{k}+x_{l})/2\pi\mathbbm{i}-d\hskip 14.22636pt1\leq k\leq n,d\in\mathbb{Z}_{\geq\beta\cdot M_{k}}.$
The residues at integers $d\leq-\beta\cdot L_{l}$ vanish. The integral is then equal to minus the sum of the remaining residues:
(A.0.4) $-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi\mathbbm{i}-d)\Gamma(1+(-z_{k}+x_{l})/2\pi\mathbbm{i}+d)q_{+}^{z_{k}/2\pi\mathbbm{i}-d}$
$\cdot$ $\frac{e^{-\pi\mathbbm{i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot(-(-z_{k}+x_{l})/2\pi\mathbbm{i}-d)+d-\beta\cdot M_{k})}}{(d-\beta\cdot M_{k})!}\frac{\prod_{i\neq k}\Gamma((-z_{i}+z_{k})/2\pi\mathbbm{i}+\beta\cdot M_{i}-d)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)}$
$\cdot$ $\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi^{n}}$
$=$ $-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi\mathbbm{i})\Gamma(1+(-z_{k}+x_{l})/2\pi\mathbbm{i})q_{+}^{z_{k}/2\pi\mathbbm{i}-d}$
$\cdot$ $\frac{e^{(n-1)(x_{l}-z_{k})/2}}{(d-\beta\cdot M_{k})!\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\prod_{i\neq k}\Gamma(1-(-z_{i}+z_{k})/2\pi\mathbbm{i}-\beta\cdot M_{i}+d)}$
$\cdot$ $\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi\prod_{i\neq k}\sin((-z_{i}+z_{k})/2\mathbbm{i})}$
$=$ $\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}e^{(n-1)(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2\mathbbm{i})}$
$\cdot$ $\frac{q_{+}^{z_{k}/2\pi\mathbbm{i}-d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\prod_{i=1}^{n}\Gamma(1-(-z_{i}+z_{k})/2\pi\mathbbm{i}-\beta\cdot M_{i}+d)}$
where we have again simplified using the Gamma function identity (A.0.1).
Let $B_{k}^{-}$ denote the $k$th fixed locus of $X_{-}$. By (4.1.3), observe that $H_{\beta,k}^{-}:=H_{X_{-},\beta}|_{B_{k}^{-}}$ is given by
$\sum_{d\geq\beta\cdot M_{k}}\frac{q_{-}^{-z_{k}/2\pi\mathbbm{i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}-z_{k})/2\pi\mathbbm{i}-\beta\cdot M_{j}+d)}.$
The result then follows immediately from (4.1.5).

### v2

@@
 Wall Crossing and the Fourier–Mukai Transform for Grassmann Flops
-Abstract.
+[…]
+[…]
+Wall Crossing and the Fourier–Mukai Transform for Grassmann Flops
+[…]
+Wall Crossing and the Fourier–Mukai Transform
+for Grassmann Flops
+[…]
+Nathan PRIDDIS a, Mark SHOEMAKER b and Yaoxiong WEN c
+[…]
+N. Priddis, M. Shoemaker and Y. Wen
+[…]
+[…]
+[…]
+b) Department of Mathematics, Colorado State University,
+[…]
+[…]
+[…]
+[…]
+Received April 26, 2024, in final form January 27, 2025; Published online February 06, 2025
+[…]
 We prove the crepant transformation conjecture for relative Grassmann flops over a smooth base $B$.
@@
 We verify that the symplectic transformation is compatible with Iritani’s integral structure, that is, that it is induced by a Fourier–Mukai transform in $K$-theory.
-1. Introduction
+[…]
+Fourier–Mukai; Grassmannian flops; wall-crossing; Gromov–Witten theory; variation of GIT
+[…]
+14N35; 14E16; 53D45
+1 Introduction
 The connection between birational geometry and quantum cohomology has been a source of great activity and interest in Gromov–Witten theory since the early development of the theory.
-As first described in Ruan’s crepant resolution conjecture in [Rua99] and later extended to more general crepant transformations in [LR01, CIT09, CR13, LLW10], it is expected that for a general crepant rational map $f:X_{-}\dashrightarrow X_{+}$, the quantum cohomology of $X_{-}$ should be equivalent to that of $X_{+}$ in a specific sense.
+As first described in Ruan’s crepant resolution conjecture in [31] and later extended to more general crepant transformations in [16, 18, 25, 27], it is expected that for a general crepant rational map $f\colon X_{-}\dashrightarrow X_{+}$, the quantum cohomology of $X_{-}$ should be equivalent to that of $X_{+}$ in a specific sense.
 More specifically, suppose we have rational maps of smooth projective varieties
 […]
-such that $f_{-}^{*}(K_{X_{-}})=f_{+}^{*}(K_{X_{+}})$.
+such that $f_{-}^{*}(K_{X_{-}})=f_{+}^{*}\bigl{(}K_{X_{+}}\bigr{)}$.
 The crepant transformation conjecture predicts that generating functions of genus zero Gromov–Witten invariants of $X_{-}$ and $X_{+}$ are equal after analytic continuation.
-Versions of the conjecture have now been proven in many specific instances, and for certain classes of crepant transformations (see e.g.
-[LR01, BG09, CIJ18, ILLW12]).
-One can further refine the crepant transformation conjecture to take into account Iritani’s integral structure (see [Iri09]) or the rational structure of Katzarkov–Kontsevich–Pantev (see [KKP08]).
+Versions of the conjecture have now been proven in many specific instances, and for certain classes of crepant transformations (see, e.g., [8, 14, 23, 27]).
+One can further refine the crepant transformation conjecture to take into account Iritani’s integral structure (see [22]) or the rational structure of Katzarkov–Kontsevich–Pantev (see [24]).
 With respect to this structure, the correspondence should be compatible with the Fourier–Mukai transform
-$\mathbb{FM}={f_{+}}_{*}f_{-}^{*}:K^{0}(X_{-})\to K^{0}(X_{+}).$
+$\mathbb{FM}={f_{+}}_{*}f_{-}^{*}\colon\ K^{0}(X_{-})\to K^{0}\bigl{(}X_{+}\bigr{)}.$
 In this paper, we prove a crepant transformation conjecture for relative Grassmann flops and show it is compatible with the Fourier–Mukai transform.
-1.1. Grassmann flops
-A rich source of examples of crepant transformations as in (1.0.1) arises from variation of GIT. Namely, suppose a reductive group $G$ acts on a variety $V$. Then we can define two GIT quotients
-$X_{\pm}=[V\sslash_{\pm}G],$
-where $+/-:G\to\mathbb{C}^{*}$ are characters lying in adjacent maximal chambers with respect to the wall and chamber structure on $\operatorname{Hom}(G,\mathbb{C}^{*})\otimes_{\mathbb{Z}}\mathbb{R}$.
+1.1 Grassmann flops
+A rich source of examples of crepant transformations as in (1.1) arises from variation of GIT.
+Namely, suppose a reductive group $G$ acts on a variety $V$.
+Then we can define two GIT quotients $X_{\pm}=[V\sslash_{\pm}G]$, where $+/-\colon G\to\mathbb{C}^{*}$ are characters lying in adjacent maximal chambers with respect to the wall and chamber structure on $\operatorname{Hom}(G,\mathbb{C}^{*})\otimes_{\mathbb{Z}}\mathbb{R}$ [^1] For those unfamiliar with the notation, by $[V\sslash_{\pm}G]$ we mean the quotient stack $[V^{\rm ss}(\pm)/G]$.
 Under certain numerical conditions, the rational map $X_{+}\dashrightarrow X_{-}$ will be crepant.
@@
 In fact, we work in the setting of relative Grassmann flops, as we describe below.
-The setup is as follows: given a smooth projective variety $B$, let $F=\oplus_{i=1}^{n}M_{i}$ and $E=\oplus_{j=1}^{n}L_{j}$ be sums of line bundles on $B$.
+The setup is as follows: given a smooth projective variety $B$, let $F=\bigoplus_{i=1}^{n}M_{i}$ and $E=\bigoplus_{j=1}^{n}L_{j}$ be sums of line bundles on $B$.
 For $r<n$, define the vector bundle $V\to B$ to be the total space of
-$\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
-There is a natural (fiberwise) action of $G=\operatorname{GL}_{r}$ on $V$.
-Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$ respectively.
+$\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
+There is a natural (fiberwise) action of $G={\rm GL}_{r}$ on $V$.
+Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$, respectively.
 Then
-$X_{+}=$ $\operatorname{tot}(S_{+}\otimes p^{*}F^{\vee})\to\operatorname{Gr}(r,E)$
-$X_{-}=$ $\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}(r,F^{\vee})$
-where $p:\operatorname{Gr}(r,E)\to B$ (resp.
-$p:\operatorname{Gr}(r,F^{\vee})\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp.
+$X_{+}=\operatorname{tot}\bigl{(}S_{+}\otimes p^{*}F^{\vee}\bigr{)}\to\operatorname{Gr}(r,E),\qquad X_{-}=\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)},$
+where $p\colon\operatorname{Gr}(r,E)\to B$ (resp.
+$p\colon\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp.
 $F^{\vee}$) and $S_{\pm}$ denotes the rank-$r$ tautological bundle on the associated Grassmann bundle.
@@
 Fix integers $0<r<k<n$ and let $B$ be the Grassmannian $\operatorname{Gr}(k,n).$ Let ${F\to B}$ be the trivial rank $k$ bundle $\mathcal{O}_{B}\otimes\mathbb{C}^{k}$ and let $E\to B$ be the rank $k$ tautological bundle on $B$.
-In this case, $X_{+}$ is
-$X_{+}=\operatorname{tot}(\hat{S}_{r}^{\oplus k})$
+In this case,
+$X_{+}=\operatorname{tot}\bigl{(}\hat{S}_{r}^{\oplus k}\bigr{)},$
 where $\hat{S}_{r}$ denotes the rank $r$ tautological bundle on $\operatorname{Fl}(r,k,n).$
 On the other hand, $X_{-}$ is the total space of a vector bundle over $\operatorname{Gr}(r,k)\times\operatorname{Gr}(k,n).$ If we let $\bar{S}_{r}$ denote the tautological rank $r$ bundle from the first factor and $S_{k}$ the tautological rank $k$ bundle from the second factor, then
-$X_{-}=\operatorname{tot}(\bar{S}_{r}^{\vee}\otimes S_{k}).$
+$X_{-}=\operatorname{tot}\bigl{(}\bar{S}_{r}^{\vee}\otimes S_{k}\bigr{)}.$
 This example can be realized as an instance of variation of GIT of quiver varieties. Consider the quiver
-[…]
+$k$$r$$k$$n$
 The associated quiver varieties (for various choices of stability conditions) are quotients of
-$\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{r})\times\operatorname{Hom}(\mathbb{C}^{r},\mathbb{C}^{k})\times\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{n})$
-by the natural right action of $\operatorname{GL}_{r}\times\operatorname{GL}_{k}$ where $\operatorname{GL}_{r}$ acts trivially on $\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{n})$ and $\operatorname{GL}_{k}$ acts trivially on $\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{r})$.
+$\operatorname{Hom}\bigl{(}\mathbb{C}^{k},\mathbb{C}^{r}\bigr{)}\times\operatorname{Hom}\bigl{(}\mathbb{C}^{r},\mathbb{C}^{k}\bigr{)}\times\operatorname{Hom}\bigl{(}\mathbb{C}^{k},\mathbb{C}^{n}\bigr{)}$
+by the natural right action of ${\rm GL}_{r}\times{\rm GL}_{k}$ where ${\rm GL}_{r}$ acts trivially on $\operatorname{Hom}\bigl{(}\mathbb{C}^{k},\mathbb{C}^{n}\bigr{)}$ and ${\rm GL}_{k}$ acts trivially on $\operatorname{Hom}\bigl{(}\mathbb{C}^{k},\mathbb{C}^{r}\bigr{)}$.
 Then $X_{+}$ and $X_{-}$ from above are GIT quotients of this representation associated to the characters
-$+:(g_{1},g_{2})\mapsto\det(g_{1})\cdot\det(g_{2})$
-and
-$-:(g_{1},g_{2})\mapsto\det(g_{1})^{-1}\cdot\det(g_{2})$
+$+\colon\ (g_{1},g_{2})\mapsto\det(g_{1})\cdot\det(g_{2})\qquad\text{and}\qquad-\colon\ (g_{1},g_{2})\mapsto\det(g_{1})^{-1}\cdot\det(g_{2}),$
 respectively.
-1.2. Statement of theorem
-With $X_{\pm}$ as above, consider the infinite dimensional (symplectic) vector space
-$\mathcal{H}_{X_{\pm}}=H^{*}_{\mathbb{T}}(X_{\pm})(\!(z^{-1})\!)[\![Q,q_{\pm}]\!].$
+1.2 Statement of theorem
+With $X_{\pm}$ as above, consider the infinite-dimensional (symplectic) vector space
+$\mathcal{H}_{X_{\pm}}=H^{*}_{\mathbb{T}}(X_{\pm})\bigl{(}\!\bigl{(}z^{-1}\bigr{)}\!\bigr{)}[\![Q,q_{\pm}]\!].$
 Here $Q$ and $q_{\pm}$ are Novikov parameters with respect to the base $B$ and fiber directions of $X_{\pm}\to B$, respectively.
-Using the relative quasimap theory of [Oh21], one can define $I$-functions
-$I_{X_{\pm}}(Q,q_{\pm},z)$
-lying in $\mathcal{H}_{X_{\pm}}$.
+Using the relative quasimap theory of [29], one can define $I$-functions $I_{X_{\pm}}(Q,q_{\pm},z)$ lying in $\mathcal{H}_{X_{\pm}}$.
 These functions are shown to lie on Givental’s Lagrangian cone $\mathcal{L}_{X_{\pm}}\subset\mathcal{H}_{X_{\pm}}$ for the respective spaces and may therefore be expressed as generating functions of genus-zero Gromov–Witten invariants of $X_{\pm}$.
 See Section 3 for details.
-One can then extend $\mathcal{H}_{X_{\pm}}$ by ~$\widetilde{\mathcal{H}}_{X_{\pm}}=\mathcal{H}_{X_{\pm}}[\log z][z^{-1/2}]$, and following Iritani in [Iri09], define the map
-$\Psi:K^{0}_{\mathbb{T}}(X)$$\to\widetilde{\mathcal{H}}_{X}$
-$A$ $\mapsto z^{-\mu}z^{\rho}\left(\hat{\Gamma}_{X}\cup(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\right)$
-where $\mathbbm{i}=\sqrt{-1}$, $\mu$ is the grading operator, $\rho=c_{1}^{\mathbb{T}}(X_{\pm})$, and $\hat{\Gamma}_{X}$ is the Gamma class—a characteristic class associated with the Gamma function $\Gamma(1+z)$.
-We prove the following:
+One can then extend $\mathcal{H}_{X_{\pm}}$ by ~$\widetilde{\mathcal{H}}_{X_{\pm}}=\mathcal{H}_{X_{\pm}}[\log z]\bigl{[}z^{-1/2}\bigr{]}$, and following Iritani in [22], define the map
+$\Psi\colon\ K^{0}_{\mathbb{T}}(X)\to\widetilde{\mathcal{H}}_{X},\qquad A\mapsto z^{-\mu}z^{\rho}\bigl{(}\hat{\Gamma}_{X}\cup(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\bigr{)},$
+where ${\rm i}=\sqrt{-1}$, $\mu$ is the grading operator, $\rho=c_{1}^{\mathbb{T}}(X_{\pm})$, and $\hat{\Gamma}_{X}$ is the Gamma class – a characteristic class associated with the Gamma function $\Gamma(1+z)$.
+We prove the following.
 Theorem 1.2 (Theorem 5.4).
-There exists a linear symplectic isomorphism
-$\mathbb{U}:\widetilde{\mathcal{H}}_{X_{-}}\to\widetilde{\mathcal{H}}_{X_{+}}$
-such that the following diagram commutes:
-(1.2.1) ${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\mathbb{U}}$
+There exists a linear symplectic isomorphism $\mathbb{U}\colon\widetilde{\mathcal{H}}_{X_{-}}\to\widetilde{\mathcal{H}}_{X_{+}}$ such that the following diagram commutes:
+${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\mathbb{U}}$ (1.2)
 There is a path $\hat{\gamma}$ from a neighborhood of $q_{+}=0$ to a neighborhood of $q_{-}=q_{+}^{-1}=0$ such that
-(1.2.2) $\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}},$
+$\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}},$ (1.3)
 where $\widetilde{I_{X_{+}}}$ denotes the analytic continuation of $I_{X_{+}}$ along $\hat{\gamma}.$
-Theorem 1.2 generalizes previous known cases of the crepant transformation conjecture in two directions.
-First, it is to our knowledge one of the first examples of a wall crossing result for a variation of GIT quotients by a non-abelian group—in this case $\operatorname{GL}_{r}$.
+Theorem 1.2 generalizes previous known cases of the crepant transformation conjecture in two directions (see Section 1.4 for more detailed comparisons with previous work).
+First, to our knowledge it is one of the first examples of a wall crossing result based on analytic continuation of generating functions for a variation of GIT quotients by a non-abelian group – in this case ${\rm GL}_{r}$.
 Second, it provides an example of the conjecture for fiber bundles over an arbitrary smooth base $B$.
-Thus our result is new even in the special case of the quotient by $\operatorname{GL}_{1}=\mathbb{C}^{*}$, where the two GIT quotients are toric bundles over $B$.
-1.3. Strategy of proof
+Thus, in the special case of the quotient by ${\rm GL}_{1}=\mathbb{C}^{*}$, where the two GIT quotients are toric bundles over B, our result generalizes a particular case of the main result in [14].
+We expect these techniques to be applicable beyond Grassmann flops, to more general variation of GIT quotients with respect to actions of ${\rm GL}_{n}$.
+In this way, the current paper may be viewed as a proof of concept.
+It would also be very interesting to extend these methods to quotients by other non-abelian groups.
+We hope to revisit this question in future work.
+1.3 Strategy of proof
 The proof relies on the abelian/non-abelian correspondence in Gromov–Witten theory, which allows us to construct the $I$-functions $I_{X_{\pm}}$ for the GIT quotients $X_{\pm}$ in terms of $I$-functions of the associated abelian quotients
@@
 where $T$ is the maximal torus in $G$ consisting of diagonal matrices. More precisely, the function $I_{X_{\pm}}(Q,q_{\pm},z)$ is obtained as the specialization of
-$I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,\pm},\ldots,q_{r,\pm},z)$
-at $q_{1,\pm}=\cdots=q_{r,\pm}=q_{\pm}.$ Here $I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ is a twist of the $I$-function for the associated abelian quotient by the root bundle for $G$, which we denote by $\mathfrak{g}/\mathfrak{t}$.
+$I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,\pm},\dots,q_{r,\pm},z)\qquad\text{at $q_{1,\pm}=\cdots=q_{r,\pm}=q_{\pm}$},$
+where $I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ is a twist of the $I$-function for the associated abelian quotient by the root bundle for $G$, which we denote by $\mathfrak{g}/\mathfrak{t}$.
 See Section 3 for details.
@@
 It is not clear how to compute $\widetilde{I_{X_{+}}}$ directly.
-Instead we compute the analytic continuation of $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\ldots,q_{r,+},z)$ along each of the variables $q_{k,+}$ independently.
-That is to say, we analytically continue $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\ldots,q_{r,+},z)$ along a concatenated path $\gamma_{1}\star\cdots\star\gamma_{r}$, where each $\gamma_{k}$ leaves the variables $q_{i,+}$ constant for $i\neq k$, and goes from 00 to $\infty$ in the $q_{k,+}$ direction.
+Instead, we compute the analytic continuation of $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\dots,q_{r,+},z)$ along each of the variables $q_{k,+}$ independently.
+That is to say, we analytically continue $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\dots,q_{r,+},z)$ along a concatenated path $\gamma_{1}\star\cdots\star\gamma_{r}$, where each $\gamma_{k}$ leaves the variables $q_{i,+}$ constant for $i\neq k$, and goes from 00 to $\infty$ in the $q_{k,+}$ direction.
 We then homotope the path $\gamma_{1}\star\cdots\star\gamma_{r}$ to a path $\hat{\gamma}$ which lies entirely in the locus $q_{1,+}=\cdots=q_{r,+},$ thereby obtaining $\widetilde{I_{X_{+}}}$.
@@
 This involves an explicit computation of $\mathbb{FM}$ on generators of $K^{0}_{\mathbb{T}}(X_{-})$.
-The proof that the map $\mathbb{U}$ in (1.2.1) satisfies (1.2.2) then involves a nontrivial identity of anti-symmetric functions (Lemma 5.2).
-1.4. Relation to other work
-The structure of the proof follows that given in [CIT09, CR10, CIR14, CIJ18].
-In particular the formulation of the correspondence and the proof via localization is closely related to [CIJ18].
+The proof that the map $\mathbb{U}$ in (1.2) satisfies (1.3) then involves a nontrivial identity of anti-symmetric functions (Lemma 5.2).
+1.4 Relation to other work
+The structure of the proof follows that given in [10, 11, 14, 16].
+In particular, the formulation of the correspondence and the proof via localization is closely related to [14].
 The new ingredients in the current article are the homotopy of the path of analytic continuation, the generalization to the relative setting, and the combinatorial identity verifying agreement with $\mathbb{FM}$, which does not appear in the case of abelian variation of GIT.
-In the final stages of writing this paper, we learned of a related result by Lutz–Shafi–Webb (see [LSW]).
+In the special case of the quotient by ${\rm GL}_{1}=\mathbb{C}^{*}$, where the two GIT quotients are toric bundles over $B$, (a compactification of) this example was studied as the local model in [26] and a similar result was obtained at the level of $D$-modules and quantum cohomology using different methods.
+It is possible that the comparison of $I$-functions in this paper may be obtained from their results, although it is not immediately apparent how to pursue such a strategy.
+On the other hand, our comparison with the $K$-theoretic Fourier–Mukai transform appears to be new even in this case.
+The Gromov–Witten theory of spaces related by variation of GIT quotients by nonabelian groups has also been studied in [21] in the very general setting of GIT quotients of a projective variety by a complex reductive group.
+The setting of the current paper is different, as our GIT quotients are of quasi-projective varieties, and the form of the comparison is also different.
+The generating functions being compared in loc.
+cit.
+are different than in this paper, and the difference between the two generating functions is expressed as an explicit sum of residues rather than via analytic continuation.
+In the final stages of writing this paper, we learned of a related result by Lutz–Shafi–Webb (see [28]).
 While both papers use the idea of homotoping the path of analytic continuation, there are significant differences in the two results and their proofs.
@@
 In this paper, we compare the map $\mathbb{U}$ to the Fourier–Mukai transform $\mathbb{FM}$.
-In [LSW], they instead compare $\mathbb{U}$ with the symplectic transformation of the associated abelian wall crossing.
+In [28], they instead compare $\mathbb{U}$ with the symplectic transformation of the associated abelian wall crossing.
 This leads to two different proofs that the map $\mathbb{U}$ is a symplectic isomorphism and has a well-defined non-equivariant limit.
@@
 We hope to revisit this issue in the future to obtain a more direct and conceptual understanding of this compatibility.
-1.5. Acknowledgements
-We would like to thank Tom Coates, Wendelin Lutz, Jeongseok Oh, Yongbin Ruan, Ed Segal, Qaasim Shafi, and Rachel Webb for the useful discussions.
-N.
-Priddis is supported by the Simons Foundation Travel Grant 586691.
-M.
-Shoemaker is supported by the Simons Foundation Travel Grant 958189.
-Y.
-Wen is supported by a KIAS Individual Grant (MG083902) at the Korea Institute for Advanced Study.
-2. Geometric setup
-Let $E=\oplus_{j=1}^{n}L_{j}$ and $F=\oplus_{i=1}^{n}M_{i}$ be sums of line bundles over a smooth projective variety $B$. Choose $r<n$, and define the vector bundle $V\to B$ to be the total space of
-$\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
-This vector bundle comes equipped with a (right) action of $G=\operatorname{Aut}(\mathbb{C}^{r})=\operatorname{GL}_{r}$ given fiberwise by
-$(X,Y)\cdot g=(g^{-1}X,Yg).$
-Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$ respectively. Then
-$X_{+}=$ $\operatorname{tot}(S_{+}\otimes p^{*}F^{\vee})\to\operatorname{Gr}(r,E)$
-$X_{-}=$ $\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}(r,F^{\vee})$
-where $p:\operatorname{Gr}(r,E)\to B$ (resp.
-$p:\operatorname{Gr}(r,F^{\vee})\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp.
+2 Geometric setup
+Let $E=\bigoplus_{j=1}^{n}L_{j}$ and $F=\bigoplus_{i=1}^{n}M_{i}$ be sums of line bundles over a smooth projective variety $B$.
+Choose $r<n$, and define the vector bundle $V\to B$ to be the total space of
+$\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
+This vector bundle comes equipped with a (right) action of $G=\operatorname{Aut}(\mathbb{C}^{r})={\rm GL}_{r}$ given fiberwise by
+$(X,Y)\cdot g=\bigl{(}g^{-1}X,Yg\bigr{)}.$
+Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$, respectively. Then
+$X_{+}=\operatorname{tot}\bigl{(}S_{+}\otimes p^{*}F^{\vee}\bigr{)}\to\operatorname{Gr}(r,E),\qquad X_{-}=\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)},$
+where $p\colon\operatorname{Gr}(r,E)\to B$ (resp.
+$p\colon\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp.
 $F^{\vee}$) and $S_{\pm}$ denotes the rank-$r$ tautological bundle on the associated Grassmann bundle.
@@
 The birational map
-$f:X_{-}\dashrightarrow X_{+}$
-is a relative version of what is known as a Grassmann flop. We construct a resolution of this map in Section 2.3 following [BLVdB16], and compute the associated Fourier–Mukai transform in $K$-theory.
-2.1. Cohomology
+$f\colon\ X_{-}\dashrightarrow X_{+}$
+is a relative version of what is known as a Grassmann flop. We construct a resolution of this map in Section 2.3 following [9], and compute the associated Fourier–Mukai transform in $K$-theory.
+2.1 Cohomology
 Let $(\mathbb{C}^{*})^{n}$ act on $F$ by scaling, such that the $i$th factor scales $M_{i}$.
@@
 Unless otherwise specified, all cohomology groups are $\mathbb{T}$-equivariant.
-We define $z_{i}=c_{1}^{\mathbb{T}}(M_{i}^{\vee})$ and $x_{i}=c_{1}^{\mathbb{T}}(L_{i}^{\vee})$.
+We define $z_{i}=c_{1}^{\mathbb{T}}\bigl{(}M_{i}^{\vee}\bigr{)}$ and $x_{i}=c_{1}^{\mathbb{T}}\bigl{(}L_{i}^{\vee}\bigr{)}$.
 Let $R=\mathbb{C}^{r}$ denote the right representation of $G$ (viewing $\mathbb{C}^{r}$ as row vectors), given by
-(2.1.1) $v\cdot g=vg.$
+$v\cdot g=vg.$ (2.1)
 The $G$-equivariant vector bundle $R\times V\to V$ induces a vector bundle on both $X_{+}$ and $X_{-}$ which, by abuse of notation, we also denote by $R$.
-On $X_{-}$ it is (the pullback of) the tautological bundle $S_{-}\to\operatorname{Gr}(r,F^{\vee})$ and on $X_{+}$ it is (the pullback of) the dual of the tautological bundle $S_{+}\to\operatorname{Gr}(r,E)$.
-Denote the equivariant Chern roots of the dual $R^{\vee}$ by $y_{1},\ldots,y_{r}$.
-Let
-$\pi_{\pm}:X_{\pm}\to B$
-denote the projection.
-To simplify notation we will denote $\pi_{\pm}^{*}(x_{i})$ (resp.
+On $X_{-}$ it is (the pullback of) the tautological bundle $S_{-}\to\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}$ and on $X_{+}$ it is (the pullback of) the dual of the tautological bundle $S_{+}\to\operatorname{Gr}(r,E)$.
+Denote the equivariant Chern roots of the dual $R^{\vee}$ by $y_{1},\dots,y_{r}$.
+Let $\pi_{\pm}\colon X_{\pm}\to B$ denote the projection.
+To simplify notation, we will denote $\pi_{\pm}^{*}(x_{i})$ (resp.
 $\pi_{\pm}^{*}(z_{j})$) simply by $x_{i}$ (resp.
@@
 The equivariant cohomology of $X_{+}$ (resp.
-$X_{-}$), denoted $H^{*}_{\mathbb{T}}(X_{\pm})$, is generated as an algebra by elements of
-$H^{*}(B)[z_{i},x_{i}]_{i=1}^{n}$
-and $\sigma_{1},\ldots,\sigma_{r}$, the elementary symmetric polynomials in $y_{1},\ldots,y_{r}$, subject to the relations:
-(2.1.2) $I_{+}=\left\{\left[\frac{\prod_{i=1}^{n}(1-x_{i})}{\prod_{j=1}^{r}(1+y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}$
-in $H^{*}_{\mathbb{T}}(X_{+})$, and
-(2.1.3) $I_{-}=\left\{\left[\frac{\prod_{i=1}^{n}(1+z_{i})}{\prod_{j=1}^{r}(1-y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}.$
+$X_{-}$), denoted $H^{*}_{\mathbb{T}}(X_{\pm})$, is generated as an algebra by elements of $H^{*}(B)[z_{i},x_{i}]_{i=1}^{n}$ and $\sigma_{1},\dots,\sigma_{r}$, the elementary symmetric polynomials in $y_{1},\dots,y_{r}$, subject to the relations:
+$I_{+}=\left\{\left[\frac{\prod_{i=1}^{n}(1-x_{i})}{\prod_{j=1}^{r}(1+y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}$
+in $H^{*}_{\mathbb{T}}\bigl{(}X_{+}\bigr{)}$, and
+$I_{-}=\left\{\left[\frac{\prod_{i=1}^{n}(1+z_{i})}{\prod_{j=1}^{r}(1-y_{j})}\right]_{l}=0\ \middle|\ l>n-r\right\}$
 in $H^{*}_{\mathbb{T}}(X_{-})$.
-In the above formulas, $\left[-\right]_{l}$ denotes the degree $l$ part.
+(See [20, Example 14.6.6] for an equivalent presentation in the context of Chow rings, the above presentation follows by the same argument.) In the above formulas, $\left[-\right]_{l}$ denotes the degree $l$ part.
 We will denote by $R_{\mathbb{T}}$ the ring $H^{*}_{\mathbb{T}}(pt)$, and by $S_{\mathbb{T}}$ its localization with respect to non-zero homogeneous elements.
 We will also make use of the completion $\hat{S}_{\mathbb{T}}$, defined by
-$\hat{S}_{\mathbb{T}}=\{\sum_{d\in\mathbb{Z}}a_{d}|\text{there exists $d_{0}\in\mathbb{Z}$ such that $a_{d}=0$ for $d<d_{0}$}\},$
+$\hat{S}_{\mathbb{T}}=\biggl{\{}\sum_{d\in\mathbb{Z}}a_{d}\,\bigg{|}\,\text{there exists $d_{0}\in\mathbb{Z}$ such that $a_{d}=0$ for $d<d_{0}$}\biggr{\}},$
 where $a_{d}$ lies in the degree $d$ homogeneous part of $S_{\mathbb{T}}.$ We will often work with the localized equivariant cohomology ring $H^{*}_{\mathbb{T}}(X_{\pm})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}},$ or its completion $H^{*}_{\mathbb{T}}(X_{\pm})\otimes_{R_{\mathbb{T}}}\hat{S}_{\mathbb{T}},$
 The connected components of the $\mathbb{T}$-fixed locus of $X_{+}$ (resp.
-$X_{-}$) are indexed by subsets of $\{1,\ldots,n\}$ of size $r$, corresponding to choosing an $r\times r$ minor of $\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ (resp.
-$\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$).
+$X_{-}$) are indexed by subsets of $\{1,\dots,n\}$ of size $r$, corresponding to choosing an $r\times r$ minor of $\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ (resp.
+$\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$).
 Thus each connected component is isomorphic to $B$.
@@
 $D_{-}$) denote the set indexing these fixed loci.
-We may represent an element of $D_{\pm}$ by a function $\delta:\{1,\ldots,r\}\to\{1,\ldots,n\}$ such that $\delta(i)<\delta(j)$ for all $i<j$.
+We may represent an element of $D_{\pm}$ by a function $\delta\colon\{1,\dots,r\}\to\{1,\dots,n\}$ such that $\delta(i)<\delta(j)$ for all $i<j$.
 We will further denote $\delta(i)$ by $\delta_{i}$.
 Lemma 2.1.
-  1. (1)
-    For $\delta^{-}\in D_{-}$, let $B_{\delta^{-}}$ denote the associated fixed locus, isomorphic to $B$. Then $R|_{B_{\delta^{-}}}\cong\oplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}$.
-  2. (2)
-    For $\delta^{+}\in D_{+}$, let $B_{\delta^{+}}$ denote the associated fixed locus, isomorphic to $B$. Then $R|_{B_{\delta^{+}}}\cong\oplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}$.
+  1. $1$
+    For $\delta^{-}\in D_{-}$, let $B_{\delta^{-}}$ denote the associated fixed locus, isomorphic to $B$. Then
+    $R|_{B_{\delta^{-}}}\cong\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}.$
+  2. $2$
+    For $\delta^{+}\in D_{+}$, let $B_{\delta^{+}}$ denote the associated fixed locus, isomorphic to $B$. Then
+    $R|_{B_{\delta^{+}}}\cong\bigoplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}.$
 Proof.
 We will prove the first statement. The fixed locus may be expressed as the GIT quotient
-$\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})\sslash_{\theta_{-}}G.$
-There is a tautological map of vector bundles $q^{*}(\oplus_{i=1}^{r}M_{\delta^{-}_{i}})\to R^{\vee}$ over the stack
-$[\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})/G,]$
-where $q:[\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})/G,]\to B$ is the projection.
-The $\theta_{-}$-stable locus of $\mathcal{H}om(\oplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})$ is exactly the locus where this map is an isomorphism.
+$\mathcal{H}{\rm om}\biggl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r}\biggr{)}\sslash_{\theta_{-}}G.$
+There is a tautological map of vector bundles $q^{*}\bigl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}}\bigr{)}\to R^{\vee}$ over the stack
+$\biggl{[}\mathcal{H}{\rm om}\biggl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r}\biggr{)}/G\biggr{]},$
+where $q\colon\bigl{[}\mathcal{H}{\rm om}\bigl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r}\bigr{)}/G\bigr{]}\to B$ is the projection.
+The locus where this map is an isomorphism is exactly the $\theta_{-}$-stable locus of $\mathcal{H}{\rm om}\bigl{(}\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}},\mathcal{O}_{B}\otimes\mathbb{C}^{r}\bigr{)}$.
 ∎
 Corollary 2.2.
-  1. (1)
+  1. $1$
     In $H^{*}_{\mathbb{T}}(X_{-})$, up to a choice of ordering, the restriction of $y_{i}$ to $B_{\delta^{-}}$ is $-z_{\delta^{-}_{i}}$.
-  2. (2)
-    In $H^{*}_{\mathbb{T}}(X_{+})$, up to a choice of ordering, the restriction of $y_{i}$ to $B_{\delta^{+}}$ is $-x_{\delta^{+}_{i}}$.
-2.2. The associated abelian quotients
+  2. $2$
+    In $H^{*}_{\mathbb{T}}\bigl{(}X_{+}\bigr{)}$, up to a choice of ordering, the restriction of $y_{i}$ to $B_{\delta^{+}}$ is $-x_{\delta^{+}_{i}}$.
+2.2 The associated abelian quotients
 Let $T\subset G$ denote the maximal torus in $G$ consisting of diagonal matrices.
@@
 We define the abelian quotients $X_{T,\pm}$ associated to $X_{\pm}$ to be the GIT quotients $V\sslash_{\pm}T$.
-They are fiber products over $B$ of $r$ copies of $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)$ and $r$ copies of $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)$, resp.:
-$X_{T,{-}}$ $=\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)\times_{B}\cdots\times_{B}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)$
-$X_{T,{+}}$ $=\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)\times_{B}\cdots\times_{B}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right).$
+They are fiber products over $B$ of $r$ copies of $\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)}$ and $r$ copies of $\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}$, respectively:
+$X_{T,{-}}=\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)}\times_{B}\cdots\times_{B}\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)},$
+$X_{T,{+}}=\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}\times_{B}\cdots\times_{B}\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}.$
 There is a series of intermediate GIT quotients defined as follows. For $0\leq j\leq r$, let $\theta_{j}\in\operatorname{Hom}(T,\mathbb{C}^{*})$ denote the character
-$\theta_{j}((t_{1},\ldots,t_{r}))=\prod_{i=1}^{j}t_{i}\cdot\prod_{i=j+1}^{r}t_{i}^{-1},$
+$\theta_{j}((t_{1},\dots,t_{r}))=\prod_{i=1}^{j}t_{i}\cdot\prod_{i=j+1}^{r}t_{i}^{-1},$
 and let $X_{T,j}$ denote the GIT quotient $V\sslash_{\theta_{j}}T$:
-$X_{T,j}=\prod_{i=1}^{j}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)\times_{B}\prod_{i=j+1}^{r}\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right)$
+$X_{T,j}=\prod_{i=1}^{j}\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}\times_{B}\prod_{i=j+1}^{r}\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)},$
 where the products above are fiber products over $B$. Then $X_{T,{+}}=X_{T,r}$ and $X_{T,{-}}=X_{T,0}$.
 The equivariant cohomology ring $H^{*}_{\mathbb{T}}(X_{T,j})$ is given by the quotient of
-$H^{*}_{\mathbb{T}}(B)[z_{i},x_{i}]_{i=1}^{n}[y_{1},\ldots,y_{r}]$
+$H^{*}_{\mathbb{T}}(B)[z_{i},x_{i}]_{i=1}^{n}[y_{1},\dots,y_{r}]$
 by the relations
 […]
-The $\mathbb{T}$-fixed loci of $X_{T,\pm}$ are indexed by (not necessarily injective) functions $f:\{1,\ldots,r\}\to\{1,\ldots,n\}$.
-For such an $f$, the associated fixed locus $B_{f}$ is the subspace of $X_{T,+}$ such that, on the $i$th factor of $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)$ in the fiber product, the homogeneous coordinates are all zero except for in the $f(i)$th factor of $\mathbb{P}(E)$.
+The $\mathbb{T}$-fixed loci of $X_{T,\pm}$ are indexed by (not necessarily injective) functions $f\colon\{1,\dots,r\}\to\{1,\dots,n\}$.
+For such an $f$, the associated fixed locus $B_{f}$ is the subspace of $X_{T,+}$ such that, on the $i$th factor of $\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)}$ in the fiber product, the homogeneous coordinates are all zero except for in the $f(i)$th factor of $\mathbb{P}(E)$.
 In other words,
-$B\cong B_{f}=\mathbb{P}(L_{f(1)})\times_{B}\cdots\times_{B}\mathbb{P}(L_{f(r)})\subset\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E)\subset X_{T,+}.$
+$B\cong B_{f}=\mathbb{P}\bigl{(}L_{f(1)}\bigr{)}\times_{B}\cdots\times_{B}\mathbb{P}\bigl{(}L_{f(r)}\bigr{)}\subset\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E)\subset X_{T,+}.$
 An analogous description holds for $X_{T,-}$ and indeed all $X_{T,j}$.
-We will make use of the following weak form of the abelian/non-abelian correspondence in cohomology:
+We will make use of the following weak form of the abelian/non-abelian correspondence in cohomology.
 Proposition 2.3.
 The pullback
-$H^{*}_{\mathbb{T}}(X_{\pm})\to(H^{*}_{\mathbb{T}}(V^{s,\pm}(G)/T))^{W}$
-is an isomorphism, where $V^{s,\pm}(G)$ denotes the stable locus of $V$ with respect to $\det^{\pm}:G\to\mathbb{C}^{*}$ and $W$ denotes the Weyl group. Furthermore, the pullback map
-$H^{*}_{\mathbb{T}}(X_{T,\pm})\to H^{*}_{\mathbb{T}}(V^{s,\pm}(G)/T)$
+$H^{*}_{\mathbb{T}}(X_{\pm})\to\bigl{(}H^{*}_{\mathbb{T}}\bigl{(}V^{s,\pm}(G)/T\bigr{)}\bigr{)}^{W}$
+is an isomorphism, where $V^{s,\pm}(G)$ denotes the stable locus of $V$ with respect to $\det^{\pm}\colon G\to\mathbb{C}^{*}$ and $W$ denotes the Weyl group. Furthermore, the pullback map
+$H^{*}_{\mathbb{T}}(X_{T,\pm})\to H^{*}_{\mathbb{T}}\bigl{(}V^{s,\pm}(G)/T\bigr{)}$
 is surjective.
-We note that although not originally stated in this context, the result holds equivariantly, as one can check from the explicit presentation of $H^{*}_{\mathbb{T}}(X_{\pm})$ given in Section 2.1.
-The proposition allows us to lift cohomology classes in $X_{\pm}$ to $W$-invariant classes in $H^{*}_{\mathbb{T}}(X_{T,\pm})$.
-For future use we will use $F_{\pm}$ to denote the set of functions $f:\{1,\ldots,r\}\to\{1,\ldots,n\}$, i.e.
-$F_{\pm}$ indexes the set of fixed loci of $X_{T,\pm}$.
-We will use $In_{\pm}$ to denote the set of injective functions $\{1,\ldots,r\}\to\{1,\ldots,n\}$, noting that $In_{\pm}$ indexes the fixed loci of $V^{s,\pm}(G)/T.$
-2.3. Fourier–Mukai transform
-We can define a resolution of the birational map
-$f:X_{-}\dashrightarrow X_{+}$
-as follows. Let
-$\tilde{V}=\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$
-and let $\tilde{G}=GL_{r}\times GL_{r}$. Consider the right action of $\tilde{G}$ on $\tilde{V}$ given by
-$(X,U,Y)\cdot(g_{1},g_{2})=(g_{1}^{-1}X,g_{2}^{-1}Ug_{1},Yg_{2}).$
-Let $\tilde{\theta}\in\hom(\tilde{G},\mathbb{C}^{*})$ denote the character
+See [6, Proposition 1] for details. The proposition allows us to lift cohomology classes in $X_{\pm}$ to $W$-invariant classes in $H^{*}_{\mathbb{T}}(X_{T,\pm})$.
+For future use we will use $F_{\pm}$ to denote the set of functions $f\colon\{1,\dots,r\}\to\{1,\dots,n\}$, i.e., $F_{\pm}$ indexes the set of fixed loci of $X_{T,\pm}$.
+We will use ${\rm In}_{\pm}$ to denote the set of injective functions $\{1,\dots,r\}\to\{1,\dots,n\}$, noting that ${\rm In}_{\pm}$ indexes the fixed loci of $V^{s,\pm}(G)/T.$
+2.3 Fourier–Mukai transform
+We can define a resolution of the birational map $f\colon X_{-}\dashrightarrow X_{+}$ as follows. Let
+$\tilde{V}=\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$
+and let $\tilde{G}={\rm GL}_{r}\times{\rm GL}_{r}$. Consider the right action of $\tilde{G}$ on $\tilde{V}$ given by
+$\bigl{(}\tilde{M},\tilde{D},\tilde{N}\bigr{)}\cdot(g_{1},g_{2})=\bigl{(}g_{1}^{-1}\tilde{M},g_{2}^{-1}\tilde{D}g_{1},\tilde{N}g_{2}\bigr{)}.$
+Let $\tilde{\theta}\in\operatorname{Hom}\bigl{(}\tilde{G},\mathbb{C}^{*}\bigr{)}$ denote the character
 $\tilde{\theta}(g_{1},g_{2})=\det(g_{1})^{-1}\det(g_{2}).$
-Define $\tilde{X}$ to be the GIT quotient
-$\tilde{X}=\tilde{V}\sslash_{\tilde{\theta}}\tilde{G}.$
-The morphism $f_{-}:\tilde{X}\to X_{-}$ is induced by the morphism of stacks
-$[\tilde{V}/\tilde{G}]\to[V/G]$
-defined by the map
-$\tilde{V}$ $\to V$
-$(X,U,Y)$ $\mapsto(X,YU)$
+Define $\tilde{X}$ to be the GIT quotient, $\tilde{X}=\tilde{V}\sslash_{\tilde{\theta}}\tilde{G}$.
+The morphism $f_{-}\colon\tilde{X}\to X_{-}$ is induced by the morphism of stacks $\bigl{[}\tilde{V}/\tilde{G}\bigr{]}\to[V/G]$ defined by the map
+$\tilde{V}\to V,\qquad\bigl{(}\tilde{M},\tilde{D},\tilde{N}\bigr{)}\mapsto\bigl{(}\tilde{M},\tilde{N}\tilde{D}\bigr{)}$
 and the homomorphism
-$\tilde{G}$ $\to G$
-$(g_{1},g_{2})$ $\mapsto g_{1}.$
-There is a similar description of the map $f_{+}:\tilde{X}\to X_{+}.$ The $\mathbb{T}$-action on $E$ and $F$ induces an action on $\tilde{X}$ for which the maps $f_{\pm}:\tilde{X}\to X_{\pm}$ are equivariant.
+$\tilde{G}\to G,\qquad(g_{1},g_{2})\mapsto g_{1}.$
+There is a similar description of the map $f_{+}\colon\tilde{X}\to X_{+}.$ The $\mathbb{T}$-action on $E$ and $F$ induces an action on $\tilde{X}$ for which the maps $f_{\pm}\colon\tilde{X}\to X_{\pm}$ are equivariant.
 We define
-(2.3.1) $\mathbb{FM}:K^{0}_{\mathbb{T}}(X_{-})\to K^{0}_{\mathbb{T}}(X_{+})$
+$\mathbb{FM}\colon\ K^{0}_{\mathbb{T}}(X_{-})\to K^{0}_{\mathbb{T}}(X_{+})$ (2.3)
 to be $\mathbb{FM}=(f_{+})_{*}\circ f_{-}^{*}.$
-In [BLVdB16], it is proven that the associated functor $\mathbb{FM}:D^{b}(X_{-})\to D^{b}(X_{+})$, between derived categories is an equivalence in the case that $B$ is a point.
-In [BCF+21] it is shown that this equivalence may also be described using windows.
+Lemma 2.4.
+Let $H=\operatorname{tot}\bigl{(}F^{\vee}\otimes E\bigr{)}=\operatorname{tot}(\mathcal{H}{\rm om}(F,E))$. The space $\tilde{X}$ is equal to the fiber product $X_{+}\times_{H}X_{-}.$
+Proof.
+It suffices to show that $\tilde{X}$ satisfies the universal property of fiber products.
+Suppose that $q\colon S\to B$ is a $B$-scheme.
+A morphism $S\to X_{+}$ over $B$ corresponds to a rank $r$ vector bundle $U_{+}\to S$ together with linear maps $M_{+}\colon F\to U_{+}$ and $N_{+}\colon U_{+}\to E$ such that $N_{+}$ is full rank.
+A morphism $S\to X_{-}$ over $B$ corresponds to a rank $r$ vector bundle $U_{-}\to S$ together with linear maps $M_{-}\colon F\to U_{-}$, $N_{-}\colon U_{-}\to E$ such that $M_{-}$ is full rank.
+On the other hand, a morphism $S\to\tilde{X}$ corresponds to a pair of rank $r$ vector bundles $U_{+/-}\to S$ (obtained via the induced map $S\to B{\rm GL}_{r}\times B{\rm GL}_{r}$) together with linear maps $\tilde{M}\colon F\to U_{-}$, $\tilde{D}\colon U_{-}\to U_{+}$, and $\tilde{N}\colon U_{+}\to E,$ such that the first and last are both full rank.
+Suppose we have morphisms $S\to X_{+/-}$ which agree when composed with the natural maps $X_{+/-}\to H$.
+This yields two triples $(U_{+},M_{+},N_{+})$ and $(U_{-},M_{-},N_{-})$ such that $N_{+}\circ M_{+}=N_{-}\circ M_{-}$.
+We define a morphism $S\to\tilde{X}$ as follows.
+Note first that since $N_{+}$ is full rank, the condition $N_{+}\circ M_{+}=N_{-}\circ M_{-}$ implies that $N_{-}$ factors through a map $\tilde{D}\colon U_{-}\to U_{+}$.
+Then define $\tilde{M}=M_{-}$ and $\tilde{N}=N_{+}.$ The triple $\tilde{M}$, $\tilde{D}$, $\tilde{N}$ defines a map $S\to\tilde{X}$ as in the previous paragraph.
+It is easy to check that the compositions $S\to\tilde{X}\to X_{+/-}$ agree with the original morphisms $S\to X_{+/-}$ from the previous paragraph.
+Thus $\tilde{X}$ satisfies the universal property as desired.
+∎
+In [9, Theorem D], it is proven that the associated functor $\mathbb{FM}\colon D^{b}(X_{-})\to D^{b}\bigl{(}X_{+}\bigr{)}$, between derived categories is an equivalence in the case that $B$ is a point.
+[^2] In [9], the Fourier–Mukai kernel used is $\mathcal{O}_{X_{+}\times_{H}X_{-}}$.
+This agrees with our kernel $\mathcal{O}_{\tilde{X}}$ by Lemma 2.4.
+In [3], it is shown that this equivalence may also be described using windows.
 The arguments naturally extend to our setting.
-We will use this fact only to conclude that the map (2.3.1) is an isomorphism and preserves the pairing.
-We compute $\mathbb{FM}$ using localization as in [Tho92].
-The fixed loci of $\tilde{X}$ are indexed by a pair of size-$r$ subsets of $\{1,\ldots,n\}$, corresponding to choosing $r\times r$ minors of both $\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ and $\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$, i.e.
-they are indexed by elements of $D_{-}\times D_{+}$.
-Let $R_{1},R_{2}$ be the representations of $\tilde{G}$ obtained by composing the projection $\pi_{1},\pi_{2}:\tilde{G}\to G$ to $G$ with the right representation $R$ of $G$.
-Again we abuse notation and also use $R_{1/2}$ to denote the induced vector bundles on $\tilde{X}$.
+We will use this fact only to conclude that the map (2.3) is an isomorphism and preserves the pairing.
+We compute $\mathbb{FM}$ using localization as in [32].
+The fixed loci of $\tilde{X}$ are indexed by a pair of size-$r$ subsets of $\{1,\dots,n\}$, corresponding to choosing $r\times r$ minors of both $\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)$ and $\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})$, i.e., they are indexed by elements of $D_{-}\times D_{+}$.
+Let $R_{1}$, $R_{2}$ be the representations of $\tilde{G}$ obtained by composing the projection $\pi_{1},\pi_{2}\colon\tilde{G}\to G$ to $G$ with the right representation $R$ of $G$.
+Again, we abuse notation and also use $R_{1/2}$ to denote the induced vector bundles on $\tilde{X}$.
 With these conventions, we observe that $f_{-/+}^{*}(R)=R_{1/2}$.
@@
 For $(\delta^{-},\delta^{+})\in D_{-}\times D_{+}$, let $B_{(\delta^{-},\delta^{+})}$ denote the associated fixed locus, isomorphic to $B$.
-Then $f_{+}$ (resp.
-$f_{-}$) maps $B_{(\delta^{-},\delta^{+})}$ isomorphically onto $B_{\delta^{+}}$ (resp.
+Then $f_{+}$ ((((resp.
+$f_{-})$ maps $B_{(\delta^{-},\delta^{+})}$ isomorphically onto $B_{\delta^{+}}$ ((((resp.
 $B_{\delta^{-}})$.
-Furthermore, $(R_{1})|_{B_{(\delta^{-},\delta^{+})}}\cong\oplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}$ and $(R_{2})|_{B_{(\delta^{-},\delta^{+})}}\cong\oplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}$.
+Furthermore, $(R_{1})|_{B_{(\delta^{-},\delta^{+})}}\cong\bigoplus_{i=1}^{r}M_{\delta^{-}_{i}}^{\vee}$ and $(R_{2})|_{B_{(\delta^{-},\delta^{+})}}\cong\bigoplus_{j=1}^{r}L_{\delta^{+}_{j}}^{\vee}$.
 In order to ease notation, in this section we will also use $L_{i}$ or $M_{j}$ to denote the pullbacks of these respective line bundles to $X_{+}$, $X_{-}$, and $\tilde{X}.$ With this convention, for $\delta^{-}\in D_{-}$, let $e_{\delta^{-}}$ denote the class
-$e_{\delta^{-}}=$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\mathcal{H}om(M_{j_{-}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})^{\vee}$
-(2.3.2) $\cong$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}$
-in $K^{0}_{\mathbb{T}}(X_{-})$, where for $A\in K^{0}_{\mathbb{T}}(X_{-})$, we denote $\wedge^{\bullet}A:=\oplus_{i=0}^{\operatorname{rk}(A)}(-1)^{i}\wedge^{i}A$.
+$e_{\delta^{-}}=\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\mathcal{H}{\rm om}(M_{j_{-}},\mathcal{O}_{B}\otimes\mathbb{C}^{r})^{\vee}\cong\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}$ (2.4)
+in $K^{0}_{\mathbb{T}}(X_{-})$, where for $A\in K^{0}_{\mathbb{T}}(X_{-})$, we denote $\wedge^{\bullet}A:=\bigoplus_{i=0}^{\operatorname{rk}(A)}(-1)^{i}\wedge^{i}A$.
 Corollary 2.6.
-The $\mathbb{T}$-localized equivariant $K$-theory of $X_{-}$ is spanned by the classes $\{\pi_{-}^{*}(A)\otimes e_{\delta^{-}}|A\in K^{0}(B),\delta^{-}\in D_{-}\}$.
+The $\mathbb{T}$-localized equivariant $K$-theory of $X_{-}$ is spanned by the classes $\bigl{\{}\pi_{-}^{*}(A)\otimes e_{\delta^{-}}|A\in K^{0}(B),\delta^{-}\in D_{-}\bigr{\}}$.
 Proof.
 By Lemma 2.1, one checks that the class $\pi_{-}^{*}(A)\otimes e_{\delta^{-}}$ restricts to zero on $B_{\delta_{0}^{-}}$ for $\delta_{0}^{-}\neq\delta^{-}$.
-Furthermore, it restricts to the invertible class $\wedge^{\bullet}N_{B_{\delta^{-}}|\operatorname{Gr}(r,F^{\vee})}^{\vee}$ on $B_{\delta^{-}}$.
-The result then follows from the localization theorem in $K$-theory.
+Furthermore, generalizing the argument of [12, Section 5.1] to the relative setting, we see that the relative tangent bundle of $\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}$ is obtained from the following relative Euler sequence for $\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}$:
+$0\rightarrow S_{-}\otimes S_{-}^{\vee}\rightarrow S_{-}^{\vee}\otimes F^{\vee}\rightarrow T\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}/B\rightarrow 0$ (2.5)
+Therefore, $\pi_{-}^{*}(A)\otimes e_{\delta^{-}}$ restricts to $A\otimes\wedge^{\bullet}N_{B_{\delta^{-}}|\operatorname{Gr}(r,F^{\vee})}^{\vee}$ where $\wedge^{\bullet}N_{B_{\delta^{-}}|\operatorname{Gr}(r,F^{\vee})}^{\vee}$ is an invertible class on $B_{\delta^{-}}$.
+The result then follows from the localization theorem in $K$-theory as in [30, Section 2.3].
 ∎
@@
 We have
-(2.3.3) $\mathbb{FM}(e_{\delta^{-}})=\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}.$
+$\mathbb{FM}(e_{\delta^{-}})=\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}.$ (2.6)
 Remark 2.8.
-Note that although the right hand side of the equation above is the same formal expression as the right hand side of (2.3.2), the right hand side of (2.3.3) is to be interpreted as a class in $K^{0}_{\mathbb{T}}(X_{+})$.
+Note that although the right-hand side of the equation above is the same formal expression as the right-hand side of (2.4), the right-hand side of (2.6) is to be interpreted as a class in $K^{0}_{\mathbb{T}}\bigl{(}X_{+}\bigr{)}$.
 Proof.
-Note first that, in $K^{0}_{\mathbb{T}}(\tilde{X})$, the class
-$\prod_{i_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{i_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}\cdot\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}(R_{2}^{\vee}\otimes R_{1})^{\vee}$
-restricts to $\wedge^{\bullet}N_{B_{\delta^{-},\delta^{+}}|\tilde{X}}^{\vee}$ on $B_{(\delta^{-},\delta^{+})}$, and restricts to zero on $B_{(\delta^{-}_{1},\delta^{+}_{1})}$ for $(\delta^{-}_{1},\delta^{+}_{1})\neq(\delta^{-},\delta^{+})$.
+Note first that, in $K^{0}_{\mathbb{T}}\bigl{(}\tilde{X}\bigr{)}$, the class
+$\prod_{i_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{i_{-}}^{\vee}\otimes R_{1}^{\vee}\bigr{)}^{\vee}\cdot\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}\bigl{(}R_{2}^{\vee}\otimes R_{1}\bigr{)}^{\vee}$
+restricts to $\wedge^{\bullet}N_{B_{\delta^{-},\delta^{+}}|\tilde{X}}^{\vee}$ on $B_{(\delta^{-},\delta^{+})}$, and restricts to zero on $B_{(\delta^{-}_{1},\delta^{+}_{1})}$ for $\bigl{(}\delta^{-}_{1},\delta^{+}_{1}\bigr{)}\neq(\delta^{-},\delta^{+})$.
 It follows that
-$f_{-}^{*}(e_{\delta^{-}})=$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}$
-$=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\left(\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}}{\prod_{i_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{i_{-}}^{\vee}\otimes R_{1}^{\vee})^{\vee}\cdot\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}(R_{2}^{\vee}\otimes R_{1})^{\vee}}\right)$
-$=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\left(\frac{1}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}(R_{2}^{\vee}\otimes R_{1})^{\vee}}\right)$
-$=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\left(\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R_{2}^{\vee})^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}(M_{j}^{\vee}\otimes R_{2}^{\vee})^{\vee}}\right)$
-where the second equality is by the localization theorem applied to $K^{0}_{\mathbb{T}}(\tilde{X})$, the fourth is Lemma 2.4. Pushing forward via $f_{+}$, we obtain
-$(f_{+})_{*}f_{-}^{*}(e_{\delta^{-}})=$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{+}})_{*}i_{\delta^{+}}^{*}\left(\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R)^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}(M_{j}^{\vee}\otimes R^{\vee})^{\vee}}\right)$
-$=$ $\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}$
-where the last equality is the localization theorem applied to $K^{0}_{\mathbb{T}}(X_{+})$.
+$f_{-}^{*}(e_{\delta^{-}})=\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R_{1}^{\vee}\bigr{)}^{\vee}$
+${}=\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}$
+[…]
+${}=\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\Biggl{(}\frac{1}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\wedge^{\bullet}\bigl{(}R_{2}^{\vee}\otimes R_{1}\bigr{)}^{\vee}}\Biggr{)}$
+${}=\sum_{\delta^{+}\in D_{+}}(i_{\delta^{-},\delta^{+}})_{*}i_{\delta^{-},\delta^{+}}^{*}\Biggl{(}\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R_{2}^{\vee}\bigr{)}^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R_{2})^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}\bigl{(}M_{j}^{\vee}\otimes R_{2}^{\vee}\bigr{)}^{\vee}}\Biggr{)},$
+where the second equality is by the localization theorem applied to $K^{0}_{\mathbb{T}}\bigl{(}\tilde{X}\bigr{)}$, the fourth is Lemma 2.5. Pushing forward via $f_{+}$, we obtain
+$(f_{+})_{*}f_{-}^{*}(e_{\delta^{-}})={}$ $\sum_{\delta^{+}\in D_{+}}(i_{\delta^{+}})_{*}i_{\delta^{+}}^{*}\Biggl{(}\frac{\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}}{\prod_{j_{+}\notin\delta^{+}}\wedge^{\bullet}(L_{j_{+}}\otimes R)^{\vee}\cdot\prod_{1\leq j\leq n}\wedge^{\bullet}\bigl{(}M_{j}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}}\Biggr{)}$
+[…]
+where the last equality is the localization theorem applied to $K^{0}_{\mathbb{T}}(X_{+})$. ∎
 For $(\delta^{-},\delta^{+})\in D_{-}\times D_{+}$, define
-(2.3.4) $C_{\delta^{-},\delta^{+}}=$ $\prod_{i=1}^{r}e^{(n-r)(x_{\delta^{+}_{i}}-z_{\delta^{-}_{i}})/2}\prod_{j_{-}\notin\delta^{-}}\frac{\sin((x_{\delta^{+}_{i}}-z_{j_{-}})/2\mathbbm{i})}{\sin((z_{\delta^{-}_{i}}-z_{j_{-}})/2\mathbbm{i})},$
-where $\mathbbm{i}=\sqrt{-1}$.
+$C_{\delta^{-},\delta^{+}}={}$ $\prod_{i=1}^{r}{\rm e}^{(n-r)(x_{\delta^{+}_{i}}-z_{\delta^{-}_{i}})/2}\prod_{j_{-}\notin\delta^{-}}\frac{\sin\bigl{(}\bigl{(}x_{\delta^{+}_{i}}-z_{j_{-}}\bigr{)}/2{\rm i}\bigr{)}}{\sin\bigl{(}\bigl{(}z_{\delta^{-}_{i}}-z_{j_{-}}\bigr{)}/2{\rm i}\bigr{)}},$ (2.7)
+where ${\rm i}=\sqrt{-1}$.
 Definition 2.9.
-Define $\mathbb{U}_{H}:H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to be the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{\delta^{-}})\in H^{*}_{\mathbb{T}}(B_{\delta^{-}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to
-(2.3.5) $\sum_{\delta^{+}\in D_{+}}C_{\delta^{-},\delta^{+}}\frac{\phi_{\delta^{+},\delta^{-}}(\alpha)}{e_{\mathbb{T}}(N_{\delta^{+}})}$
-where $\phi_{\delta^{+},\delta^{-}}:H^{*}_{\mathbb{T}}(B_{\delta^{-}})\to H^{*}_{\mathbb{T}}(B_{\delta^{+}})$ is the canonical isomorphism induced by the projections $\pi_{\delta^{+}}:B_{\delta^{+}}\to B$ and $\pi_{\delta^{-}}:B_{\delta^{-}}\to B$.
+Define $\mathbb{U}_{H}\colon H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to be the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{\delta^{-}})\in H^{*}_{\mathbb{T}}(B_{\delta^{-}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ to
+$\sum_{\delta^{+}\in D_{+}}C_{\delta^{-},\delta^{+}}\frac{\phi_{\delta^{+},\delta^{-}}(\alpha)}{e_{\mathbb{T}}(N_{\delta^{+}})},$ (2.8)
+where $\phi_{\delta^{+},\delta^{-}}\colon H^{*}_{\mathbb{T}}(B_{\delta^{-}})\to H^{*}_{\mathbb{T}}(B_{\delta^{+}})$ is the canonical isomorphism induced by the projections $\pi_{\delta^{+}}\colon B_{\delta^{+}}\to B$ and $\pi_{\delta^{-}}\colon B_{\delta^{-}}\to B$.
 Proposition 2.10.
@@
 Proof.
-We will prove $\operatorname{ch}^{\mathbb{T}}(\mathbb{FM}(e_{\delta^{-}}))=\mathbb{U}_{H}(\operatorname{ch}^{\mathbb{T}}(e_{\delta^{-}}))$.
+We will prove $\operatorname{ch}^{\mathbb{T}}(\mathbb{FM}(e_{\delta^{-}}))=\mathbb{U}_{H}\bigl{(}\operatorname{ch}^{\mathbb{T}}(e_{\delta^{-}})\bigr{)}$.
 The general statement then follows by Corollary 2.6 and the projection formula.
 Observe that
-$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{-}}^{*}e_{\delta^{-}})=$ $\operatorname{ch}^{\mathbb{T}}\left(i_{\delta^{-}}^{*}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}\right)$
-$=$ $\operatorname{ch}^{\mathbb{T}}\left(\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes M_{\delta^{-}_{i}})^{\vee}\right)$
-$=$ $\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\left(1-e^{z_{\delta^{-}_{i}}-z_{j_{-}}}\right)$
+$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{-}}^{*}e_{\delta^{-}})$ $=\operatorname{ch}^{\mathbb{T}}\biggl{(}i_{\delta^{-}}^{*}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}\biggr{)}$
+$=\operatorname{ch}^{\mathbb{T}}\Biggl{(}\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes M_{\delta^{-}_{i}}\bigr{)}^{\vee}\Biggr{)}=\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\Bigl{(}1-{\rm e}^{z_{\delta^{-}_{i}}-z_{j_{-}}}\Bigr{)}$
 and
-$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{+}}^{*}\mathbb{FM}(e_{\delta^{-}}))=$ $\operatorname{ch}^{\mathbb{T}}i_{\delta^{+}}^{*}\left(\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes R^{\vee})^{\vee}\right)$
-$=$ $\operatorname{ch}^{\mathbb{T}}\left(\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}(M_{j_{-}}^{\vee}\otimes L_{\delta^{+}_{i}})^{\vee}\right)$
-$=$ $\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\left(1-e^{x_{\delta^{+}_{i}}-z_{j_{-}}}\right)$
+$\operatorname{ch}^{\mathbb{T}}(i_{\delta^{+}}^{*}\mathbb{FM}(e_{\delta^{-}}))$ $=\operatorname{ch}^{\mathbb{T}}i_{\delta^{+}}^{*}\biggl{(}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes R^{\vee}\bigr{)}^{\vee}\biggr{)}$
+$=\operatorname{ch}^{\mathbb{T}}\Biggl{(}\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\wedge^{\bullet}\bigl{(}M_{j_{-}}^{\vee}\otimes L_{\delta^{+}_{i}}\bigr{)}^{\vee}\Biggr{)}=\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\Bigl{(}1-{\rm e}^{x_{\delta^{+}_{i}}-z_{j_{-}}}\Bigr{)}.$
 The ratio of the two expressions above, after identifying $B_{\delta^{-}}$ and $B_{\delta^{+}}$, is
-$\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{\left(1-e^{x_{\delta^{+}_{i}}-z_{j_{-}}}\right)}{\left(1-e^{z_{\delta^{-}_{i}}-z_{j_{-}}}\right)}$
-$=$ $\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{e^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\left(e^{-(x_{\delta^{+}_{i}}-z_{j_{-}})/2}-e^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\right)}{e^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\left(e^{-(z_{\delta^{-}_{i}}-z_{j_{-}})/2}-e^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\right)}$
-$=$ $C_{\delta^{-},\delta^{+}}.$
-By the localization theorem in cohomology (see [AB84]), it follows that the diagram commutes. ∎
+$\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{\Bigl{(}1-{\rm e}^{x_{\delta^{+}_{i}}-z_{j_{-}}}\Bigr{)}}{\Bigl{(}1-{\rm e}^{z_{\delta^{-}_{i}}-z_{j_{-}}}\Bigr{)}}$
+$\qquad{}=\prod_{i=1}^{r}\prod_{j_{-}\notin\delta^{-}}\frac{{\rm e}^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\Bigl{(}{\rm e}^{-(x_{\delta^{+}_{i}}-z_{j_{-}})/2}-{\rm e}^{(x_{\delta^{+}_{i}}-z_{j_{-}})/2}\Bigr{)}}{{\rm e}^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\Bigl{(}{\rm e}^{-(z_{\delta^{-}_{i}}-z_{j_{-}})/2}-{\rm e}^{(z_{\delta^{-}_{i}}-z_{j_{-}})/2}\Bigr{)}}=C_{\delta^{-},\delta^{+}}.$
+By the localization theorem in cohomology (see [2]), it follows that the diagram commutes. ∎
 Corollary 2.11.
 The map $\mathbb{U}_{H}$ is invertible and has a well-defined non-equivariant limit. Furthermore, we have
-${\pi_{-}}_{*}\left(\alpha\cup\operatorname{Td}_{X_{-}/B}\right)={\pi_{+}}_{*}\left(\mathbb{U}_{H}(\alpha)\cup\operatorname{Td}_{X_{+}/B}\right).$
+${\pi_{-}}_{*}\bigl{(}\alpha\cup\operatorname{Td}_{X_{-}/B}\bigr{)}={\pi_{+}}_{*}\bigl{(}\mathbb{U}_{H}(\alpha)\cup\operatorname{Td}_{X_{+}/B}\bigr{)}.$
 Proof.
@@
 Let
-$\overline{\mathbb{FM}}:\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{-})\to\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{+})$
-denote the Fourier–Mukai transform in this case, where $\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{\pm})$ denotes the completion of $K^{0}_{\mathbb{T}}(\overline{X}_{\pm})$ along the augmentation ideal (see [EG00]).
-By definition of the $\mathbb{T}$ action on $B$, $E$, and $F$, the (completion of the) equivariant cohomology of $B$ is given by
-$\hat{H}^{*}_{\mathbb{T}}(B)=H^{*}(B)[\![z_{i},x_{i}]\!]_{i=1}^{n}$
-where recall that $x_{i}=c_{1}^{\mathbb{T}}(L_{i}^{\vee})$ and $z_{i}=c_{1}^{\mathbb{T}}(M_{i}^{\vee})$. The equivariant line bundles $L_{i}\to B$ and $M_{i}\to B$ for $1\leq i\leq n$ determine a map
-$v:[B/(\mathbb{C}^{*})^{2n}]\to[\{pt\}/(\mathbb{C}^{*})^{2n}].$
+$\overline{\mathbb{FM}}\colon\ \hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{-}\bigr{)}\to\hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{+}\bigr{)}$
+denote the Fourier–Mukai transform in this case, where $\hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}$ denotes the completion of $K^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}$ along the augmentation ideal (see [19]).
+By definition of the $\mathbb{T}$ action on $B$, $E$ and $F$, the (completion of the) equivariant cohomology of $B$ is given by
+$\hat{H}^{*}_{\mathbb{T}}(B)=H^{*}(B)[\![z_{i},x_{i}]\!]_{i=1}^{n},$
+where recall that $x_{i}=c_{1}^{\mathbb{T}}\bigl{(}L_{i}^{\vee}\bigr{)}$ and $z_{i}=c_{1}^{\mathbb{T}}\bigl{(}M_{i}^{\vee}\bigr{)}$.
+The equivariant line bundles $L_{i}\to B$ and $M_{i}\to B$ for $1\leq i\leq n$ determine a map
+$v\colon\ \big{[}B/(\mathbb{C}^{*})^{2n}\big{]}\to\big{[}\{pt\}/(\mathbb{C}^{*})^{2n}\big{]}.$
 This induces the following commutative diagram of stacks:
-${\left[X_{\pm}/(\mathbb{C}^{*})^{2n}\right]}$${\left[\overline{X}_{\pm}/(\mathbb{C}^{*})^{2n}\right]}$${\left[B/(\mathbb{C}^{*})^{2n}\right]}$[${\left[\{pt\}/(\mathbb{C}^{*})^{2n}\right].}$$\scriptstyle{\pi_{\pm}}$$\scriptstyle{\tilde{v}_{\pm}}$$\scriptstyle{\overline{\pi}_{\pm}}$$\scriptstyle{v}$
+${\bigl{[}X_{\pm}/(\mathbb{C}^{*})^{2n}\bigr{]}}$${\bigl{[}\overline{X}_{\pm}/(\mathbb{C}^{*})^{2n}\bigr{]}}$${\bigl{[}B/(\mathbb{C}^{*})^{2n}\bigr{]}}$[${\bigl{[}\{pt\}/(\mathbb{C}^{*})^{2n}\bigr{]}.}$$\scriptstyle{\pi_{\pm}}$$\scriptstyle{\tilde{v}_{\pm}}$$\scriptstyle{\overline{\pi}_{\pm}}$$\scriptstyle{v}$
 Taking cohomology, we obtain
@@
 $\hat{H}^{*}_{\mathbb{T}}(X_{\pm}):=\prod_{i=0}^{\infty}H^{i}_{\mathbb{T}}(X_{\pm})$
-with $H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{\pm},$ we have the identification
-${\pi_{\pm}}_{*}=\operatorname{id}_{B}\otimes_{\mathbb{C}}{\overline{\pi}_{\pm}}_{*}$
+with $H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\dots,\sigma_{r}]/\hat{I}_{\pm},$ we have the identification
+${\pi_{\pm}}_{*}=\operatorname{id}_{B}\otimes_{\mathbb{C}}{\overline{\pi}_{\pm}}_{*},$
 where $\operatorname{id}_{B}$ denotes the identity map on $H^{*}_{\mathbb{T}}(B)$.
 After taking completions with respect to the augmentation ideal, the Chern map
-$\hat{\operatorname{ch}}^{\mathbb{T}}:\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{\pm})\to\hat{H}^{*}_{\mathbb{T}}(\overline{X}_{\pm})$
-is an isomorphism (see [EG00]; here we use $\hat{\operatorname{ch}}^{\mathbb{T}}$ to denote the induced map defined on the completion $\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{\pm})$).
+$\hat{\operatorname{ch}}^{\mathbb{T}}\colon\ \hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}\to\hat{H}^{*}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}$
+is an isomorphism (see [19]; here we use $\hat{\operatorname{ch}}^{\mathbb{T}}$ to denote the induced map defined on the completion $\hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}$).
 Because it is induced by a derived equivalence, the map $\overline{\mathbb{FM}}$ is also an isomorphism.
-Proposition 2.9 implies that $\overline{\mathbb{U}}_{H}:=\hat{\operatorname{ch}}^{\mathbb{T}}\circ\overline{\mathbb{FM}}\circ(\hat{\operatorname{ch}}^{\mathbb{T}})^{-1}$ is also an isomorphism, and may be defined before localizing with respect to $\mathbb{T}$.
+Proposition 2.10 implies that $\overline{\mathbb{U}}_{H}:=\hat{\operatorname{ch}}^{\mathbb{T}}\circ\overline{\mathbb{FM}}\circ\bigl{(}\hat{\operatorname{ch}}^{\mathbb{T}}\bigr{)}^{-1}$ is also an isomorphism, and may be defined before localizing with respect to $\mathbb{T}$.
 In other words, $\overline{\mathbb{U}}_{H}$ gives a well-defined map
-$\overline{\mathbb{U}}_{H}:\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{-}\to\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{+}.$
-Again using the identification $\hat{H}^{*}_{\mathbb{T}}(X_{\pm})=H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{\pm}$ together with the explicit formula for $\mathbb{U}_{H}$ in (2.3.5), we also see that $\mathbb{U}_{H}=\operatorname{id}_{B}\otimes_{\mathbb{C}}\overline{\mathbb{U}}_{H}$ which implies the first statement.
-Given $\overline{\alpha}\in\hat{H}^{*}_{\mathbb{T}}(\overline{X}_{-})$, let $A=(\operatorname{ch}^{\mathbb{T}})^{-1}(\overline{\alpha})\in\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{-})$.
-Applying the equivariant Hirzebruch–Riemann–Roch theorem (see (3.1.2) below) twice, we see that for $A\in\hat{K}^{0}_{\mathbb{T}}(X_{-})$,
-(2.3.7) ${\overline{\pi}_{-}}_{*}\left(\overline{\alpha}\operatorname{Td}_{\overline{X}_{-}}\right)$ $=\chi(A)$
-$=\chi(\overline{\mathbb{FM}}(A))$
-$={\overline{\pi}_{+}}_{*}\left(\operatorname{ch}^{\mathbb{T}}(\overline{\mathbb{FM}}(A))\operatorname{Td}_{\overline{X}_{+}}\right)$
-$={\overline{\pi}_{+}}_{*}\left(\mathbb{U}_{H}(\overline{\alpha})\operatorname{Td}_{\overline{X}_{+}}\right).$
-The pullback $\tilde{v}_{\pm}^{*}(\operatorname{Td}_{\overline{X}_{\pm}})$ is the relative Todd class $\operatorname{Td}_{X_{\pm}}$.
-This fact together with (2.3.7) allows us to upgrade (2.3.6) to the following commutative diagram:
+$\overline{\mathbb{U}}_{H}\colon\ \mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\dots,\sigma_{r}]/\hat{I}_{-}\to\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\dots,\sigma_{r}]/\hat{I}_{+}.$
+Again using the identification $\hat{H}^{*}_{\mathbb{T}}(X_{\pm})=H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\dots,\sigma_{r}]/\hat{I}_{\pm}$ together with the explicit formula for $\mathbb{U}_{H}$ in (2.8), we also see that $\mathbb{U}_{H}=\operatorname{id}_{B}\otimes_{\mathbb{C}}\overline{\mathbb{U}}_{H}$ which implies the first statement.
+Given $\overline{\alpha}\in\hat{H}^{*}_{\mathbb{T}}\bigl{(}\overline{X}_{-}\bigr{)}$, let $A=\bigl{(}\operatorname{ch}^{\mathbb{T}}\bigr{)}^{-1}(\overline{\alpha})\in\hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{-}\bigr{)}$.
+Applying the equivariant Hirzebruch–Riemann–Roch theorem (see (3.2) below) twice, we see that for $A\in\hat{K}^{0}_{\mathbb{T}}(X_{-})$,
+${\overline{\pi}_{-}}_{*}\bigl{(}\overline{\alpha}\operatorname{Td}_{\overline{X}_{-}}\bigr{)}$ $=\chi(A)=\chi\bigl{(}\overline{\mathbb{FM}}(A)\bigr{)}={\overline{\pi}_{+}}_{*}\bigl{(}\operatorname{ch}^{\mathbb{T}}\bigl{(}\overline{\mathbb{FM}}(A)\bigr{)}\operatorname{Td}_{\overline{X}_{+}}\bigr{)}$
+$={\overline{\pi}_{+}}_{*}\bigl{(}\mathbb{U}_{H}(\overline{\alpha})\operatorname{Td}_{\overline{X}_{+}}\bigr{)}.$ (2.10)
+The pullback $\tilde{v}_{\pm}^{*}\bigl{(}\operatorname{Td}_{\overline{X}_{\pm}}\bigr{)}$ is the relative Todd class $\operatorname{Td}_{X_{\pm}}$.
+This fact together with (2.10) allows us to upgrade (2.9) to the following commutative diagram:
 […]
 The second claim follows. ∎
-3. Gromov–Witten theory and $I$-functions
+3 Gromov–Witten theory and $\bm{I}$-functions
 Let $X$ be a smooth semi-projective variety.
 In what follows, $X$ will also be equipped with an action by a torus $\mathbb{T}$.
-Following [CIJS15], we assume that all $\mathbb{T}$-weights appearing in the representation $H^{0}(X,\mathcal{O}_{X})$ lie in a strictly convex cone, and that $H^{0}(X,\mathcal{O}_{X})^{\mathbb{T}}=\mathbb{C}.$ This implies in particular that $X^{\mathbb{T}}$ is compact.
+Following [15], we assume that all $\mathbb{T}$-weights appearing in the representation $H^{0}(X,\mathcal{O}_{X})$ lie in a strictly convex cone, and that $H^{0}(X,\mathcal{O}_{X})^{\mathbb{T}}=\mathbb{C}.$ This implies in particular that $X^{\mathbb{T}}$ is compact.
 Following standard practice, we define integrals over $X$ via localization. For $\gamma\in H^{*}_{\mathbb{T}}(X)$, define
-$\int_{X}\gamma:=\sum_{F}\int_{F}\frac{(\gamma)|_{F}}{e_{\mathbb{T}}(N_{F|X})}\in S_{\mathbb{T}}$
+$\int_{X}\gamma:=\sum_{F}\int_{F}\frac{(\gamma)|_{F}}{e_{\mathbb{T}}\bigl{(}N_{F|X}\bigr{)}}\in S_{\mathbb{T}},$
 where the sum is over all connected components $F$ in $X^{\mathbb{T}}$ and the class $e_{\mathbb{T}}(-)$ denotes the $\mathbb{T}$-equivariant Euler class.
 Define a pairing $\left\langle-,-\right\rangle_{X}$ on $H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ by
-$\left\langle\alpha,\beta\right\rangle_{X}:=\sum_{F}\int_{F}\frac{(\alpha\cup\beta)|_{F}}{e_{\mathbb{T}}(N_{F|X})}.$
+$\left\langle\alpha,\beta\right\rangle_{X}:=\sum_{F}\int_{F}\frac{(\alpha\cup\beta)|_{F}}{e_{\mathbb{T}}\bigl{(}N_{F|X}\bigr{)}}.$
 For a ring $R$, let
-$R[\![Q]\!]=\left\{\sum_{d\in\operatorname{Eff}(X)}a_{d}Q^{d}\mid a_{d}\in R\right\}$
+$R[\![Q]\!]=\biggl{\{}\sum_{d\in\operatorname{Eff}(X)}a_{d}Q^{d}\mid a_{d}\in R\biggr{\}}$
 denote the Novikov ring of $X$. Let
-$\mathcal{H}_{X}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}(\!(z^{-1})\!)[\![Q]\!]$
+$\mathcal{H}_{X}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\bigl{(}\!\bigl{(}z^{-1}\bigr{)}\!\bigr{)}[\![Q]\!]$
 denote Givental’s symplectic vector space, equipped with the symplectic form
-(3.0.1) $\Omega(f,g)=\operatorname{Res}_{z=0}\left\langle f(-z),g(z)\right\rangle_{X}dz.$
+$\Omega(f,g)=\operatorname{Res}_{z=0}\left\langle f(-z),g(z)\right\rangle_{X}{\rm d}z.$
 This defines a polarization
 $\mathcal{H}_{X}=\mathcal{H}_{X}^{+}\oplus\mathcal{H}_{X}^{-}$
-with $\mathcal{H}_{X}^{+}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}[z][\![Q]\!]$ and $\mathcal{H}_{X}^{-}=z^{-1}H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}[\![z^{-1}]\!][\![Q]\!]$.
-For $\alpha_{1},\ldots,\alpha_{m}\in H^{*}_{\mathbb{T}}(X)$, $k_{1},\ldots,k_{m}\geq 0$, and $d\in\operatorname{Eff}(X)$, define the degree-$d$ genus-0 Gromov–Witten invariant of $X$:
-$\left\langle\alpha_{1}\psi_{1}^{k_{1}},\ldots,\alpha_{m}\psi_{m}^{k_{m}}\right\rangle_{0,m,d}^{X}=\int_{[\mathcal{M}_{0,m}(X,d)]^{vir}}\prod_{i=1}^{m}ev_{i}^{*}(\alpha_{i})\psi_{i}^{k_{i}}.$
-If $X$ is not compact, the integral on the right hand side is defined after localization.
+with $\mathcal{H}_{X}^{+}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}[z][\![Q]\!]$ and $\mathcal{H}_{X}^{-}=z^{-1}H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\bigl{[}\!\bigl{[}z^{-1}\bigr{]}\!\bigr{]}[\![Q]\!]$.
+For $\alpha_{1},\dots,\alpha_{m}\in H^{*}_{\mathbb{T}}(X)$, $k_{1},\dots,k_{m}\geq 0$, and $d\in\operatorname{Eff}(X)$, define the degree-$d$ genus-0 Gromov–Witten invariant of $X$:
+$\big{\langle}\alpha_{1}\psi_{1}^{k_{1}},\dots,\alpha_{m}\psi_{m}^{k_{m}}\bigr{\rangle}_{0,m,d}^{X}=\int_{[\mathcal{M}_{0,m}(X,d)]^{vir}}\prod_{i=1}^{m}ev_{i}^{*}(\alpha_{i})\psi_{i}^{k_{i}}.$
+If $X$ is not compact, the integral on the right-hand side is defined after localization.
 Definition 3.1.
 Define the $\mathbb{T}$-equivariant genus-zero descendent potential to be
-(3.0.2) $\mathcal{F}_{X}({\bm{t}}(z))=\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle{\bm{t}}(\psi_{1}),\ldots,{\bm{t}}(\psi_{m})\right\rangle_{0,m,d}^{X},$
+$\mathcal{F}_{X}({\bm{t}}(z))=\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle{\bm{t}}(\psi_{1}),\dots,{\bm{t}}(\psi_{m})\right\rangle_{0,m,d}^{X},$
 for ${\bm{t}}(z)=\sum_{a=0}^{s}t_{a}z^{a}\in H^{*}_{\mathbb{T}}(X)[z]$. Extending linearly, $\mathcal{F}_{X}$ defines a function on $\mathcal{H}_{X}^{+}$.
@@
 After applying the dilaton shift $\sum_{a=0}^{s}t_{a}z^{a}=z1+\sum_{a=0}^{s}q_{a}z^{a}$, we define Givental’s overruled Lagrangian cone $\mathcal{L}_{X}$ to be the graph of $\mathcal{F}_{X}$ with respect to the above polarization:
-$\mathcal{L}_{X}=\{(p,q)\in T^{*}\mathcal{H}_{X}^{+}\mid p=d_{q}\mathcal{F}_{X}\}.$
+$\mathcal{L}_{X}=\bigl{\{}(p,q)\in T^{*}\mathcal{H}_{X}^{+}\mid p=d_{q}\mathcal{F}_{X}\bigr{\}}.$
 Let $\{\phi_{i}\}_{i\in I}$ be a basis for $H^{*}_{\mathbb{T}}(X)$ and let $\{\phi^{i}\}_{i\in I}$ denote the dual basis. Points of $\mathcal{L}_{X}$ are of the form
-$-z1+{\bm{t}}(z)+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},{\bm{t}}(\psi_{2}),\ldots,{\bm{t}}(\psi_{m+1})\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
+$-z1+{\bm{t}}(z)+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},{\bm{t}}(\psi_{2}),\dots,{\bm{t}}(\psi_{m+1})\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
 Definition 3.3.
 The $J$-function of $X$ is the slice of $\mathcal{L}_{X}$ obtained by intersecting with $-z1+t+\mathcal{H}_{X}^{-}$ for $t=\sum_{i\in I}t^{i}\phi_{i}\in H^{*}_{\mathbb{T}}(X)$:
-$J^{X}(t,z)=-z1+t+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\sum_{i\in I}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},t,\ldots,t\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
+$J^{X}(t,z)=-z1+t+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\sum_{i\in I}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},t,\dots,t\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
 The function $J^{X}(t,z)$ determines the entire Lagrangian cone. Via the string equation, dilaton equation, and topological recursion relations, any point on $\mathcal{L}_{X}$ may be expressed as
-$J^{X}(\tau,z)+\sum_{i\in I}C^{i}(z)\frac{\partial}{\partial{t^{i}}}J^{X}(t,z)|_{t=\tau}$
+$J^{X}(\tau,z)+\sum_{i\in I}C^{i}(z)\frac{\partial}{\partial{t^{i}}}J^{X}(t,z)|_{t=\tau},$
 where $C^{i}(z)\in S_{\mathbb{T}}[z][\![Q]\!]$ are polynomials in $z$.
-3.1. Integral lattice
+3.1 Integral lattice
+As predicted by mirror symmetry, the genus zero Gromov–Witten theory ($A$-model) of a space $X$ should be equivalent to the Gauss–Manin connection ($B$-model) of a family of mirror manifolds $\check{X_{t}}$.
+The integral cohomology $H^{*}\bigl{(}\check{X_{t}},\mathbb{Z}\bigr{)}$ naturally endows the Gauss–Manin connection with an integral lattice of flat sections.
+Thus by mirror symmetry, one should expect an integral structure to appear in Gromov–Witten theory as well.
+Such a structure was defined by Iritani in [22] where it was verified to be compatible with mirror symmetry in the toric setting.
+In this section, we describe Iritani’s integral structure on $\tilde{H}_{X}$.
+We define a lattice in $\tilde{H}_{X}$ as the image of $K_{\mathbb{T}}(X)$ under a certain modification of the Chern character.
+A key ingredient is a characteristic class known as the $\hat{\Gamma}$ class, which plays the role of a square root of the Todd class.
+This integral structure will allow us to compare the symplectic transformation $\mathbb{U}$ with the Fourier–Mukai transform $\mathbb{FM}$ from Section 2.3.
 For $\phi\in H^{2p}_{\mathbb{T}}(X)$ a homogeneous element, let
@@
 First, we define $\mu$ to be the operator
-$\mu(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X)}{2}\right)\phi,$
+$\mu(\phi)=\biggl{(}\frac{\deg(\phi)}{2}-\frac{\dim(X)}{2}\biggr{)}\phi,$
 where $\dim$ denotes the dimension over $\mathbb{C}$.
-Secondly, we define $\deg_{0}:H^{*}_{\mathbb{T}}(X)\to H^{*}_{\mathbb{T}}(X)$ be the degree operator, defined by
+Secondly, we define $\deg_{0}\colon H^{*}_{\mathbb{T}}(X)\to H^{*}_{\mathbb{T}}(X)$ be the degree operator, defined by
 $\deg_{0}(\phi)=\deg(\phi)\cdot\phi.$
-Finally, we define $\rho$ denote the operator of multiplication by $c_{1}(T_{X})$.
+Finally, we define $\rho$ to be the operator of multiplication by $c_{1}(T_{X})$.
 For $A\to X$ a vector bundle, define the multiplicative class $\hat{\Gamma}(A)$ to be
-$\hat{\Gamma}(A)=\prod_{i=1}^{s}\Gamma(1+a_{i})$
-where $a_{1},\ldots,a_{s}$ are the Chern roots of $A$, and $\Gamma(z)$ is the Gamma function expanded as a power series at $z=0$. Denote $\hat{\Gamma}(TX)$ by $\hat{\Gamma}_{X}$.
+$\hat{\Gamma}(A)=\prod_{i=1}^{s}\Gamma(1+a_{i}),$
+where $a_{1},\dots,a_{s}$ are the Chern roots of $A$, and $\Gamma(z)$ is the Gamma function expanded as a power series at $z=0$. Denote $\hat{\Gamma}(TX)$ by $\hat{\Gamma}_{X}$.
 Define the map
-$\Psi:K^{0}_{\mathbb{T}}(X)$$\to\widetilde{\mathcal{H}}_{X}$
-$A$ $\mapsto z^{-\mu}z^{\rho}\left(\hat{\Gamma}_{X}\cup(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\right)$
+$\Psi\colon\ K^{0}_{\mathbb{T}}(X)\to\widetilde{\mathcal{H}}_{X},\qquad A\mapsto z^{-\mu}z^{\rho}\bigl{(}\hat{\Gamma}_{X}\cup(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\bigr{)},$
 where
-~$\widetilde{\mathcal{H}}_{X}=\mathcal{H}_{X}[\log z][z^{-1/2}].$
+~$\widetilde{\mathcal{H}}_{X}=\mathcal{H}_{X}[\log z]\big{[}z^{-1/2}\big{]}.$
 The image of $\Psi$ defines a lattice in ~$\widetilde{\mathcal{H}}_{X}$. For future use, we will also define the map
-(3.1.1) $\psi:\hat{H}^{*}_{\mathbb{T}}(X)$$\to\widetilde{\mathcal{H}}_{X}$
-$\phi$ $\mapsto z^{-\mu}z^{\rho}\left(\hat{\Gamma}_{X}\cup(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\phi\right)$
+$\psi\colon\ \hat{H}^{*}_{\mathbb{T}}(X)\to\widetilde{\mathcal{H}}_{X},\qquad\phi\mapsto z^{-\mu}z^{\rho}\bigl{(}\hat{\Gamma}_{X}\cup(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\phi\bigr{)}$ (3.1)
 so that $\Psi=\psi\circ\operatorname{ch}^{\mathbb{T}}$.
 Define the equivariant Euler characteristic to be
-$\chi(A):=\sum_{i=0}^{\dim(X)}(-1)^{i}\operatorname{ch}^{\mathbb{T}}(H^{i}(X,A))$
-where, as before, we use the equivariant Chern character with respect to $\mathbb{T}$. If $X$ is compact, then by [EG00, Tho92] we have the following equivariant Hirzebruch–Riemann–Roch theorem
-(3.1.2) $\chi(A)=\int_{X}\operatorname{ch}^{\mathbb{T}}(A)\operatorname{Td}_{X},$
+$\chi(A):=\sum_{i=0}^{\dim(X)}(-1)^{i}\operatorname{ch}^{\mathbb{T}}\bigl{(}H^{i}(X,A)\bigr{)},$
+where, as before, we use the equivariant Chern character with respect to $\mathbb{T}$. If $X$ is compact, then by [19, 32] we have the following equivariant Hirzebruch–Riemann–Roch theorem
+$\chi(A)=\int_{X}\operatorname{ch}^{\mathbb{T}}(A)\operatorname{Td}_{X},$ (3.2)
 as an equality in $\hat{S}_{\mathbb{T}}$ (as defined in Section 2.1).
@@
 Then the projection formula in cohomology implies
-$\int_{X}\alpha=\int_{Z}s^{*}(\alpha)/e_{\mathbb{T}}(V)$
-where $s:Z\to V$ denotes the zero section. On the other hand, the projection formula in $K$-theory implies
-$\chi(A)=\chi(s^{*}(A)/\wedge^{\bullet}(V^{\vee})).$
-Applying the Hirzebruch–Riemann–Roch theorem (3.1.2) to $Z$ then implies (3.1.2) for $X$.
-Following [CIJ18], consider a modified Euler characteristic $\chi_{z}$ given by
-$\chi_{z}(C,D):=\left(\frac{2\pi\mathbbm{i}}{z}\right)^{\lambda\partial_{\lambda}}\chi(C,D)$
-where $\lambda\partial_{\lambda}:=\sum_{i}\lambda_{i}\frac{\partial}{\partial\lambda_{i}}$ for some basis $\lambda_{1},\ldots,\lambda_{s}$ of $H^{2}(B\mathbb{T})$, .
-Lemma 3.4 (Proposition 2.10 of [Iri09]).
+$\int_{X}\alpha=\int_{Z}s^{*}(\alpha)/e_{\mathbb{T}}(V),$
+where $s\colon Z\to V$ denotes the zero section. On the other hand, the projection formula in $K$-theory implies
+$\chi(A)=\chi\bigl{(}s^{*}(A)/\wedge^{\bullet}\bigl{(}V^{\vee}\bigr{)}\bigr{)}.$
+Applying the Hirzebruch–Riemann–Roch theorem (3.2) to $Z$ then implies (3.2) for $X$.
+Following [14], consider a modified Euler characteristic $\chi_{z}$ given by
+$\chi_{z}(C,D):=\biggl{(}\frac{2\pi{\rm i}}{z}\biggr{)}^{\lambda\partial_{\lambda}}\chi(C,D),$
+where $\lambda\partial_{\lambda}:=\sum_{i}\lambda_{i}\frac{\partial}{\partial\lambda_{i}}$ for some basis $\lambda_{1},\dots,\lambda_{s}$ of $H^{2}(B\mathbb{T})$.
+Lemma 3.4 ([22, Proposition 2.10]).
 The map $\Psi$ relates the pairings in $K$-theory and cohomology as follows:
@@
 Proof.
-Following [CIJ18, Proposition 3.2], we use the following equalities: $z^{-\lambda d\lambda}\left\langle\alpha,\beta\right\rangle_{X}=\left\langle z^{-\mu}\alpha,z^{\mu}\beta\right\rangle_{X}$, $\left\langle z^{\rho}\alpha,z^{\rho}\beta\right\rangle_{X}=\left\langle\alpha,\beta\right\rangle_{X}$, and $e^{2\pi\mathbbm{i}\mu}\rho=-\rho e^{2\pi\mathbbm{i}\mu}$.
-The proof then follows from the equivariant Hirzebruch–Riemann–Roch theorem.
-$\chi_{z}(C,D)=z^{-\lambda\partial\lambda}\frac{1}{(2\pi\mathbbm{i})^{\dim X}}\int_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C^{\vee}\otimes D)\operatorname{Td}_{X}$
-$=z^{-\lambda\partial\lambda}\frac{1}{(2\pi\mathbbm{i})^{\dim X}}\int_{X}e^{\pi\mathbbm{i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C^{\vee}\otimes D)$
-$=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\int_{X}e^{-\pi\mathbbm{i}\dim(X)/2}e^{\pi\mathbbm{i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C^{\vee}\otimes D)$
-$=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\left\langle e^{\pi\mathbbm{i}\rho}e^{\pi\mathbbm{i}\mu}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\right\rangle_{X}$
-$=\frac{1}{(2\pi)^{\dim X}}\left\langle z^{-\mu}z^{-\rho}e^{\pi\mathbbm{i}\rho}e^{\pi\mathbbm{i}\mu}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\right\rangle_{X}$
-$=\frac{1}{(2\pi)^{\dim X}}\left\langle(e^{-\pi\mathbbm{i}}z)^{-\mu}(e^{-\pi\mathbbm{i}}z)^{\rho}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\right\rangle_{X}$
-$=\frac{1}{(2\pi)^{\dim X}}\left\langle\Psi(C)(e^{-\pi\mathbbm{i}}z),\Psi(D)(z)\right\rangle_{X}$
+Following [14, Proposition 3.2], we use the following equalities:
+$z^{-\lambda{\rm d}\lambda}\left\langle\alpha,\beta\right\rangle_{X}=\left\langle z^{-\mu}\alpha,z^{-\mu}\beta\right\rangle_{X},\left\langle z^{\rho}\alpha,z^{\rho}\beta\right\rangle_{X}=\left\langle\alpha,\beta\right\rangle_{X}\qquad\text{and}\qquad{\rm e}^{2\pi{\rm i}\mu}\rho=-\rho{\rm e}^{2\pi{\rm i}\mu}.$
+The proof then follows from the equivariant Hirzebruch–Riemann–Roch theorem,
+$\chi_{z}(C,D)=z^{-\lambda\partial\lambda}\frac{1}{(2\pi{\rm i})^{\dim X}}\int_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}\bigl{(}C^{\vee}\otimes D\bigr{)}\operatorname{Td}_{X}$
+$\qquad{}=z^{-\lambda\partial\lambda}\frac{1}{(2\pi{\rm i})^{\dim X}}\int_{X}{\rm e}^{\pi{\rm i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}\bigl{(}C^{\vee}\otimes D\bigr{)}$
+$\qquad{}=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\int_{X}{\rm e}^{-\pi{\rm i}\dim(X)/2}{\rm e}^{\pi{\rm i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}\bigl{(}C^{\vee}\otimes D\bigr{)}$
+$\qquad{}=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\bigl{\langle}{\rm e}^{\pi{\rm i}\rho}{\rm e}^{\pi{\rm i}\mu}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\bigr{\rangle}_{X}$
+$\qquad{}=\frac{1}{(2\pi)^{\dim X}}\bigl{\langle}z^{-\mu}z^{-\rho}{\rm e}^{\pi{\rm i}\rho}{\rm e}^{\pi{\rm i}\mu}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\bigr{\rangle}_{X}$
+$\qquad{}=\frac{1}{(2\pi)^{\dim X}}\bigl{\langle}\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)}^{-\mu}\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)}^{\rho}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\bigr{\rangle}_{X}$
+$\qquad{}=\frac{1}{(2\pi)^{\dim X}}\bigl{\langle}\Psi(C)\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)},\Psi(D)(z)\bigr{\rangle}_{X}.$
 Let us return now to the specific geometry of the relative Grassmann flop of the previous section.
 Let $X_{+}$ and $X_{-}$ be as in Section 2.
-Let $\psi_{\pm}$ denote (3.1.1) for these particular spaces.
+Let $\psi_{\pm}$ denote (3.1) for these particular spaces.
 Definition 3.5.
-Define $\mathbb{U}:\mathcal{H}_{X_{-}}\to\mathcal{H}_{X_{+}}$ to be
-$\mathbb{U}=\psi_{+}\circ\mathbb{U}_{H}\circ\psi_{-}^{-1},$
-where $\mathbb{U}_{H}$ is the map in (2.3.5).
+Define $\mathbb{U}\colon\mathcal{H}_{X_{-}}\to\mathcal{H}_{X_{+}}$ to be $\mathbb{U}=\psi_{+}\circ\mathbb{U}_{H}\circ\psi_{-}^{-1}$, where $\mathbb{U}_{H}$ is the map in (2.8).
 Corollary 3.6.
@@
 Proof.
-By the Leray–Hirsch theorem, any class in $\hat{H}^{*}_{\mathbb{T}}(X_{-})$ may be decomposed as a sum of classes of the form $\pi_{-}^{*}(\beta)\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C})$ for some $\overline{C}\in K^{0}_{\mathbb{T}}(\overline{X}_{-})$ where the map $\tilde{v}_{-}$ appears in the proof of Corollary 2.10.
-We may therefore assume without loss of generality that $f(z)=\psi_{-}(\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C}))$ and $g(z)=\psi_{-}(\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{D}))$.
-Using the same chain of equalities from the previous proof, the projection formula, and (2.3.8), one computes that
-$\left\langle\psi_{-}(\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C}))(e^{-\pi\mathbbm{i}}z),\psi_{-}(\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{D}))(z)\right\rangle_{X_{-}}$
-$=$ $K\cdot\chi_{z}(\overline{C},\overline{D})$
+By the Leray–Hirsch theorem, any class in $\hat{H}^{*}_{\mathbb{T}}(X_{-})$ may be decomposed as a sum of classes of the form $\pi_{-}^{*}(\beta)\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C})$ for some $\overline{C}\in K^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{-}\bigr{)}$, where the map $\tilde{v}_{-}$ appears in the proof of Corollary 2.11.
+We may therefore assume without loss of generality that $f(z)=\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{C}\bigr{)}\bigr{)}$ and $g(z)=\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{D}\bigr{)}\bigr{)}$.
+Using the same chain of equalities from the previous proof, the projection formula, and (2.11), one computes that
+$\bigl{\langle}\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{C}\bigr{)}\bigr{)}\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)},\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{D}\bigr{)}\bigr{)}(z)\bigr{\rangle}_{X_{-}}=K\cdot\chi_{z}\bigl{(}\overline{C},\overline{D}\bigr{)}$
 and
-$\left\langle\mathbb{U}\psi_{-}(\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C}))(e^{-\pi\mathbbm{i}}z),\mathbb{U}\psi_{-}(\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{D}))(z)\right\rangle_{X_{+}}$
-$=$ $K\cdot\chi_{z}(\overline{\mathbb{FM}}(\overline{C}),\overline{\mathbb{FM}}(\overline{D}))$
+$\bigl{\langle}\mathbb{U}\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{C}\bigr{)}\bigr{)}\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)},\mathbb{U}\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{D}\bigr{)}\bigr{)}(z)\bigr{\rangle}_{X_{+}}$
+$\qquad{}=K\cdot\chi_{z}\bigl{(}\overline{\mathbb{FM}}\bigl{(}\overline{C}\bigr{)},\overline{\mathbb{FM}}\bigl{(}\overline{D}\bigr{)}\bigr{)},$
 where
-$K=\left(z^{-\lambda\partial\lambda}\frac{(2\pi)^{\dim X}}{(2\pi\mathbbm{i})^{\dim B}}\int_{B}\left((-2\pi\mathbbm{i})^{\deg_{0}/2}\beta_{1}\right)\cup\left((2\pi\mathbbm{i})^{\deg_{0}/2}\beta_{2}\operatorname{Td}_{B}\right)\right).$
+$K=\biggl{(}z^{-\lambda\partial\lambda}\frac{(2\pi)^{\dim X}}{(2\pi{\rm i})^{\dim B}}\int_{B}\bigl{(}(-2\pi{\rm i})^{\deg_{0}/2}\beta_{1}\bigr{)}\cup\bigl{(}(2\pi{\rm i})^{\deg_{0}/2}\beta_{2}\operatorname{Td}_{B}\bigr{)}\biggr{)}.$
 Because $\overline{\mathbb{FM}}$ is induced from an equivalence of categories,
-$\chi_{z}(\overline{C},\overline{D})=\chi_{z}(\overline{\mathbb{FM}}(\overline{C}),\overline{\mathbb{FM}}(\overline{D})).$
-The claim follows.
-3.2. $I$-functions for Grassmann bundles
-In this section, we record $I$-functions for $X_{\pm}$ and the associated abelian quotients. In each case an $I$-function is obtained by modifying the $J$-function of the base $B$.
-3.2.1. $r=1$
+$\chi_{z}\bigl{(}\overline{C},\overline{D}\bigr{)}=\chi_{z}\bigl{(}\overline{\mathbb{FM}}\bigl{(}\overline{C}\bigr{)},\overline{\mathbb{FM}}\bigl{(}\overline{D}\bigr{)}\bigr{)}.$
+The claim follows. ∎
+3.2 $\bm{I}$-functions for Grassmann bundles
+In this section, we record $I$-functions for $X_{\pm}$ and the associated abelian quotients. In each case, an $I$-function is obtained by modifying the $J$-function of the base $B$.
+3.2.1 $\bm{r=1}$
 We begin with the case $r=1$, in which
-$X_{+}=$ $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)$
-$X_{-}=$ $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right).$
+$X_{+}=\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)},\qquad X_{-}=\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)}.$
 In this case $X_{\pm}$ is a toric bundle over $B$.
-$I$-functions for split toric bundles were computed by Brown in [Bro14] in terms of the $J$-function for the base $B$.
+$I$-functions for split toric bundles were computed by Brown in [7] in terms of the $J$-function for the base $B$.
 Let $J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}$ be the $J$-function for $B$.
 Define the $H^{*}_{\mathbb{T}}(X_{\pm})$-valued function $I_{X_{\pm}}(Q,q,z)$ to be
-(3.2.1) $I_{X_{+}}(Q,q_{+},z)=$ $\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{e\in\mathbb{Z}}q_{+}^{-y/z+e}M_{\beta,-e}$
-(3.2.2) $I_{X_{-}}(Q,q_{-},z)=$ $\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{d\in\mathbb{Z}}q_{-}^{y/z+d}M_{\beta,d}$
+$I_{X_{+}}(Q,q_{+},z)=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{e\in\mathbb{Z}}q_{+}^{-y/z+e}M_{\beta,-e},$
+$I_{X_{-}}(Q,q_{-},z)=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{d\in\mathbb{Z}}q_{-}^{y/z+d}M_{\beta,d},$
 where
-(3.2.3) $M_{\beta,d}=\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y+hz)}{\prod_{h=-\infty}^{d-\beta\cdot M_{i}}(z_{i}+y+hz)}\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y+hz)}{\prod_{h=-\infty}^{-d+\beta\cdot L_{k}}(-x_{k}-y+hz)}.$
+$M_{\beta,d}=\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y+hz)}{\prod_{h=-\infty}^{d-\beta\cdot M_{i}}(z_{i}+y+hz)}\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y+hz)}{\prod_{h=-\infty}^{-d+\beta\cdot L_{k}}(-x_{k}-y+hz)}.$ (3.3)
 Remark 3.7.
 Despite having the same formal expression under the change of variables $q_{+}=q_{-}^{-1}$, we emphasize that $I_{X_{+}}$ and $I_{X_{-}}$ take very different forms due to the fact that they lie in different cohomology rings.
-In particular, in $I_{X_{+}}$, the expression $M_{\beta,-e}$ vanishes whenever $e<\operatorname{min}_{k}(-\beta\cdot L_{k})$ by (2.2.1).
-On the other hand in $I_{X_{-}}$ the expression $M_{\beta,d}$ vanishes when $d<\operatorname{min}_{i}(\beta\cdot M_{i})$.
-The following is a special case of Brown’s result for toric bundles (see [Bro14]).
+In particular, in $I_{X_{+}}$, the expression $M_{\beta,-e}$ vanishes whenever $e<\operatorname{min}_{k}(-\beta\cdot L_{k})$ by (2.2).
+On the other hand, in $I_{X_{-}}$ the expression $M_{\beta,d}$ vanishes when $d<\operatorname{min}_{i}(\beta\cdot M_{i})$.
+The following is a special case of Brown’s result for toric bundles (see [7]).
 Theorem 3.8.
 The function $I_{X_{\pm}}(Q,q_{\pm},-z)$ lies in the $\mathbb{T}$-equivariant Lagrangian cone $\mathcal{L}_{X_{\pm}}$.
-3.2.2. The associated abelian quotients
+3.2.2 The associated abelian quotients
 Now assume $r>1$. By iterating the result of the previous section, we obtain $I$-functions for $X_{T,\pm}$, and for all $X_{T,j}$ for $0\leq j\leq s$. Let
-$L_{\beta,\vec{d}}=$ $\prod_{j=1}^{r}\left(\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}.\right)$
-Define $I_{X_{T,j}}(Q,d_{1},\ldots,d_{j},e_{j+1},\ldots,e_{r},z)$ to be
-(3.2.4) $I_{X_{T,j}}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{\stackrel{{\scriptstyle(e_{1},\ldots,e_{j})\in\mathbb{Z}^{j}}}{{(d_{j+1},\ldots,d_{r})\in\mathbb{Z}^{s-j}}}}\prod_{i=1}^{j}q_{i,+}^{-y_{i}/z+e_{i}}\prod_{i=j+1}^{r}q_{i,-}^{y_{i}/z+d_{i}}L_{\beta,d_{1},\ldots,d_{j},-e_{j+1},\ldots,-e_{r}}.$
-Again by [Bro14], we see that $I_{X_{T,j}}(Q,d_{1},\ldots,d_{j},e_{j+1},\ldots,e_{r},-z)$ lies on the Lagrangian cone for $X_{T,j}$.
-3.2.3. The non-abelian quotients
+$L_{\beta,\vec{d}}={}$ $\prod_{j=1}^{r}\Biggl{(}\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}\Biggr{)}.$
+Define $I_{X_{T,j}}(Q,d_{1},\dots,d_{j},e_{j+1},\dots,e_{r},z)$ to be
+$I_{X_{T,j}}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{\stackrel{{\scriptstyle(e_{1},\dots,e_{j})\in\mathbb{Z}^{j}}}{{(d_{j+1},\dots,d_{r})\in\mathbb{Z}^{s-j}}}}\prod_{i=1}^{j}q_{i,+}^{-y_{i}/z+e_{i}}\prod_{i=j+1}^{r}q_{i,-}^{y_{i}/z+d_{i}}L_{\beta,d_{1},\dots,d_{j},-e_{j+1},\dots,-e_{r}}.$
+Again by [7], we see that $I_{X_{T,j}}(Q,d_{1},{\dots},d_{j},e_{j+1},{\dots},e_{r},-z)$ lies on the Lagrangian cone for $X_{T,j}$.
+3.2.3 The non-abelian quotients
 The abelian/non-abelian correspondence in Gromov–Witten theory is the principle that one can recover the Gromov–Witten theory of the non-abelian quotient from that of the associated abelian quotient.
-There is a long history of results in this direction, e.g.
-[BCFK05, BCFK08, CFKS08, Web23, Web21, CLS22].
+There is a long history of results in this direction, e.g., [4, 5, 13, 17, 33, 34].
 The modern formulation of the expected relationship is, roughly, that the (genus zero) Gromov–Witten theory for the non-abelian quotient may be obtained from that of the associated abelian quotient by the following steps:
-  1. (1)
+  1. 1)
     twisting the theory of the abelian quotient by the Euler class of the root bundle $\mathfrak{g}/\mathfrak{t}$;
-  2. (2)
+  2. 2)
     specializing those Novikov parameters which are identified via the natural map
@@
 Let
-$N_{\beta,\vec{d}}=\prod_{j=1}^{r}$ $\left(\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}\right)$
-(3.2.5) $\cdot$ $\prod_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{r}\frac{\prod_{h=-\infty}^{d_{i}-d_{j}}(y_{i}-y_{j}+hz)}{\prod_{h=-\infty}^{0}(y_{i}-y_{j}+hz)}.$
+$N_{\beta,\vec{d}}=\prod_{j=1}^{r}\Biggl{(}\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}\Biggr{)}$
+×${}\times\prod_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{r}\frac{\prod_{h=-\infty}^{d_{i}-d_{j}}(y_{i}-y_{j}+hz)}{\prod_{h=-\infty}^{0}(y_{i}-y_{j}+hz)}.$ (3.4)
 Define $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,\vec{q}_{+},z)$ and $I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,\vec{q}_{-},z)$ to be
-$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\ldots,-e_{r}}$
-$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\ldots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\ldots,d_{r}}.$
+$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\dots,-e_{r}},$
+$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\dots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\dots,d_{r}}.$
 Then, define
-$I_{X_{\pm}}(Q,q_{\pm},z)=I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,\vec{q}_{+},z)|_{q_{1,\pm}=\cdots=q_{s,\pm}=q_{\pm}}.$
-The following proposition is a corollary of the work of Coates–Lutz–Shafi in [CLS22] on the abelian/non-abelian correspondence.
-One could also prove it directly by generalizing the work of Oh in [Oh21] on $I$-functions of Grassmann bundles.
+$I_{X_{\pm}}(Q,q_{\pm},z)=I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,\vec{q}_{\pm},z)|_{q_{1,\pm}=\cdots=q_{s,\pm}=q_{\pm}}.$
+The following proposition is a corollary of Coates–Lutz–Shafi in [17] on the abelian/non-abelian correspondence.
+One could also prove it directly by generalizing the work of Oh in [29] on $I$-functions of Grassmann bundles.
 Proposition 3.9.
@@
 Proof.
-As described in [CLS22], $I_{X_{\pm}}$ is the Givental–Martin modification of $I_{X_{T,\pm}}$, so the proposition essentially follows from [CLS22, Theorem 1.11].
-More precisely, we apply the argument of Theorem 5.11 of [CLS22], which provides an abelian/non-abelian correspondence for twisted theories over a flag bundle.
+As described in [17], $I_{X_{\pm}}$ is the Givental–Martin modification of $I_{X_{T,\pm}}$, so the proposition essentially follows from [17, Theorem 1.11].
+More precisely, we apply the argument of [17, Theorem 5.11], which provides an abelian/non-abelian correspondence for twisted theories over a flag bundle.
 Because that theorem is not stated in exactly the generality which we need here, we summarize the argument below and explain the necessary adjustments to the case at hand.
@@
 The space $X_{+}$ is the total space of a vector bundle over the Grassmann bundle
-$Z:=\operatorname{Gr}(r,E)=\operatorname{tot}\left(\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}G.$
+$Z:=\operatorname{Gr}(r,E)=\operatorname{tot}\left(\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}G.$
 The associated abelian quotient is the toric bundle given by the $r$-fold fiber product of $\mathbb{P}(E)$ over $B$:
-$Z_{T}:=\operatorname{tot}\left(\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}T=\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E).$
-Let $I_{Z_{T}}$ denote the Brown $I$-function for $Z_{T}$ as in [Bro14], and let $I_{Z_{T}}^{e_{\mathbb{T}}^{-1}((S_{T}\otimes p^{*}F^{\vee}))}$ denote the corresponding $((S_{T}\otimes p^{*}F^{\vee}),e^{-1}_{\mathbb{T}}(-))$-twisted $I$-function, where we use $S_{T}$ to denote the vector bundle on $Z_{T}$ induced by the right standard representation $R$ from (2.1.1) restricted to $T\subset G$.
+$Z_{T}:=\operatorname{tot}\left(\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E)\right)\sslash_{+}T=\mathbb{P}(E)\times_{B}\cdots\times_{B}\mathbb{P}(E).$
+Let $I_{Z_{T}}$ denote the Brown $I$-function for $Z_{T}$ as in [7], and let $I_{Z_{T}}^{e_{\mathbb{T}}^{-1}((S_{T}\otimes p^{*}F^{\vee}))}$ denote the corresponding $\bigl{(}\bigl{(}S_{T}\otimes p^{*}F^{\vee}\bigr{)},{\rm e}^{-1}_{\mathbb{T}}(-)\bigr{)}$-twisted $I$-function, where we use $S_{T}$ to denote the vector bundle on $Z_{T}$ induced by the right standard representation $R$ from (2.1) restricted to $T\subset G$.
 Then $I_{Z_{T}}^{e_{\mathbb{T}}^{-1}((S_{T}\otimes p^{*}F^{\vee}))}$ is exactly the $I$-function $I_{X_{T,+}}$ from the previous section.
-Following the notation and discussion in Section 5.2 of [CLS22] (in particular Theorem 5.11), we apply a further twist to $I_{X_{T,+}}=I_{Z_{T}}^{e^{-1}_{\mathbb{T}}((S_{T}\otimes p^{*}F^{\vee})}$ by $(\Phi,\mathbf{c}^{\prime})$ where $\Phi\to Z_{T}$ is the root bundle with respect to $G$ and $\mathbf{c}^{\prime}$ is the $\mathbb{C}^{*}$-equivariant Euler class with respect to the $\mathbb{C}^{*}$-action scaling the fibers of $\Phi.$ Denote the equivariant parameter of this $\mathbb{C}^{*}$ action by $\lambda$.
-Specializing variables to $q_{1,+}=\cdots=q_{s,+}=q_{+},$ projecting via the map on Givental spaces defined in [CLS22]
-$p:\mathcal{H}_{Z_{T}}^{W}\to\mathcal{H}_{Z},$
+Following the notation and discussion in [17, Section 5.2] (in particular, Theorem 5.11), we apply a further twist to
+$I_{X_{T,+}}=I_{Z_{T}}^{{\rm e}^{-1}_{\mathbb{T}}((S_{T}\otimes p^{*}F^{\vee})}$
+by $(\Phi,\mathbf{c}^{\prime})$ where $\Phi\to Z_{T}$ is the root bundle with respect to $G$ and $\mathbf{c}^{\prime}$ is the $\mathbb{C}^{*}$-equivariant Euler class with respect to the $\mathbb{C}^{*}$-action scaling the fibers of $\Phi$.
+Denote the equivariant parameter of this $\mathbb{C}^{*}$ action by $\lambda$.
+Specializing variables to $q_{1,+}=\cdots=q_{s,+}=q_{+},$ projecting via the map on Givental spaces defined in [17]
+$p\colon\ \mathcal{H}_{Z_{T}}^{W}\to\mathcal{H}_{Z},$
 and taking the nonequivariant limit $\lambda\mapsto 0$, we obtain $I_{X_{+}}(Q,q_{+},z)$.
-By the proof of [CLS22, Theorem 5.11], this function lies in the $((S_{+}\otimes p^{*}F^{\vee}),e^{-1}_{\mathbb{T}})$-twisted Lagrangian cone of $Z=\operatorname{Gr}(r,E)$.
+By the proof of [17, Theorem 5.11], this function lies in the $\bigl{(}\bigl{(}S_{+}\otimes p^{*}F^{\vee}\bigr{)},{\rm e}^{-1}_{\mathbb{T}}\bigr{)}$-twisted Lagrangian cone of $Z=\operatorname{Gr}(r,E)$.
 While the theorem in loc.
 cit.
-is only stated for the twisted theory $(L,e_{\mathbb{C}^{*}}(-))$, where $L$ is a line bundle and $e_{\mathbb{C}^{*}}(-)$ is an equivariant Euler class, by [CLS22, Remark 3.7] the result holds for other multiplicative classes as well.
+is only stated for the twisted theory $(L,e_{\mathbb{C}^{*}}(-))$, where $L$ is a line bundle and $e_{\mathbb{C}^{*}}(-)$ is an equivariant Euler class, by [17, Remark 3.7] the result holds for other multiplicative classes as well.
 The further extension to the case that $L$ is replaced by the sum of line bundles also follows the same argument.
-By definition, the $((S_{+}\otimes p^{*}F^{\vee}),e^{-1}_{\mathbb{T}})$-twisted Lagrangian cone of $\operatorname{Gr}(r,E)$ is equal to the $\mathbb{T}$-equivariant Lagrangian cone of $X_{+}$.
-4. Analytic continuation: the abelian case
-In this section we explicitly compute the analytic continuation of the $I$-functions of the associated abelian quotients $X_{T,\pm}$ and show that they are related by symplectic transformation.
-This was proven in [CIJ18] in the case $B$ is a point.
-4.1. Projective bundles
+By definition, the $\bigl{(}\bigl{(}S_{+}\otimes p^{*}F^{\vee}\bigr{)},{\rm e}^{-1}_{\mathbb{T}}\bigr{)}$-twisted Lagrangian cone of $\operatorname{Gr}(r,E)$ is equal to the $\mathbb{T}$-equivariant Lagrangian cone of $X_{+}$.
+∎
+4 Analytic continuation: The abelian case
+In this section, we explicitly compute the analytic continuation of the $I$-functions of the associated abelian quotients $X_{T,\pm}$ and show that they are related by symplectic transformation.
+This was proven in [14] in the case $B$ is a point.
+4.1 Projective bundles
 Let us consider the special case of $r=1$ first.
 We will analytically continue the function $I_{X_{+}}$ to $q_{+}=\infty$, and show that the resulting function is related to $I_{X_{-}}$ by a symplectic transformation.
-We begin by writing $I_{X_{\pm}}$ as $\psi(H_{X_{\pm}})$ where $\psi$ is as in (3.1.1). The relative Euler sequence for $X_{\pm}$ is
-$0\to\mathcal{O}_{X_{\pm}}\to R\otimes p^{*}E\oplus R^{\vee}\otimes p^{*}F^{\vee}\to TX_{\pm}/B\to 0$
-where $R$ is the vector bundle defined in Section 2.1. Because the Gamma class is multiplicative, we have $\hat{\Gamma}$, we have that
-$\hat{\Gamma}_{X_{\pm}}=$ $\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)$
-$=$ $\hat{\Gamma}_{B}\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y))\prod_{i=1}^{n}\Gamma(1+(z_{i}+y)).$
-Note that the hypergeometric modification (3.2.3) is equal to
-(4.1.1) $\frac{1}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}\prod_{j=1}^{n}\frac{\Gamma(1+(-x_{j}-y)/z)\Gamma(1+(z_{j}+y)/z)}{\Gamma(1+(-x_{j}-y)/z+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/z-\beta\cdot M_{j}+d)}$
+We begin by writing $I_{X_{\pm}}$ as $\psi(H_{X_{\pm}})$ where $\psi$ is as in (3.1). By (2.5) in the case $r=1$ we obtain the relative Euler sequence
+$0\to\mathcal{O}_{X_{-}}\to S_{-}\otimes p^{*}E\oplus S_{-}^{\vee}\otimes p^{*}F^{\vee}\to TX_{-}/B\to 0$
+and similarly for $X_{+}$. Because the Gamma class is multiplicative, we have that
+$\hat{\Gamma}_{X_{\pm}}=\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)=\hat{\Gamma}_{B}\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y))\prod_{i=1}^{n}\Gamma(1+(z_{i}+y)).$
+Note that the hypergeometric modification (3.3) is equal to
+$\frac{1}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}\prod_{j=1}^{n}\frac{\Gamma(1+(-x_{j}-y)/z)\Gamma(1+(z_{j}+y)/z)}{\Gamma(1+(-x_{j}-y)/z+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/z-\beta\cdot M_{j}+d)}.$
 Define $H^{B}(Q,z)$ to be
-$H^{B}$ $:=(2\pi\mathbbm{i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}$
-$=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
-where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}(TB\oplus E\oplus F^{\vee})$ and
-$\mu_{X}(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\right)\phi.$
-Define the $H^{*}_{\mathbb{T}}(X_{+})$-valued function
-(4.1.2) $H_{X_{+},\beta}(q_{+})=$
-$\sum_{e\in\mathbb{Z}}\frac{q_{+}^{-y/2\pi\mathbbm{i}+e}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi\mathbbm{i}+\beta\cdot L_{j}+e)\Gamma(1+(z_{j}+y)/2\pi\mathbbm{i}-\beta\cdot M_{j}-e)}.$
+$H^{B}:=(2\pi{\rm i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
+where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}\bigl{(}TB\oplus E\oplus F^{\vee}\bigr{)}$ and
+$\mu_{X}(\phi)=\biggl{(}\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\biggr{)}\phi.$
+Define the $H^{*}_{\mathbb{T}}\bigl{(}X_{+}\bigr{)}$-valued function
+$H_{X_{+},\beta}(q_{+})$
+$\qquad{}=\sum_{e\in\mathbb{Z}}\frac{q_{+}^{-y/2\pi{\rm i}+e}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi{\rm i}+\beta\cdot L_{j}+e)\Gamma(1+(z_{j}+y)/2\pi{\rm i}-\beta\cdot M_{j}-e)}.$
 Similarly, define the $H^{*}_{\mathbb{T}}(X_{-})$-valued function
-(4.1.3) $H_{X_{-},\beta}(q_{-})=$
-$\sum_{d\in\mathbb{Z}}\frac{q_{-}^{y/2\pi\mathbbm{i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/2\pi\mathbbm{i}-\beta\cdot M_{j}+d)}.$
+$H_{X_{-},\beta}(q_{-})$ (4.1)
+$\qquad{}=\sum_{d\in\mathbb{Z}}\frac{q_{-}^{y/2\pi{\rm i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi{\rm i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}+y)/2\pi{\rm i}-\beta\cdot M_{j}+d)}.$
 Then a simple check shows that
-(4.1.4) $I_{X_{\pm}}(Q,q_{\pm},z)$ $=\psi\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{\pm},\beta}\right).$
+$I_{X_{\pm}}(Q,q_{\pm},z)$ ${}=\psi\biggl{(}\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{\pm},\beta}\biggr{)}.$ (4.2)
 We will fix $\beta$ and relate $H_{X_{+},\beta}$ and $H_{X_{-},\beta}$ via analytic continuation and symplectic transformation.
 Recall Definition 2.9 in the special case $r=1$.
-In this case $\mathbb{U}_{H}:H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{B_{k}|X_{-}})\in H^{*}_{\mathbb{T}}(B_{k}^{-})$ to
-(4.1.5) $\sum_{l=1}^{n}e^{(n-1)\cdot(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2\mathbbm{i})}\phi_{kl}(\alpha)/e_{\mathbb{T}}(N_{B_{l}|X_{+}})$
-where $\phi_{kl}:H^{*}_{\mathbb{T}}(B_{k}^{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(B_{l}^{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{k}^{+}:B_{k}^{+}\to B$ and $\pi_{l}^{-}:B_{l}^{-}\to B$.
-By restricting $H_{X_{+},\beta}$ to the fixed points of the $\mathbb{T}$-action and using Mellin–Barnes integrals, we may analytically continue $H_{X_{+},\beta}$ to $q_{+}=\infty.$ Because the method of computation is by now somewhat standard (see e.g.
-[CIJ18]), we relegate the proof to Appendix A.
+In this case, $\mathbb{U}_{H}\colon H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the linear map defined by sending $\alpha/e_{\mathbb{T}}(N_{B_{k}|X_{-}})\in H^{*}_{\mathbb{T}}(B_{k}^{-})$ to
+$\sum_{l=1}^{n}{\rm e}^{(n-1)\cdot(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2{\rm i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2{\rm i})}\phi_{kl}(\alpha)/e_{\mathbb{T}}(N_{B_{l}|X_{+}}),$ (4.3)
+where $\phi_{kl}\colon H^{*}_{\mathbb{T}}(B_{k}^{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(B_{l}^{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{k}^{+}\colon B_{k}^{+}\to B$ and $\pi_{l}^{-}\colon B_{l}^{-}\to B$.
+By restricting $H_{X_{+},\beta}$ to the fixed points of the $\mathbb{T}$-action and using Mellin–Barnes integrals, we may analytically continue $H_{X_{+},\beta}$ to $q_{+}=\infty.$ Because the method of computation is by now somewhat standard (see, e.g., [14]), we relegate the proof to Appendix A.
 We obtain the following.
@@
 We have the equality
-(4.1.6) $\mathbb{U}_{H}H_{X_{-},\beta}=\widetilde{H_{X_{+},\beta}}$
+$\mathbb{U}_{H}H_{X_{-},\beta}=\widetilde{H_{X_{+},\beta}}$ (4.4)
 for all $\beta$, where $\widetilde{H_{X_{+},\beta}}$ denotes the analytic continuation of $H_{X_{+},\beta}$ along the path described above, under the substitution $q_{+}^{-1}=q_{-}$.
@@
 where $\widetilde{I_{X_{+}}}$ denotes the analytic continuation of $I_{X_{+}}$ along the path $\gamma$ and we set $q_{+}^{-1}=q_{-}$. Furthermore, the following diagram commutes:
-${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\mathbb{U}}$
+${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\mathbb{U}}$
 Proof.
-The proof of the equality is immediate from Definition 3.5, (4.1.4) and (4.1.6).
+The proof of the equality is immediate from Definition 3.5, (4.2) and (4.4).
 The fact that $\mathbb{U}$ is symplectic and has a well-defined non-equivariant limit is Corollary 3.6.
@@
 ∎
-4.2. Products of projective bundles
+4.2 Products of projective bundles
 We next consider, for general $r>0$, the wall crossing between $I$-functions of the associated abelian quotients $X_{T,\pm}$.
@@
 Observe that the wall crossing from $X_{T,j}$ to $X_{T,j+1}$ is a particular example of the wall crossing from the previous section.
-In this case for any $j$ we have the relationship between Gamma functions:
+In this case, for any $j$ we have the relationship between Gamma functions:
 $\hat{\Gamma}(TX_{T,j})=\hat{\Gamma}_{B}\hat{\Gamma}(TX_{T}/B)$
 with
-$\hat{\Gamma}(TX_{T}/B)=\prod_{k=1}^{r}\prod_{j=1}^{n}\left(\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\right).$
+$\hat{\Gamma}(TX_{T}/B)=\prod_{k=1}^{r}\prod_{j=1}^{n}\bigl{(}\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\bigr{)}.$
 Define $H^{B}_{T}(Q,z)$ to be
-$H^{B}_{T}$ $:=(2\pi\mathbbm{i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X_{T}}}z^{\mu_{X_{T}}}J^{B}$
-$=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,T,\beta},$
-where $\rho_{X_{T}}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}(TB\oplus(E\oplus F^{\vee})^{\oplus r})$ and
-$\mu_{X_{T}}(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X_{T,\pm})}{2}\right)\phi.$
+$H^{B}_{T}:=(2\pi{\rm i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X_{T}}}z^{\mu_{X_{T}}}J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,T,\beta},$
+where $\rho_{X_{T}}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}\bigl{(}TB\oplus\bigl{(}E\oplus F^{\vee}\bigr{)}^{\oplus r}\bigr{)}$ and
+$\mu_{X_{T}}(\phi)=\biggl{(}\frac{\deg(\phi)}{2}-\frac{\dim(X_{T,\pm})}{2}\biggr{)}\phi.$
 Define the $H^{*}_{\mathbb{T}}(X_{T,+})$-valued, resp. $H^{*}_{\mathbb{T}}(X_{T,-})$-valued, functions
 $H_{X_{T,+},\beta}=$
-$\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{q_{k,+}^{(-y_{k}/2\pi\mathbbm{i}+e_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})}$
+$\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\!\frac{q_{k,+}^{(-y_{k}/2\pi{\rm i}+e_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(y_{k}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})},$
 $H_{X_{T,-},\beta}=$
-$\sum_{(d_{1},\ldots,d_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{q_{k,-}^{(y_{k}/2\pi\mathbbm{i}+d_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}+d_{k})}$
+$\sum_{(d_{1},\dots,d_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\!\frac{q_{k,-}^{(y_{k}/2\pi{\rm i}+d_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}+d_{k})}.$
 Then as in the case of $r=1$, we have
-(4.2.1) $I_{X_{T,\pm}}(Q,q_{\pm},z)$ $=\psi_{X_{T,\pm}}\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}H_{B,T,\beta}H_{X_{T,\pm},\beta}\right)$
+$I_{X_{T,\pm}}(Q,q_{\pm},z)$ ${}=\psi_{X_{T,\pm}}\biggl{(}\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}H_{B,T,\beta}H_{X_{T,\pm},\beta}\biggr{)}.$
 Recall that the $\mathbb{T}$-fixed loci of $X_{T,\pm}$ are indexed by $F_{\pm}$.
-Let $\mathbb{U}_{H}^{T}:H^{*}_{\mathbb{T}}(X_{T,-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{T,+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined by sending $\alpha/e_{\mathbb{T}}(N_{f^{-}})\in H^{*}_{\mathbb{T}}(B_{f^{-}})$ to
-$\sum_{f^{+}\in F_{+}}\prod_{i=1}^{r}e^{(n-1)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+x_{f^{+}_{i}})/2\mathbbm{i})}{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+z_{f^{-}_{i}})/2\mathbbm{i})}\frac{\phi_{f^{+},f^{-}}(\alpha)}{e_{\mathbb{T}}(N_{f^{+}})}$
-where $\phi_{f^{+},f^{-}}:H^{*}_{\mathbb{T}}(B_{f^{-}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(B_{f^{+}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{f^{+}}:B_{f^{+}}\to B$ and $\pi_{f^{-}}:B_{f^{-}}\to B$, and $N_{f^{\pm}}$ is the normal bundle $N_{B_{f^{\pm}}|X_{\pm}}$.
+Let $\mathbb{U}_{H}^{T}\colon H^{*}_{\mathbb{T}}(X_{T,-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{T,+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined by sending $\alpha/e_{\mathbb{T}}\bigl{(}N_{f^{-}}\bigr{)}\in H^{*}_{\mathbb{T}}\bigl{(}B_{f^{-}}\bigr{)}$ to
+$\sum_{f^{+}\in F_{+}}\prod_{i=1}^{r}{\rm e}^{(n-1)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin\bigl{(}\bigl{(}-z_{j}+x_{f^{+}_{i}}\bigr{)}/2{\rm i}\bigr{)}}{\prod_{j\neq{f^{-}_{i}}}\sin\bigl{(}\bigl{(}-z_{j}+z_{f^{-}_{i}}\bigr{)}/2{\rm i}\bigr{)}}\frac{\phi_{f^{+},f^{-}}(\alpha)}{e_{\mathbb{T}}(N_{f^{+}})},$
+where $\phi_{f^{+},f^{-}}\colon H^{*}_{\mathbb{T}}\bigl{(}B_{f^{-}}\bigr{)}\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}\bigl{(}B_{f^{+}}\bigr{)}\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{f^{+}}\colon B_{f^{+}}\to B$ and $\pi_{f^{-}}\colon B_{f^{-}}\to B$, and $N_{f^{\pm}}$ is the normal bundle $N_{B_{f^{\pm}}|X_{\pm}}$.
 A straightforward generalization of the analytic continuation in Appendix A (essentially repeating it $r$ times, one for each $q_{i}$ variable) yields the following:
 $\mathbb{U}_{H}^{T}H_{X_{T,-},\beta}=\widetilde{H_{X_{T,+},\beta}}$
-for all $\beta$, after identifying $q_{k,+}^{-1}=q_{k,-}$. Here the path of analytic continuation is a concatenation
-$\gamma_{1}^{T}\star\cdots\star\gamma_{r}^{T}$
-where $\gamma_{k}^{T}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$, and in the $\log(q_{k,+})$ plane follows the path $\gamma$ of analytic continuation from the previous section.
-Define
-$\mathbb{U}^{T}=\psi_{X_{T,+}}\circ\mathbb{U}_{H}^{T}\circ\psi_{X_{T,-}}^{-1}$
-in analogy to the previous section. By Theorem 4.2, we obtain:
+for all $\beta$, after identifying $q_{k,+}^{-1}=q_{k,-}$.
+Here the path of analytic continuation is a concatenation $\gamma_{1}^{T}\star\cdots\star\gamma_{r}^{T}$, where $\gamma_{k}^{T}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$, and in the $\log(q_{k,+})$ plane follows the path $\gamma$ of analytic continuation from the previous section.
+Define $\mathbb{U}^{T}=\psi_{X_{T,+}}\circ\mathbb{U}_{H}^{T}\circ\psi_{X_{T,-}}^{-1}$ in analogy to the previous section. By Theorem 4.2, we obtain the following.
 Theorem 4.3.
-The linear transformation $\mathbb{U}^{T}$ is symplectic, has a well-defined non-equivariant limit, and satisfies:
+The linear transformation $\mathbb{U}^{T}$ is symplectic, has a well-defined non-equivariant limit, and satisfies
 $\mathbb{U}^{T}I_{X_{T,-}}=\widetilde{I_{X_{T,+}}}.$
-5. Grassmann flops
-In this section we compute the analytic continuation and symplectic transformation for the $I$-functions $X_{+}$ and $X_{-}$ appearing in the Grassmann flop.
-5.1. Analytic continuation
+5 Grassmann flops
+In this section, we compute the analytic continuation and symplectic transformation for the $I$-functions $X_{+}$ and $X_{-}$ appearing in the Grassmann flop.
+5.1 Analytic continuation
 The computation proceeds by first computing the analytic continuation for the twisted $I$-functions of the abelian quotients, similar to the previous section, and then deforming the path to the locus $q_{1,+}=\cdots=q_{r,+}$.
-Let $S_{-}$ denote the rank-$r$ tautological bundle on $X_{-}$. We have the exact sequence
+Let $S_{-}$ denote the rank-$r$ tautological bundle on $X_{-}$. By (2.5), we have the short exact sequence
 $0\to S_{-}\otimes S_{-}^{\vee}\to S_{-}\otimes p^{*}E\oplus S_{-}^{\vee}\otimes p^{*}F^{\vee}\to TX_{-}/B\to 0$
 and similarly for $X_{+}$.
-We view $H^{*}_{\mathbb{T}}(X_{\pm})$ as a subspace of $(H^{*}_{\mathbb{T}}(V^{s}(G)/T))$ via Proposition 2.3.
-After pulling back to $V^{s}(G)/T$, we may write the Gamma class $\hat{\Gamma}_{X_{\pm}}$ as $\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)$ where in this case
-$\hat{\Gamma}(TX_{\pm}/B)=\frac{\prod_{k=1}^{r}\prod_{j=1}^{n}\left(\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\right)}{\hat{\Gamma}(\mathfrak{g}/\mathfrak{t})}$
+We view $H^{*}_{\mathbb{T}}(X_{\pm})$ as a subspace of $\bigl{(}H^{*}_{\mathbb{T}}(V^{s}(G)/T)\bigr{)}$ via Proposition 2.3.
+After pulling back to $V^{s}(G)/T$, we may write the Gamma class $\hat{\Gamma}_{X_{\pm}}$ as $\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)$, where in this case
+$\hat{\Gamma}(TX_{\pm}/B)=\frac{\prod_{k=1}^{r}\prod_{j=1}^{n}\bigl{(}\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\bigr{)}}{\hat{\Gamma}(\mathfrak{g}/\mathfrak{t})}$
 with
@@
 Recall that the $I$-functions for $X_{\pm}$ are specializations at $q_{1,\pm}=\cdots=q_{r,\pm}$ of $I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ where
-$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\ldots,-e_{r}}$
-$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\ldots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\ldots,d_{r}}.$
-with $N_{\beta,\vec{d}}$ defined in (3.2.5).
+$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\dots,-e_{r}},$
+$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\dots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\dots,d_{r}}$
+with $N_{\beta,\vec{d}}$ defined in (3.4).
 The hypergeometric factor $N_{\beta,\vec{d}}$ can be rewritten as
-$\frac{1}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}\prod_{i\neq k}\frac{\Gamma(1+(y_{k}-y_{i})/z+d_{k}-d_{i})}{\Gamma(1+(y_{k}-y_{i})/z)}\cdot$
-$\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{\hat{\Gamma}(1+(-y_{k}-x_{j})/z)\Gamma(1+(y_{k}+z_{j})/z)}{\Gamma(1+(-y_{k}-x_{j})/z+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/z-\beta\cdot M_{j}+d_{k})},$
+$\frac{1}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}\prod_{i\neq k}\frac{\Gamma(1+(y_{k}-y_{i})/z+d_{k}-d_{i})}{\Gamma(1+(y_{k}-y_{i})/z)}$
+×$\qquad{}\times\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{\hat{\Gamma}(1+(-y_{k}-x_{j})/z)\Gamma(1+(y_{k}+z_{j})/z)}{\Gamma(1+(-y_{k}-x_{j})/z+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/z-\beta\cdot M_{j}+d_{k})}.$
 Generalizing the definition of $H^{B}(Q,z)$ to the case of $r>1$, we define
-$H^{B}$ $:=(2\pi\mathbbm{i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}$
-$=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
-where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}(TB\oplus(E\oplus F^{\vee})^{\oplus r})$ and $\mu_{X}$ acts on a homogeneous element $\phi$ by
-$\mu_{X}(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\right)\phi.$
+$H^{B}:=(2\pi{\rm i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
+where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}\bigl{(}TB\oplus\bigl{(}E\oplus F^{\vee}\bigr{)}^{\oplus r}\bigr{)}$ and $\mu_{X}$ acts on a homogeneous element $\phi$ by
+$\mu_{X}(\phi)=\biggl{(}\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\biggr{)}\phi.$
 Denote by $H_{X_{+},\beta}$ the function
-$\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\left(q_{k,+}^{(-y_{k}/2\pi\mathbbm{i}+e_{k})}\prod_{i<k}\Gamma(1+(y_{k}-y_{i})/2\pi\mathbbm{i}-e_{k}+e_{i})\Gamma(1-(y_{k}-y_{i})/2\pi\mathbbm{i}+e_{k}-e_{i})\right.$
-$\left.\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})^{-1}\right)$
-$=$ $\frac{\pi^{r(r-1)}}{\prod_{i<k\leq r}\sin((y_{k}-y_{i})/2\mathbbm{i})}\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}$
-$\prod_{k=1}^{r}\left(q_{k,+}^{(-y_{k}/2\pi\mathbbm{i}+e_{k})}\prod_{i<k}((y_{k}-y_{i})/2\pi\mathbbm{i}-e_{k}+e_{i})\right.$
-$\left.\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})^{-1}\right)$
-As before we have that
-(5.1.1) $I_{X_{+}}(Q,q_{+},z)$ $=\psi_{+}\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{+},\beta}\right)\vline_{q_{1,+}=\cdots=q_{r,+}=q_{+}}.$
+$\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\!\Biggl{(}\!q_{k,+}^{(-y_{k}/2\pi{\rm i}+e_{k})}\!\prod_{i<k}\!\Gamma(1+(y_{k}-y_{i})/2\pi{\rm i}-e_{k}+e_{i})\Gamma(1-(y_{k}-y_{i})/2\pi{\rm i}+e_{k}-e_{i})\!$
+×$\quad\quad{}\times\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})^{-1}\!\Biggr{)}$
+$\quad{}=\frac{\pi^{r(r-1)/2}}{\prod_{i<k\leq r}\sin((y_{k}-y_{i})/2{\rm i})}\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}$
+×$\qquad{}\times\prod_{k=1}^{r}\Biggl{(}q_{k,+}^{(-y_{k}/2\pi{\rm i}+e_{k})}\prod_{i<k}((y_{k}-y_{i})/2\pi{\rm i}-e_{k}+e_{i})$
+×$\qquad{}\times\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})^{-1}\Biggr{)}.$
+As before, we have that
+$I_{X_{+}}(Q,q_{+},z)$ ${}=\psi_{+}\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{+},\beta}\right)\vline_{q_{1,+}=\cdots=q_{r,+}=q_{+}}.$ (5.1)
 An analogous statement holds for $X_{-}$.
-Restrict $H_{X_{+},\beta}$ to a fixed locus $B_{\delta}$, where $\delta\in D_{+}$ is a size $r$ subset of $\{1,\ldots,n\}$. Then
-(5.1.2) $H^{+}_{\beta,\delta}:=H_{X_{+},\beta}|_{B_{\delta}}=\frac{\pi^{r(r-1)}}{\prod_{i<k\leq r}\sin((x_{\delta_{i}}-x_{\delta_{k}})/2\mathbbm{i})}K^{+}_{\beta,\delta}$
+Restrict $H_{X_{+},\beta}$ to a fixed locus $B_{\delta}$, where $\delta\in D_{+}$ is a size $r$ subset of $\{1,\dots,n\}$. Then
+$H^{+}_{\beta,\delta}:=H_{X_{+},\beta}|_{B_{\delta}}=\frac{\pi^{r(r-1)/2}}{\prod_{i<k\leq r}\sin((x_{\delta_{i}}-x_{\delta_{k}})/2{\rm i})}K^{+}_{\beta,\delta},$ (5.2)
 where $K^{+}_{\beta,\delta}$ is
-$\sum_{e_{1}\geq-\beta\cdot L_{\delta_{1}}}\cdots\sum_{e_{r}\geq-\beta\cdot L_{\delta_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}\prod_{k=1}^{r}\left(q_{k,+}^{(x_{\delta_{k}}/2\pi\mathbbm{i}+e_{k})}\prod_{i<k}((-x_{\delta_{k}}+x_{\delta_{i}})/2\pi\mathbbm{i}-e_{k}+e_{i})\right.$
-$\left.\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})^{-1}\right).$
+$\sum_{e_{1}\geq-\beta\cdot L_{\delta_{1}}}\cdots\sum_{e_{r}\geq-\beta\cdot L_{\delta_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}\prod_{k=1}^{r}\Biggl{(}q_{k,+}^{(x_{\delta_{k}}/2\pi{\rm i}+e_{k})}\prod_{i<k}((-x_{\delta_{k}}+x_{\delta_{i}})/2\pi{\rm i}-e_{k}+e_{i})$
+×$\quad{}\times\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})^{-1}\Biggr{)}.$
 For any $f\in F_{-}$, define $K^{-}_{\beta,f}$ to be
-$\sum_{d_{1}\geq\beta\cdot M_{f_{1}}}\cdots\sum_{d_{r}\geq\beta\cdot M_{f_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}d_{k}}\prod_{k=1}^{r}\left(q_{k,-}^{(-z_{f_{k}}/2\pi\mathbbm{i}+d_{k})}\prod_{i<k}((-z_{f_{k}}+z_{f_{i}})/2\pi\mathbbm{i}+d_{k}-d_{i})\right.$
-$\left.\prod_{j=1}^{n}\Gamma(1+(z_{f_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d_{k})^{-1}\Gamma(1+(-z_{f_{k}}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}+d_{k})^{-1}\right).$
+$\sum_{d_{1}\geq\beta\cdot M_{f_{1}}}\cdots\sum_{d_{r}\geq\beta\cdot M_{f_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}d_{k}}\prod_{k=1}^{r}\Biggl{(}q_{k,-}^{(-z_{f_{k}}/2\pi{\rm i}+d_{k})}\prod_{i<k}((-z_{f_{k}}+z_{f_{i}})/2\pi{\rm i}+d_{k}-d_{i})$
+×$\quad{}\times\prod_{j=1}^{n}\Gamma(1+(z_{f_{k}}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}-d_{k})^{-1}\Gamma(1+(-z_{f_{k}}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}+d_{k})^{-1}\Biggr{)}.$
 Then for $\delta^{-}\in D_{-}$, after choosing an ordering on the elements and viewing $\delta^{-}$ as a function, we have
-(5.1.3) $H^{-}_{\beta,\delta^{-}}:=H_{X_{-},\beta}|_{B_{\delta^{-}}}=\frac{\pi^{r(r-1)}}{\prod_{i<k\leq r}\sin((z_{\delta^{-}_{i}}-z_{\delta^{-}_{k}})/2\mathbbm{i})}K^{-}_{\beta,\delta^{-}}$
+$H^{-}_{\beta,\delta^{-}}:=H_{X_{-},\beta}|_{B_{\delta^{-}}}=\frac{\pi^{r(r-1)/2}}{\prod_{i<k\leq r}\sin\bigl{(}\bigl{(}z_{\delta^{-}_{i}}-z_{\delta^{-}_{k}}\bigr{)}/2{\rm i}\bigr{)}}K^{-}_{\beta,\delta^{-}}.$ (5.3)
 The computation of the analytic continuation of $K_{\beta,\delta}^{+}$ is almost identical to that of $H_{\beta,T,f^{+}}$ from the previous section.
-In this case the path of analytic continuation will be given by $\gamma_{1}\star\cdots\star\gamma_{r}$ where $\gamma_{k}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$ and in the $\log(q_{k,+})$ plane, will go through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi\mathbbm{i}$ when $\mathfrak{R}(\log(q_{k,+}))$ is close to 0.
+In this case the path of analytic continuation will be given by $\gamma_{1}\star\cdots\star\gamma_{r}$ where $\gamma_{k}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$ and in the $\log(q_{k,+})$ plane, will go through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi{\rm i}$ when $\mathfrak{Re}(\log(q_{k,+}))$ is close to 0.
 See Appendix A for details.
 Let
-$C^{K}_{f^{-},f^{+}}=\prod_{i=1}^{r}e^{(n-r)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+x_{f^{+}_{i}})/2\mathbbm{i})}{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+z_{f^{-}_{i}})/2\mathbbm{i})}$
+$C^{K}_{f^{-},f^{+}}=\prod_{i=1}^{r}{\rm e}^{(n-r)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin\bigl{(}\bigl{(}-z_{j}+x_{f^{+}_{i}}\bigr{)}/2{\rm i}\bigr{)}}{\prod_{j\neq{f^{-}_{i}}}\sin\bigl{(}\bigl{(}-z_{j}+z_{f^{-}_{i}}\bigr{)}/2{\rm i}\bigr{)}}.$
 Generalizing the computation from Appendix A, we see that the analytic continuation of $K^{+}_{\beta,\delta^{+}}$ along the path $\gamma_{1}\star\cdots\star\gamma_{r}$ is given by
-(5.1.4) $\sum_{f^{-}\in F_{-}}C^{K}_{f^{-},\delta^{+}}K^{-}_{\beta,f^{-}},$
-after identifying $q_{k,+}^{-1}=q_{k,-}$. We emphasize that here the sum is over all functions $f^{-}:\{1,\ldots,r\}\to\{1,\ldots,n\}$.
-Then by (5.1.2) and (5.1.3), if we define
-$C^{H}_{f^{-},\delta^{+}}=C^{K}_{f^{-},\delta^{+}}\prod_{i<k\leq r}\frac{\sin((z_{f^{-}_{i}}-z_{f^{-}_{k}})/2\mathbbm{i})}{\sin((x_{\delta^{+}_{i}}-x_{\delta^{+}_{k}})/2\mathbbm{i})},$
+$\sum_{f^{-}\in F_{-}}C^{K}_{f^{-},\delta^{+}}K^{-}_{\beta,f^{-}},$ (5.4)
+after identifying $q_{k,+}^{-1}=q_{k,-}$. We emphasize that here the sum is over all functions $f^{-}\colon\{1,\dots,r\}\allowbreak\to\{1,\dots,n\}$.
+Then by (5.2) and (5.3), if we define
+$C^{H}_{f^{-},\delta^{+}}=C^{K}_{f^{-},\delta^{+}}\prod_{i<k\leq r}\frac{\sin\bigl{(}\bigl{(}z_{f^{-}_{i}}-z_{f^{-}_{k}}\bigr{)}/2{\rm i}\bigr{)}}{\sin\bigl{(}\bigl{(}x_{\delta^{+}_{i}}-x_{\delta^{+}_{k}}\bigr{)}/2{\rm i}\bigr{)}},$
 the analytic continuation of $H^{+}_{\beta,\delta^{+}}$ is given by
@@
 Note that the numerator of $C^{H}_{f^{-},\delta^{+}}$ vanishes unless $f^{-}_{i}\neq f^{-}_{k}$ for $i\neq k$.
-Therefore we can replace the summation above to be only over all injective functions $\delta^{-}\in In_{-}$:
-(5.1.5) $\widetilde{H^{+}_{\beta,\delta^{+}}}=\sum_{\delta^{-}\in In_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}.$
-In order to connect the above computation to the actual $I$-functions for $X_{+}$ and $X_{-}$ via (5.1.1), we must homotope the path $\gamma_{1}\star\cdots\star\gamma_{r}$ of analytic continuation to a new path which lies entirely within the diagonal $q_{1,+}=\cdots=q_{r,+}$.
+Therefore, we can replace the summation above to be only over all injective functions $\delta^{-}\in{\rm In}_{-}$:
+$\widetilde{H^{+}_{\beta,\delta^{+}}}=\sum_{\delta^{-}\in{\rm In}_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}.$ (5.5)
+In order to connect the above computation to the actual $I$-functions for $X_{+}$ and $X_{-}$ via (5.1), we must homotope the path $\gamma_{1}\star\cdots\star\gamma_{r}$ of analytic continuation to a new path which lies entirely within the diagonal $q_{1,+}=\cdots=q_{r,+}$.
 Proposition 5.1.
-The path of analytic continuation $\gamma_{1}\star\cdots\star\gamma_{r}$ used in (5.1.5) is homotopic to a path $\hat{\gamma}$ contained entirely in the locus $q_{1,+}=\cdots=q_{r,+}$, via a homotopy which is
-  1. (1)
+The path of analytic continuation $\gamma_{1}\star\cdots\star\gamma_{r}$ used in (5.5) is homotopic to a path $\hat{\gamma}$ contained entirely in the locus $q_{1,+}=\cdots=q_{r,+}$, via a homotopy which is
+  1. 1)1)1 )
     contained within the domain for which ${H^{+}_{\beta,\delta^{+}}}$ is analytic, and
-  2. (2)
+  2. 2)2)2 )
     independent of $\beta$ and $\delta^{+}$.
@@
 Upon setting $q_{+}^{-1}=q_{-}$, we obtain the following:
-(5.1.6) $\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in In_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}.$
+$\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in{\rm In}_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}.$ (5.6)
 Proof.
-First consider the function $f_{\beta,\delta,k}(q_{k,+})$ defined as
-$\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi\mathbbm{i}+e_{k})}(-1)^{(r-1)e_{k}}}{\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})}$
-$=$ $\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\left(\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi\mathbbm{i}+e_{k})}(-1)^{(n+r-1)e_{k}+\sum_{j=1}^{n}\beta\cdot M_{j}}\prod_{j=1}^{n}\sin((x_{\delta_{k}}-z_{j})/2\mathbbm{i})}{\pi^{n}}\right.$
-$\cdot\left.\prod_{j=1}^{n}\frac{\Gamma((x_{\delta_{k}}-z_{j})/2\pi\mathbbm{i}+\beta\cdot M_{j}+e_{k})}{\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})}\right)$
+First, consider the function $f_{\beta,\delta,k}(q_{k,+})$ defined as
+$\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi{\rm i}+e_{k})}(-1)^{(r-1)e_{k}}}{\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})}$
+$\qquad{}=\sum_{e_{k}\geq-\beta\cdot L_{\delta_{k}}}\left(\frac{q_{k,+}^{(x_{\delta_{k}}/2\pi{\rm i}+e_{k})}(-1)^{(n+r-1)e_{k}+\sum_{j=1}^{n}\beta\cdot M_{j}}\prod_{j=1}^{n}\sin((x_{\delta_{k}}-z_{j})/2{\rm i})}{\pi^{n}}\right.$
+[…]
 This satisfies the differential equation
-$\left[\prod_{j=1}^{n}(\partial_{k}-x_{j}/2\pi\mathbbm{i}+\beta\cdot L_{j})-q_{k,+}(-1)^{n-r+1}\prod_{j=1}^{n}(\partial_{k}-z_{j}/2\pi\mathbbm{i}+\beta\cdot M_{j})\right]\Phi(q_{k,+})=0,$
+$\left[\prod_{j=1}^{n}(\partial_{k}-x_{j}/2\pi{\rm i}+\beta\cdot L_{j})-q_{k,+}(-1)^{n-r+1}\prod_{j=1}^{n}(\partial_{k}-z_{j}/2\pi{\rm i}+\beta\cdot M_{j})\right]\Phi(q_{k,+})=0,$
 where $\partial_{k}=q_{k,+}\frac{\partial}{\partial q_{k,+}}$.
-This differential equation has singularities at $0,\infty,$ and $(-1)^{n-r+1}$.
-Following (A.0.3), our path of analytic continuation in the $\log(q_{k,+})$ plane goes through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi\mathbbm{i}$ when $\mathfrak{Re}(\log(q_{k,+}))$ is close to 0.
-Consider the product
-$\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})$
-viewed as an analytic function of the variables $\log(q_{1,+}),\ldots,\log(q_{r,+})$. We see that it may be analytically continued everywhere except along the poles:
-(5.1.7) $\log(q_{k,+})=(n-r+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}.$
+This differential equation has singularities at 00, $\infty$,∞ , and $(-1)^{n-r+1}$.
+Following (A.3), our path of analytic continuation in the $\log(q_{k,+})$ plane goes through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi{\rm i}$ when $\mathfrak{Re}(\log(q_{k,+}))$ is close to 0.
+Consider the product $\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})$ viewed as an analytic function of the variables $\log(q_{1,+}),\allowbreak\dots,\log(q_{r,+})$.
+We see that it may be analytically continued everywhere except along the poles:
+$\log(q_{k,+})=(n-r+1)\pi{\rm i}+2\pi{\rm i}\mathbb{Z}.$ (5.7)
 Define the differential operator
-$\hat{\Delta}=\prod_{k=1}^{r}\left(\prod_{i<k}(-\partial_{k}+\partial_{i})\right).$
+$\hat{\Delta}=\prod_{k=1}^{r}\biggl{(}\prod_{i<k}(-\partial_{k}+\partial_{i})\biggr{)}.$
 Then
-$K^{+}_{\beta,\delta}=\hat{\Delta}\left(\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})\right),$
-and therefore $K^{+}_{\beta,\delta}$ has the same poles at (5.1.7).
+$K^{+}_{\beta,\delta}=\hat{\Delta}\Biggl{(}\prod_{k=1}^{r}f_{\beta,\delta,k}(q_{k,+})\Biggr{)},$
+and therefore $K^{+}_{\beta,\delta}$ has the same poles at (5.7).
 […]
-Figure 1.
-Let $\gamma_{1}$ denote the path that keeps $\log(q_{2,+}),\ldots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ from Figure 1 in $\log(q_{1,+})$.
-Let $\gamma_{2}$ denote the path that keeps $\log(q_{1,+})$ constant (equal to $\infty$), keeps $\log(q_{3,+}),\ldots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ in $\log(q_{2,+})$.
+Figure 1:
+Let $\gamma_{1}$ denote the path that keeps $\log(q_{2,+}),\dots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ from Figure 1 in $\log(q_{1,+})$.
+Let $\gamma_{2}$ denote the path that keeps $\log(q_{1,+})$ constant (equal to $\infty$), keeps $\log(q_{3,+}),\dots,\log(q_{r,+})$ constant (equal to $\infty$- ∞) and follows $\gamma$ in $\log(q_{2,+})$.
 Define $\gamma_{k}$ for $k\leq r$ similarly.
-The analytic continuation for $K^{+}_{\beta,\delta}$ computed in (5.1.4) involved the concatenated path
-$\gamma_{1}\star\cdots\star\gamma_{r}.$
+The analytic continuation for $K^{+}_{\beta,\delta}$ computed in (5.4) involved the concatenated path $\gamma_{1}\star\cdots\star\gamma_{r}$.
 We must show that this composition of paths is homotopy equivalent to a path $\hat{\gamma}$ which follows $\gamma$ in the plane $\log(q_{1,+})=\cdots=\log(q_{r,+})$ via a homotopy which is contained in
-$\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}\}_{k=1}^{r}.$
+$\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi{\rm i}+2\pi{\rm i}\mathbb{Z}\}_{k=1}^{r}.$
 We verify this in the case $r=2$. An induction argument proves the claim for general $r$. First, define a path $\widetilde{\gamma}_{1}$ which agrees with $\gamma_{1}$ except that
-  • for all points on the path such that $\mathfrak{Re}(\log(q_{1,+}))\geq 0$ we let $\mathfrak{Im}(\log(q_{1,+}))=(n-r)\pi\mathbbm{i}$;
+  • for all points on the path such that $\mathfrak{Re}(\log(q_{1,+}))\geq 0$ we let $\mathfrak{Im}(\log(q_{1,+}))=(n-r)\pi{\rm i}$;
   • while the value of $\mathfrak{Re}(\log(q_{2,+}))$ remains constant, the value of $\mathfrak{Im}(\log(q_{2,+}))$ is defined to be equal to $\mathfrak{Im}(\log(q_{1,+}))$ for all points on the path.
 Similarly, define a path $\widetilde{\gamma}_{2}$ which agrees with $\gamma_{2}$ except that
-  • for all points on the path such that $\mathfrak{Re}(\log(q_{2,+}))\leq 0$ we let $\mathfrak{Im}(\log(q_{2,+}))=(n-r)\pi\mathbbm{i}$;
+  • for all points on the path such that $\mathfrak{Re}(\log(q_{2,+}))\leq 0$ we let $\mathfrak{Im}(\log(q_{2,+}))=(n-r)\pi{\rm i}$;
   • while the value of $\mathfrak{Re}(\log(q_{1,+}))$ remains constant, the value of $\mathfrak{Im}(\log(q_{1,+}))$ is defined to be equal to $\mathfrak{Im}(\log(q_{2,+}))$ for all points on the path.
@@
 […]
-Figure 2. The projections of the paths $\widetilde{\gamma}_{1}$ and $\widetilde{\gamma}_{2}$ to the $\log(q_{1,+})$ and $\log(q_{2,+})$ planes respectively.
-It is clear that $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ is homotopic to $\gamma_{1}\star\gamma_{2}$ in $\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}\}_{k=1}^{r}.$ By the construction, the imaginary parts $\mathfrak{Im}(\log(q_{1,+}))$ and $\mathfrak{Im}(\log(q_{2,+}))$ are equal for all points on the path $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$.
-Finally, we homotope $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ further by modifying the real part of the path so that $\mathfrak{Re}(\log(q_{1,+}))=\mathfrak{Re}(\log(q_{2,+}))$ for all points on the path (see Figure 3).
-This second homotopy is also contained entirely in $\mathbb{C}^{2}\setminus\{\log(q_{k,+})=(n-2+1)\pi\mathbbm{i}+2\pi\mathbbm{i}\mathbb{Z}\}_{k=1}^{2}.$
-The resulting path $\hat{\gamma}$ from $(\log(q_{1,+}),\log(q_{2,+}))=(-\infty,-\infty)$ to (,)($\infty$,$\infty$)( ∞ , ∞ ) lies entirely in the plane $\log(q_{1,+})=\log(q_{2,+})$ as desired.
+Figure 2: The projections of the paths $\widetilde{\gamma}_{1}$ and $\widetilde{\gamma}_{2}$ to the $\log(q_{1,+})$ and $\log(q_{2,+})$ planes, respectively.
+It is clear that $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ is homotopic to $\gamma_{1}\star\gamma_{2}$ in $\mathbb{C}^{r}\setminus\{\log(q_{k,+})=(n-r+1)\pi{\rm i}+2\pi{\rm i}\mathbb{Z}\}_{k=1}^{r}.$ By the construction, the imaginary parts $\mathfrak{Im}(\log(q_{1,+}))$ and $\mathfrak{Im}(\log(q_{2,+}))$ are equal for all points on the path $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$.
+Finally, we homotope $\widetilde{\gamma}_{1}\star\widetilde{\gamma}_{2}$ further by modifying the real part of the path so that we have $\mathfrak{Re}(\log(q_{1,+}))=\mathfrak{Re}(\log(q_{2,+}))$ for all points on the path (see Figure 3).
+This second homotopy is also contained entirely in $\mathbb{C}^{2}\setminus\{\log(q_{k,+})=(n-2+1)\pi{\rm i}+2\pi{\rm i}\mathbb{Z}\}_{k=1}^{2}.$
+The resulting path $\hat{\gamma}$ from $(\log(q_{1,+}),\log(q_{2,+}))=(-\infty,-\infty)$ to (,)($\infty$,$\infty$)( ∞ , ∞ ) lies entirely in the plane $\log(q_{1,+})=\log(q_{2,+})$ as desired. ∎
 $\mathfrak{Re}_{\mathbb{T}}(\log(q_{1,+}))$$\mathfrak{Re}_{\mathbb{T}}(\log(q_{2,+}))$(,($\infty$,$\infty$)( - ∞ , - ∞ )(,($\infty$,$\infty$)( ∞ , ∞ )
-Figure 3. The (real part of the) original path of analytic continuation $\gamma_{1}\star\gamma_{2}$ (in purple) and the new path $\hat{\gamma}$ (in blue).
-5.2. Simplification
+Figure 3: The (real part of the) original path of analytic continuation $\gamma_{1}\star\gamma_{2}$ (in purple) and the new path $\hat{\gamma}$ (in blue).
+5.2 Simplification
 After specializing the Novikov parameters, we observe that $H^{{\pm}}_{\beta,\delta}|_{q_{1,\pm}=\cdots=q_{r,\pm}}=H^{{\pm}}_{\beta,\delta^{\prime}}|_{q_{1,\pm}=\cdots=q_{r,\pm}}$ whenever $\delta$ and $\delta^{\prime}$ are equal up to a permutation.
-This allows for a dramatic simplification to the formula (5.1.6).
+This allows for a dramatic simplification to the formula (5.6).
 The simplification relies on the following identity.
 Lemma 5.2.
-We have the following identity of (bi)-anti-symmetric functions:
-(5.2.1) $\prod_{i=1}^{r}\prod_{l<i}\left((z_{i}-z_{l})(x_{l}-x_{i})\right)=\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l}).$
+We have the following identity of ((((bi))))-anti-symmetric functions:
+$\prod_{i=1}^{r}\prod_{l<i}((z_{i}-z_{l})(x_{l}-x_{i}))=\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l}).$ (5.8)
 Proof.
-Denote the left and right hand sides of (5.2.1) by $L(\underline{x},\underline{z})$ and $R(\underline{x},\underline{z})$ respectively.
+Denote the left- and right-hand sides of (5.8) by $L(\underline{x},\underline{z})$ and $R(\underline{x},\underline{z})$, respectively.
 We first claim that $R(\underline{x},\underline{z})$ is divisible by $(z_{i}-z_{l})$ for $i\neq l$.
@@
 Since $\sigma(l)=\tilde{\sigma}(i)$ and $\sigma(i)=\tilde{\sigma}(l)$, we see that
-$\left(\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l})+\operatorname{sgn}(\tilde{\sigma})\prod_{l=1}^{r}\prod_{j\neq\tilde{\sigma}(l)}(x_{j}-z_{l})\right)|_{z_{i}=z_{l}=z}=0,$
+$\Biggl{(}\operatorname{sgn}(\sigma)\prod_{l=1}^{r}\prod_{j\neq\sigma(l)}(x_{j}-z_{l})+\operatorname{sgn}(\tilde{\sigma})\prod_{l=1}^{r}\prod_{j\neq\tilde{\sigma}(l)}(x_{j}-z_{l})\Biggr{)}\bigg{|}_{z_{i}=z_{l}=z}=0,$
 and the claim follows.
@@
 Corollary 5.3.
-For $\delta^{-}:\{1,\ldots,r\}\to\{1,\ldots,n\}$ an injective function, let $S_{r}\cdot\delta^{-}$ denote the set of all possible functions obtained by pre-composing $\delta^{-}$ with a permutation of $\{1,\ldots,r\}.$ The identity
-(5.2.2) $\sum_{f^{-}\in S_{r}\cdot\delta^{-}}C^{H}_{f^{-},\delta^{+}}=C_{\delta^{-},\delta^{+}}.$
-holds, where $C_{\delta^{-},\delta^{+}}$ was defined in (2.3.4).
+For $\delta^{-}\colon\{1,\dots,r\}\to\{1,\dots,n\}$ an injective function, let $S_{r}\cdot\delta^{-}$ denote the set of all possible functions obtained by pre-composing $\delta^{-}$ with a permutation of $\{1,\dots,r\}.$ The identity
+$\sum_{f^{-}\in S_{r}\cdot\delta^{-}}C^{H}_{f^{-},\delta^{+}}=C_{\delta^{-},\delta^{+}}$ (5.9)
+holds, where $C_{\delta^{-},\delta^{+}}$ was defined in (2.7).
 Proof.
-By reindexing the sets $\{x_{1},\ldots,x_{n}\}$ and $\{z_{1},\ldots,z_{n}\}$, we may assume without loss of generality that $\delta^{-}=\delta^{+}=\{1,\ldots,r\}$.
-Then in this case (5.2.2) simplifies to the equation
-(5.2.3) $\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\left(\prod_{\stackrel{{\scriptstyle 1\leq j\leq n}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2\mathbbm{i})}{\sin((z_{\sigma(k)}-z_{j})/2\mathbbm{i})}\prod_{i<k}\frac{\sin((z_{\sigma(k)}-z_{\sigma(i)})/2\mathbbm{i})}{\sin((x_{k}-x_{i})/2\mathbbm{i})}\right)$
-$=$ $\prod_{k=1}^{r}\prod_{j>r}\frac{\sin((x_{k}-z_{j})/2\mathbbm{i})}{\sin((z_{k}-z_{j})/2\mathbbm{i})}$
+By reindexing the sets $\{x_{1},\dots,x_{n}\}$ and $\{z_{1},\dots,z_{n}\}$, we may assume without loss of generality that $\delta^{-}=\delta^{+}=\{1,\dots,r\}$.
+Then in this case, (5.9) simplifies to the equation
+$\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\Biggl{(}\prod_{\stackrel{{\scriptstyle 1\leq j\leq n}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2{\rm i})}{\sin((z_{\sigma(k)}-z_{j})/2{\rm i})}\prod_{i<k}\frac{\sin\bigl{(}\bigl{(}z_{\sigma(k)}-z_{\sigma(i)}\bigr{)}/2{\rm i}\bigr{)}}{\sin((x_{k}-x_{i})/2{\rm i})}\Biggr{)}$
+$\qquad{}=\prod_{k=1}^{r}\prod_{j>r}\frac{\sin((x_{k}-z_{j})/2{\rm i})}{\sin((z_{k}-z_{j})/2{\rm i})}$
 or, equivalently,
-(5.2.4) $\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\left(\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2\mathbbm{i})}{\sin((z_{\sigma(k)}-z_{j})/2\mathbbm{i})}\prod_{i<k}\frac{\sin((z_{\sigma(k)}-z_{\sigma(i)})/2\mathbbm{i})}{\sin((x_{k}-x_{i})/2\mathbbm{i})}\right)=1.$
+$\sum_{\sigma\in S_{r}}\prod_{k=1}^{r}\Biggl{(}\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{\sin((x_{k}-z_{j})/2{\rm i})}{\sin((z_{\sigma(k)}-z_{j})/2{\rm i})}\prod_{i<k}\frac{\sin((z_{\sigma(k)}-z_{\sigma(i)})/2{\rm i})}{\sin((x_{k}-x_{i})/2{\rm i})}\Biggr{)}=1.$ (5.10)
 Observe that
-$\prod_{k=1}^{r}\left(\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{1}{\sin((z_{\sigma(k)}-z_{j})/2\mathbbm{i})}\prod_{i<k}\sin((z_{\sigma(k)}-z_{\sigma(i)})/2\mathbbm{i})\right)$
-$=$ $\prod_{k=1}^{r}\prod_{j>k}\frac{1}{\sin((z_{\sigma(k)}-z_{\sigma(j)})/2\mathbbm{i})}$
-$=$ $\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{j<k}\frac{1}{\sin((z_{j}-z_{k})/2\mathbbm{i})}$
-and (5.2.4) is then equivalent to the equality
-(5.2.5) $\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\sin((x_{k}-z_{j})/2\mathbbm{i})=\prod_{k=1}^{r}\prod_{i<k}\sin((z_{i}-z_{k})/2\mathbbm{i})\sin((x_{k}-x_{i})/2\mathbbm{i})$
+$\prod_{k=1}^{r}\Biggl{(}\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\frac{1}{\sin((z_{\sigma(k)}-z_{j})/2{\rm i})}\prod_{i<k}\sin((z_{\sigma(k)}-z_{\sigma(i)})/2{\rm i})\Biggr{)}$
+$\qquad{}=\prod_{k=1}^{r}\prod_{j>k}\frac{1}{\sin((z_{\sigma(k)}-z_{\sigma(j)})/2{\rm i})}=\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{j<k}\frac{1}{\sin((z_{j}-z_{k})/2{\rm i})}$
+and (5.10) is then equivalent to the equality
+$\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{k=1}^{r}\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}\sin((x_{k}-z_{j})/2{\rm i})=\prod_{k=1}^{r}\prod_{i<k}\sin((z_{i}-z_{k})/2{\rm i})\sin((x_{k}-x_{i})/2{\rm i}).$ (5.11)
 Noting that
-$\sin((a-b)/2\mathbbm{i})=\frac{e^{a/2+b/2}(e^{a}-e^{b})}{2\mathbbm{i}},$
-under the change of variables $X_{i}=e^{x_{i}},Z_{i}=e^{z_{i}}$, (5.2.5) becomes
-(5.2.6) $\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}(X_{k}-Z_{j})=\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\prod_{i<k}(Z_{i}-Z_{k})(X_{k}-X_{i}).$
-Lemma 5.2 then gives the desired equality.
-5.3. Main theorem
-Combining (5.1.6) with (5.2.2), we conclude that
-(5.3.1) $\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in D_{-}}C_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}.$
-where the path of analytic continuation is along $\hat{\gamma}$ from Proposition 5.1 and the sum is now over all unordered size $r$ subsets of $\{1,\ldots,n\}$.
-Let $\mathbb{U}_{H}:H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined in (2.3.5).
-Then (5.3.1) implies that:
-(5.3.2) $\mathbb{U}_{H}H_{X_{-},\beta}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}=\widetilde{H_{X_{+},\beta}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}$
+$\sin((a-b)/2{\rm i})=\frac{{\rm e}^{a/2+b/2}\bigl{(}{\rm e}^{a}-{\rm e}^{b}\bigr{)}}{2{\rm i}},$
+under the change of variables $X_{i}={\rm e}^{x_{i}}$, $Z_{i}={\rm e}^{z_{i}}$, (5.11) becomes
+$\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\sum_{\sigma\in S_{r}}\operatorname{sgn}(\sigma)\prod_{\stackrel{{\scriptstyle 1\leq j\leq r}}{{j\neq{\sigma(k)}}}}(X_{k}-Z_{j})=\prod_{k=1}^{r}(Z_{k}X_{k})^{(r-1)/2}\prod_{i<k}(Z_{i}-Z_{k})(X_{k}-X_{i}).$
+Lemma 5.2 then gives the desired equality. ∎
+5.3 Main theorem
+Combining (5.6) with (5.9), we conclude that
+$\widetilde{H^{+}}_{\beta,\delta^{+}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}=\sum_{\delta^{-}\in D_{-}}C_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}},$ (5.12)
+where the path of analytic continuation is along $\hat{\gamma}$ from Proposition 5.1 and the sum is now over all unordered size $r$ subsets of $\{1,\dots,n\}$.
+Let $\mathbb{U}_{H}\colon H^{*}_{\mathbb{T}}(X_{-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined in (2.8).
+Then (5.12) implies that
+$\mathbb{U}_{H}H_{X_{-},\beta}|_{q_{1,-}=\cdots=q_{r,-}=q_{-}}=\widetilde{H_{X_{+},\beta}}|_{q_{1,+}=\cdots=q_{r,+}=q_{+}}$ (5.13)
 for all $\beta$, where the path of analytic continuation is along $\hat{\gamma}$ and we set $q_{+}^{-1}=q_{-}$.
 Theorem 5.4.
-The linear transformation $\mathbb{U}$ is symplectic, has a well-defined non-equivariant limit, and satisfies:
+The linear transformation $\mathbb{U}$ is symplectic, has a well-defined non-equivariant limit, and satisfies
 $\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}},$
 where $\widetilde{I_{X_{+}}}$ denotes the analytic continuation of $I_{X_{+}}$ along the path $\hat{\gamma}$ and we set $q_{+}^{-1}=q_{-}$. Furthermore, the following diagram commutes:
-${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\mathbb{U}}$
+${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\mathbb{U}}$
 Proof.
-Equation (5.1.1) and the analogous statement for $X_{-}$ together with (5.3.2) immediately implies that $\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}}.$ The fact that $\mathbb{U}$ is symplectic and has a well-defined non-equivariant limit is Corollary 3.6.
+Equation (5.1) and the analogous statement for $X_{-}$ together with (5.13) immediately implies that $\mathbb{U}I_{X_{-}}=\widetilde{I_{X_{+}}}.$ The fact that $\mathbb{U}$ is symplectic and has a well-defined non-equivariant limit is Corollary 3.6.
 The commuting diagram the follows from Proposition 2.10.
-Remark 5.5 (Big $I$-functions and the comparison with [LSW]).
+∎
+Remark 5.5 (big $I$-functions and the comparison with [28]).
 In the special case that $B$ is a point, the wall crossing from $X_{+}$ to $X_{-}$ is the standard Grassmann flop.
-In this case, by Webb [Web23, Section 6.3], one can obtain big $I$-functions of $X_{\pm}$, denoted $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$, as a modification of $I_{X_{\pm}}(q_{\pm},z)$.
+In this case, by Webb [33, Section 6.3], one can obtain big $I$-functions of $X_{\pm}$, denoted $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$, as a modification of $I_{X_{\pm}}(q_{\pm},z)$.
 The respective big $I$-functions generate the entire Lagrangian cones $\mathcal{L}_{X_{\pm}}$, and therefore fully determine the genus-zero Gromov–Witten theory of $X_{\pm}$.
 A straightforward generalization of the computations above shows that the transformation $\mathbb{U}$ from Theorem 5.4 also identifies these big $I$-functions:
-(5.3.3) $\mathbb{U}\mathbb{I}_{X_{-}}(q_{-},\mathbf{x},z)=\widetilde{\mathbb{I}_{X_{+}}}(q_{+},\mathbf{x},z),$
+$\mathbb{U}\mathbb{I}_{X_{-}}(q_{-},\mathbf{x},z)=\widetilde{\mathbb{I}_{X_{+}}}(q_{+},\mathbf{x},z),$ (5.14)
 where $\widetilde{\mathbb{I}_{X_{+}}}$ denotes the analytic continuation in $q_{+}$ as in Theorem 5.4.
-In the case that $B$ is a point, Lutz–Shafi–Webb [LSW] have obtained a different proof that there exists a symplectic transformation, which we will denote here by $\mathbb{U}^{LSW}$, identifying the big $I$-functions of $X_{\pm}$ after the analytic continuation.
-Rather than comparing to the Fourier–Mukai transform $\mathbb{FM}$ however, the symplectic transformation in $\mathbb{U}^{LSW}$ is obtained directly from wall crossing of the associated abelian GIT quotients $X_{T,\pm}$.
+In the case that $B$ is a point, Lutz–Shafi–Webb [28] have obtained a different proof that there exists a symplectic transformation, which we will denote here by $\mathbb{U}^{\rm LSW}$, identifying the big $I$-functions of $X_{\pm}$ after the analytic continuation.
+Rather than comparing to the Fourier–Mukai transform $\mathbb{FM}$ however, the symplectic transformation in $\mathbb{U}^{\rm LSW}$ is obtained directly from wall crossing of the associated abelian GIT quotients $X_{T,\pm}$.
 Their result is related to Theorem 5.4 as follows.
-Let $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$ be the big $I$-functions of the previous paragraph, and let $g^{*}\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)$ denote the $I$-functions used in [LSW].
+Let $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)$ be the big $I$-functions of the previous paragraph, and let $g^{*}\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)$ denote the $I$-functions used in [28].
 Comparing explicit formulas, we have
-(5.3.4) $\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)|_{q_{\pm}=\mathbf{y}_{\pm}}=\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)|_{\mathbf{q}=1},$
+$\mathbb{I}_{X_{\pm}}(q_{\pm},\mathbf{x},z)|_{q_{\pm}=\mathbf{y}_{\pm}}=\mathbb{I}^{\pm}_{G}(\mathbf{y}_{\pm},\mathbf{x},z)|_{\mathbf{q}=1}.$ (5.15)
 Under the identification $q_{\pm}\mapsto\mathbf{y}_{\pm}$, the path $\delta$ of analytic continuation used in loc.
@@
 matches that given in Proposition 5.1.
-In [LSW] the abelian/non-abelian correspondence is used to show that there is a unique linear map $\mathbb{U}^{LSW}$ satisfying
-$\mathbb{U}^{LSW}\mathbb{I}^{-}_{G}(\mathbf{y}_{-}\mathbf{x},z)|_{\mathbf{q}=1}=\widetilde{\mathbb{I}^{+}_{G}}(\mathbf{y}_{+}\mathbf{x},z)|_{\mathbf{q}=1}.$
-From this, (5.3.4), and (5.3.3), we conclude that $\mathbb{U}=\mathbb{U}^{LSW}$, thus providing a consistency check between the two results.
-5.4. Central charge
-Motivated by the notion of central charge as given in e.g. [KA23, Iri09], we make a similar definition. The above wall crossing result takes an especially simple form using this language.
+In [28], the abelian/non-abelian correspondence is used to show that there is a unique linear map $\mathbb{U}^{\rm LSW}$ satisfying
+$\mathbb{U}^{\rm LSW}\mathbb{I}^{-}_{G}(\mathbf{y}_{-}\mathbf{x},z)|_{\mathbf{q}=1}=\widetilde{\mathbb{I}^{+}_{G}}(\mathbf{y}_{+}\mathbf{x},z)|_{\mathbf{q}=1}.$
+From this, (5.15) and (5.14), we conclude that $\mathbb{U}=\mathbb{U}^{\rm LSW}$, thus providing a consistency check between the two results.
+5.4 Central charge
+Motivated by the notion of central charge as given, e.g., in [1, 22], we make a similar definition. The above wall crossing result takes an especially simple form using this language.
 Definition 5.6.
 For $E\in K^{0}_{\mathbb{T}}(X_{\pm})$, define the relative quasimap central charge
-$Z_{X_{\pm}}(E)(Q,q_{\pm},z):=\left\langle I_{X_{\pm}}(Q,q_{\pm},-z),\Psi_{\pm}(E)\right\rangle_{X_{\pm}}.$
+$Z_{X_{\pm}}(E)(Q,q_{\pm},z):=\langle I_{X_{\pm}}(Q,q_{\pm},-z),\Psi_{\pm}(E)\rangle_{X_{\pm}}.$
 With this definition, we obtain as an immediate corollary that the central charge of an element $E$ and of its Fourier–Mukai transform are related by analytic continuation.
 Corollary 5.7.
-For $E\in K^{0}_{\mathbb{T}}(X_{\pm})$, the analytic continuation of $Z_{X_{+}}(\mathbb{FM}(E))(Q,q_{+},z)$ to $q_{+}=\infty$ is $Z_{X_{-}}(E)(Q,q_{-},z)$.
+For $E\in K^{0}_{\mathbb{T}}(X_{-})$, the analytic continuation of $Z_{X_{+}}(\mathbb{FM}(E))(Q,q_{+},z)$ to ${q_{+}\!=\!\infty}$ is $Z_{X_{-}}(E)(Q,q_{-},z)$.
 Proof.
-We use the diagram in Theorem 4.2 and the fact that $\mathbb{U}$ preserves the symplectic pairing.
-$\widetilde{Z_{X_{+}}}(\mathbb{FM}(E))(Q,q_{+},z)=$ $\left\langle\widetilde{I_{X_{+}}}(Q,q_{+},-z),\Psi_{+}(\mathbb{FM}(E))\right\rangle_{X_{+}}$
-$=$ $\left\langle\mathbb{U}I_{X_{-}}(Q,q_{-},-z),\mathbb{U}\Psi_{-}(E)\right\rangle_{X_{+}}$
-$=$ $\left\langle I_{X_{-}}(Q,q_{-},-z),\Psi_{-}(E)\right\rangle_{X_{-}}$
-$=$ $Z_{X_{-}}(E)(Q,q_{-},z).$
+We use the diagram in Theorem 4.2 and the fact that $\mathbb{U}$ preserves the symplectic pairing,
+$\widetilde{Z_{X_{+}}}(\mathbb{FM}(E))(Q,q_{+},z)$ $=\bigl{\langle}\widetilde{I_{X_{+}}}(Q,q_{+},-z),\Psi_{+}(\mathbb{FM}(E))\bigr{\rangle}_{X_{+}}$
+$=\langle\mathbb{U}I_{X_{-}}(Q,q_{-},-z),\mathbb{U}\Psi_{-}(E)\rangle_{X_{+}}$
+$=\langle I_{X_{-}}(Q,q_{-},-z),\Psi_{-}(E)\rangle_{X_{-}}=Z_{X_{-}}(E)(Q,q_{-},z).$
 Appendix A Proof of Proposition 4.1
 Using the identity
-(A.0.1) $\Gamma(1-z)\Gamma(z)=\frac{\pi}{\sin(\pi z)},$
+$\Gamma(1-z)\Gamma(z)=\frac{\pi}{\sin(\pi z)},$ (A.1)
 the denominators of $H_{X_{+},\beta}$ may be rewritten as
-$\frac{\prod_{i=1}^{n}\sin((-z_{i}-y)/2\mathbbm{i}))}{\pi^{n}}\frac{\prod_{i=1}^{n}(-1)^{\beta\cdot M_{i}+e}\Gamma((-z_{i}-y)/2\pi\mathbbm{i}+\beta\cdot M_{i}+e)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi\mathbbm{i}+\beta\cdot L_{j}+e)}.$
+$\frac{\prod_{i=1}^{n}\sin((-z_{i}-y)/2{\rm i}))}{\pi^{n}}\frac{\prod_{i=1}^{n}(-1)^{\beta\cdot M_{i}+e}\Gamma((-z_{i}-y)/2\pi{\rm i}+\beta\cdot M_{i}+e)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}-y)/2\pi{\rm i}+\beta\cdot L_{j}+e)}.$
 Let $B_{l}^{+}$ denote the $l$th fixed locus, isomorphic to $B$.
@@
 Thus $H_{\beta,l}^{+}:=H_{X_{+},\beta}|_{B_{l}^{+}}$ is given by
-(A.0.2) $\sum_{e\geq-\beta\cdot L_{l}}q_{+}^{e+x_{l}/2\pi\mathbbm{i}}\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi^{n}}\frac{\prod_{i=1}^{n}(-1)^{\beta\cdot M_{i}+e}\Gamma((-z_{i}+x_{l})/2\pi\mathbbm{i}+\beta\cdot M_{i}+e)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+x_{l})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e)}$
+$\sum_{e\geq-\beta\cdot L_{l}}q_{+}^{e+x_{l}/2\pi{\rm i}}$
+[…]
 This satisfies the differential equation
-$\left[\prod_{j=1}^{n}(\partial_{q}-x_{j}/2\pi\mathbbm{i}+\beta\cdot L_{j})-q_{+}(-1)^{n}\prod_{j=1}^{n}(\partial_{q}-z_{j}/2\pi\mathbbm{i}+\beta\cdot M_{j})\right]\Phi(q_{+})=0,$
+$\left[\prod_{j=1}^{n}(\partial_{q}-x_{j}/2\pi{\rm i}+\beta\cdot L_{j})-q_{+}(-1)^{n}\prod_{j=1}^{n}(\partial_{q}-z_{j}/2\pi{\rm i}+\beta\cdot M_{j})\right]\Phi(q_{+})=0,$
 where $\partial_{q}=q_{+}\frac{\partial}{\partial q_{+}}$.
 This differential equation has singularities at $0,\infty,$ and $(-1)^{n}$.
-Following [CIJ18], our path $\gamma$ of analytic continuation in the $\log(q_{+})$ will go through $\mathfrak{Im}(\log(q_{+}))=(n-1)\pi\mathbbm{i}$ when $\mathfrak{R}(\log(q_{+}))$ is close to 0.
+Following [14], our path $\gamma$ of analytic continuation in the $\log(q_{+})$ will go through $\mathfrak{Im}(\log(q_{+}))=(n-1)\pi{\rm i}$ when $\mathfrak{Re}(\log(q_{+}))$ is close to 0.
 See Figure 1 for a picture of the path.
 We will use the Mellin–Barnes method to compute this analytic continuation of $H_{\beta,l}^{+}$ to $q_{+}=\infty$.
-We may rewrite (A.0.2) as a contour integral
-(A.0.3) $\frac{1}{2\pi\mathbbm{i}}\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi^{n}}\int_{C}\Gamma(s)\Gamma(1-s)q_{+}^{s+x_{l}/2\pi\mathbbm{i}}\cdot$
-$e^{-\pi\mathbbm{i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot s)}\frac{\prod_{i=1}^{n}\Gamma((-z_{i}+x_{l})/2\pi\mathbbm{i}+\beta\cdot M_{i}+s)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+x_{l})/2\pi\mathbbm{i}+\beta\cdot L_{j}+s)},$
+We may rewrite (A.2) as a contour integral
+$\frac{1}{2\pi{\rm i}}\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2{\rm i})}{\pi^{n}}$ (A.3)
+×$\times\!\int_{C}\Gamma(s)\Gamma(1-s)q_{+}^{s+x_{l}/2\pi{\rm i}}{\rm e}^{-\pi{\rm i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot s)}\frac{\prod_{i=1}^{n}\Gamma((-z_{i}+x_{l})/2\pi{\rm i}+\beta\cdot M_{i}+s)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+x_{l})/2\pi{\rm i}+\beta\cdot L_{j}+s)},$
 where $C$ encloses all integers $e\geq-\beta\cdot L_{l}$.
 Closing the curve on the other side, we obtain residues at integers $d<-\beta\cdot L_{l}$ and at
-$-(-z_{k}+x_{l})/2\pi\mathbbm{i}-d\hskip 14.22636pt1\leq k\leq n,d\in\mathbb{Z}_{\geq\beta\cdot M_{k}}.$
+$-(-z_{k}+x_{l})/2\pi{\rm i}-d,\qquad 1\leq k\leq n,d\in\mathbb{Z}_{\geq\beta\cdot M_{k}}.$
 The residues at integers $d\leq-\beta\cdot L_{l}$ vanish. The integral is then equal to minus the sum of the remaining residues:
-(A.0.4) $-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi\mathbbm{i}-d)\Gamma(1+(-z_{k}+x_{l})/2\pi\mathbbm{i}+d)q_{+}^{z_{k}/2\pi\mathbbm{i}-d}$
-$\cdot$ $\frac{e^{-\pi\mathbbm{i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot(-(-z_{k}+x_{l})/2\pi\mathbbm{i}-d)+d-\beta\cdot M_{k})}}{(d-\beta\cdot M_{k})!}\frac{\prod_{i\neq k}\Gamma((-z_{i}+z_{k})/2\pi\mathbbm{i}+\beta\cdot M_{i}-d)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)}$
-$\cdot$ $\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi^{n}}$
-$=$ $-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi\mathbbm{i})\Gamma(1+(-z_{k}+x_{l})/2\pi\mathbbm{i})q_{+}^{z_{k}/2\pi\mathbbm{i}-d}$
-$\cdot$ $\frac{e^{(n-1)(x_{l}-z_{k})/2}}{(d-\beta\cdot M_{k})!\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\prod_{i\neq k}\Gamma(1-(-z_{i}+z_{k})/2\pi\mathbbm{i}-\beta\cdot M_{i}+d)}$
-$\cdot$ $\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\pi\prod_{i\neq k}\sin((-z_{i}+z_{k})/2\mathbbm{i})}$
-$=$ $\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}e^{(n-1)(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2\mathbbm{i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2\mathbbm{i})}$
-$\cdot$ $\frac{q_{+}^{z_{k}/2\pi\mathbbm{i}-d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\prod_{i=1}^{n}\Gamma(1-(-z_{i}+z_{k})/2\pi\mathbbm{i}-\beta\cdot M_{i}+d)}$
-where we have again simplified using the Gamma function identity (A.0.1).
-Let $B_{k}^{-}$ denote the $k$th fixed locus of $X_{-}$. By (4.1.3), observe that $H_{\beta,k}^{-}:=H_{X_{-},\beta}|_{B_{k}^{-}}$ is given by
-$\sum_{d\geq\beta\cdot M_{k}}\frac{q_{-}^{-z_{k}/2\pi\mathbbm{i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}-z_{k})/2\pi\mathbbm{i}-\beta\cdot M_{j}+d)}.$
-The result then follows immediately from (4.1.5).
+$-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi{\rm i}-d)\Gamma(1+(-z_{k}+x_{l})/2\pi{\rm i}+d)q_{+}^{z_{k}/2\pi{\rm i}-d}$
+×$\times\frac{{\rm e}^{-\pi{\rm i}(\sum_{i=1}^{n}\beta\cdot M_{i}+(n-1)\cdot(-(-z_{k}+x_{l})/2\pi{\rm i}-d)+d-\beta\cdot M_{k})}}{(d-\beta\cdot M_{k})!}\frac{\prod_{i\neq k}\Gamma((-z_{i}+z_{k})/2\pi{\rm i}+\beta\cdot M_{i}-d)}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi{\rm i}+\beta\cdot L_{j}-d)}$
+×$\times\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2{\rm i})}{\pi^{n}}$
+$=-\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}\Gamma(-(-z_{k}+x_{l})/2\pi{\rm i})\Gamma(1+(-z_{k}+x_{l})/2\pi{\rm i})q_{+}^{z_{k}/2\pi{\rm i}-d}$
+×$\times\frac{{\rm e}^{(n-1)(x_{l}-z_{k})/2}}{(d-\beta\cdot M_{k})!\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi{\rm i}+\beta\cdot L_{j}-d)}$
+×$\times\frac{\prod_{i=1}^{n}\sin((-z_{i}+x_{l})/2{\rm i})}{\prod_{i\neq k}\Gamma(1-(-z_{i}+z_{k})/2\pi{\rm i}-\beta\cdot M_{i}+d)\cdot\pi\prod_{i\neq k}\sin((-z_{i}+z_{k})/2{\rm i})}$
+$=\sum_{1\leq k\leq n}\sum_{d\geq\beta\cdot M_{k}}{\rm e}^{(n-1)(x_{l}-z_{k})/2}\frac{\prod_{i\neq k}\sin((-z_{i}+x_{l})/2{\rm i})}{\prod_{i\neq k}\sin((-z_{i}+z_{k})/2{\rm i})}$
+×$\times\frac{q_{+}^{z_{k}/2\pi{\rm i}-d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi{\rm i}+\beta\cdot L_{j}-d)\prod_{i=1}^{n}\Gamma(1-(-z_{i}+z_{k})/2\pi{\rm i}-\beta\cdot M_{i}+d)},$
+where we have again simplified using the Gamma function identity (A.1).
+Let $B_{k}^{-}$ denote the $k$th fixed locus of $X_{-}$. By (4.1), observe that $H_{\beta,k}^{-}:=H_{X_{-},\beta}|_{B_{k}^{-}}$ is given by
+$\sum_{d\geq\beta\cdot M_{k}}\frac{q_{-}^{-z_{k}/2\pi{\rm i}+d}}{\prod_{j=1}^{n}\Gamma(1+(-x_{j}+z_{k})/2\pi{\rm i}+\beta\cdot L_{j}-d)\Gamma(1+(z_{j}-z_{k})/2\pi{\rm i}-\beta\cdot M_{j}+d)}.$
+The result then follows immediately from (4.3).
+Acknowledgements
+We would like to thank Tom Coates, Wendelin Lutz, Jeongseok Oh, Yongbin Ruan, Ed Segal, Qaasim Shafi, and Rachel Webb for the useful discussions.
+We also thank the anonymous referees for valuable comments and suggestions on a previous draft.
+N.
+Priddis is supported by the Simons Foundation Travel Grant 586691.
+M.
+Shoemaker is supported by the Simons Foundation Travel Grant 958189.
+Y.
+Wen is supported by a KIAS Individual Grant (MG083902) at the Korea Institute for Advanced Study.